### v1

# Semileptonic decays $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ with the SU(3) flavor analysis

Abstract
Semileptonic three-body $D\to A\ell^{+}\nu_{\ell}$ decays, nonleptonic $A\to VP$ decays and semileptonic four-body $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ strong or electromagnetic decays are analyzed by the SU(3) flavor symmetry/breaking approach, where $A$ denotes the axial-vector meson, $V$ denotes the vector meson, and $P$ denotes pseudoscalar meson. In terms of the SU(3) flavor symmetry/breaking, the decay amplitudes of the $D\to A\ell^{+}\nu_{\ell}$ decays or the vertex coefficients of the $A\to VP$ decays are related. The relevant nonperturbative parameters are constrained by the present experimental data. Then the branching ratios of the $D\to A\ell^{+}\nu_{\ell}$, $A\to VP$ and $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays are predicted by the constrained parameters. We find that some branching ratios are large and they might be measured in the near future experiments. Moreover, the sensitivities of the branching ratios of $D\to A\ell^{+}\nu_{\ell}$ and $A\to VP$ decays to the parameters are also shown, and some decay branching ratios are sensitive to the nonperturbative parameters.
I INTRODUCTION
Semileptonic decays of the charmed mesons are very useful for determining the quark mixing parameters, the values of the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements, and searching for new physics beyond the standard model. Among the semileptonic four-body decays $D\to VP\ell^{+}\nu_{\ell}$, only the branching ratios of $D^{+}\to b^{0}_{1}(b_{1}^{0}\to\omega\pi^{0})e^{+}\nu_{e}$ and $D^{0}\to b^{-}_{1}(b_{1}^{-}\to\omega\pi^{-})e^{+}\nu_{e}$ decays have the experiential upper limits PDG2022 . Nevertheless, two semileptonic three-body decays $D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e}$ and $D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e}$ have been found by the CLEO Collaboration CLEO:2007oer and they have been much improved by the BESIII Collaboration BESIII:2019eao ; BESIII:2021uqr . The six nonleptonic decays $K_{1}\to VP$ have been measured by the ACCMOR Collaboration ACCMOR:1981yww . One can use them to study $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays.
For the semileptonic decays, since the leptons do not participate in the strong interaction, the weak and strong dynamics can be separated, the theoretical description of the semileptonic decays are relatively simple. The hadronic transition form factors contain all the strong dynamics in the initial and final hadrons, and they are crucial for testing the theoretical calculations of the involved strong interaction. The structure of the $D\to VP$ form factors is more complicated than for the $D\to V$ or $D\to P$ form factors, nevertheless, we have not found their systematical calculations up to now. In the absence of reliable calculations, the symmetry analysis can provide very useful information about the decays. SU(3) flavor symmetry is independent of the detailed dynamics, it can provide us an opportunity to relate different decay modes. Even though SU(3) flavor symmetry is only an approximate symmetry, it still gives some valuable information about the decays. SU(3) flavor symmetry has been widely used to study hadron decays, including b-hadron decays He:1998rq ; He:2000ys ; Fu:2003fy ; Hsiao:2015iiu ; He:2015fwa ; He:2015fsa ; Deshpande:1994ii ; Gronau:1994rj ; Gronau:1995hm ; Shivashankara:2015cta ; Zhou:2016jkv ; Cheng:2014rfa ; Wang:2021uzi ; Wang:2020wxn , c-hadron decays Wang:2021uzi ; Wang:2020wxn ; Grossman:2012ry ; Pirtskhalava:2011va ; Cheng:2012xb ; Savage:1989qr ; Savage:1991wu ; Altarelli:1975ye ; Lu:2016ogy ; Geng:2017esc ; Geng:2018plk ; Geng:2017mxn ; Geng:2019bfz ; Wang:2017azm ; Wang:2019dls ; Wang:2017gxe ; Muller:2015lua .
The semileptonic $D\to A\ell^{+}\nu_{\ell}$ decays have been studied, for examples, by the QCD sum rules approach Momeni:2020zrb ; Momeni:2019uag ; Zuo:2016msr ; Hu:2021lkl ; Khosravi:2008jw , the light-cone sum rules Huang:2021owr , the quark models Isgur:1988gb ; Scora:1995ty ; Cheng:2017pcq . The nonleptonic $A\to VP$ decays have been studied by many methods, for examples, the chiral unitary approach Geng:2006yb ; Cabrera:2009ep , quark models Ackleh:1996yt ; Bloch:1999vka ; Du:2021zdg , the effective Lagrangian of the strong interaction Lichard:2010zza ; Li:1995aw ; Li:1995tv ; Wess:1967jq ; Roca:2003uk ; Vojik:2010ua . The four-body semileptonic $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays have also been studies by the unitary extensions of chiral perturbation theory Wang:2020pyy . In this work, we will analyze the $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ four-body semileptonic decays by using the SU(3) flavor symmetry/breaking approach. We will first obtain the decay amplitude relations of the $D\to A\ell^{+}\nu_{\ell}$ or the vertex coefficient relations of $A\to VP$ by using the SU(3) flavor symmetry/breaking approach. Then, after ignoring the SU(3) flavor breaking effects due to poor experimental data of $D\to A\ell^{+}\nu_{\ell}$, $A\to VP$ and $D\to A(A\to VP)\ell^{+}\nu_{\ell}$, the relevant nonperturbative coefficients and the mixing angles will be constrained by the present experiential data. Finally, the branching ratios, which have not been measured yet, will be predicted by the SU(3) flavor symmetry. In addition, the sensitivities of $\mathcal{B}(D\to A\ell^{+}\nu_{\ell})$ and $\mathcal{B}(A\to VP)$ decays to the nonperturbative parameters are also displayed.
This paper will be organized as follows. In Sec. II, the results of the semileptonic $D\to A\ell^{+}\nu_{\ell}$ decays will be obtained. In Sec. III, the results of the nonleptonic $A\to VP$ decays will be calculated. The branching ratios of semileptonic four-body $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays will be discussed in Sec. IV. We summarize our results in Sec. V. Finally, all figures and tables are listed in Appendix.
II Semileptonic $D\rightarrow A\ell^{+}\nu_{\ell}$ decays
II.1 Meson multiplets
The SU(3) flavor symmetry analysis can provide the hadronic helicity amplitude relationships among various decay models. We will start by presenting the meson multiplets of the SU(3) flavor group. Charmed pseudoscalar mesons are composed of a charm quark and a light anti-quark
$D^{i}=\Big{(}D^{0}(c\bar{u}),\leavevmode\nobreak\ D^{+}(c\bar{d}),\leavevmode\nobreak\ D^{+}_{s}(c\bar{s})\Big{)}.$ (1)
Light pseudoscalar mesons and vector mesons are represented as He:2018joe
$P$ $=$ $\left(\begin{array}[]{ccc}\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta_{8}}{\sqrt{6}}+\frac{\eta_{1}}{\sqrt{3}}&\pi^{+}&K^{+}\\ \pi^{-}&-\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta_{8}}{\sqrt{6}}+\frac{\eta_{1}}{\sqrt{3}}&K^{0}\\ K^{-}&\overline{K}^{0}&-\frac{2\eta_{8}}{\sqrt{6}}+\frac{\eta_{1}}{\sqrt{3}}\end{array}\right)\,,$ (5)
$V$ $=$ $\left(\begin{array}[]{ccc}\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}&\rho^{+}&K^{*+}\\ \rho^{-}&-\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}&K^{*0}\\ K^{*-}&\overline{K}^{*0}&\phi\end{array}\right)\,.$ (9)
The pseudoscalar mesons $\eta$ and $\eta^{\prime}$ are the mixtures of $\eta_{1}=\frac{u\bar{u}+d\bar{d}+s\bar{s}}{\sqrt{3}}$ and $\eta_{8}=\frac{u\bar{u}+d\bar{d}-2s\bar{s}}{\sqrt{6}}$ with a mixing angle denoted as $\theta_{P}$ Bramon:1992kr
$\left(\begin{array}[]{c}\eta\\ \eta^{\prime}\end{array}\right)\,=\left(\begin{array}[]{cc}\mbox{cos}\theta_{P}&-\mbox{sin}\theta_{P}\\ \mbox{sin}\theta_{P}&\mbox{cos}\theta_{P}\end{array}\right)\,\left(\begin{array}[]{c}\eta_{8}\\ \eta_{1}\end{array}\right)\,.$ (16)
In our numerical analysis, we will utilize the mixing angle range of $\theta_{P}=[-20^{\circ},-10^{\circ}]$ from the Particle Data Group (PDG) PDG2022 .
The multiplets of the light axial-vector mesons are presented in the matrix of $A$ and $B$, where $A$ represents the axial-vector mesons with $J^{PC}=1^{++}$ and $B$ corresponds to the mesons with $J^{PC}=1^{+-}$ Roca:2003uk
$A$ $=$ $\left(\begin{array}[]{ccc}\frac{a^{0}_{1}}{\sqrt{2}}+\frac{f_{1}}{\sqrt{3}}+\frac{f_{8}}{\sqrt{6}}&a^{+}_{1}&K^{+}_{1A}\\ a^{-}_{1}&-\frac{a^{0}_{1}}{\sqrt{2}}+\frac{f_{1}}{\sqrt{3}}+\frac{f_{8}}{\sqrt{6}}&K^{0}_{1A}\\ K^{-}_{1A}&\overline{K}^{0}_{1A}&\frac{f_{1}}{\sqrt{3}}-\frac{2f_{8}}{\sqrt{6}}\end{array}\right)\,,$ (20)
$B$ $=$ $\left(\begin{array}[]{ccc}\frac{b^{0}_{1}}{\sqrt{2}}+\frac{h_{1}}{\sqrt{3}}+\frac{h_{8}}{\sqrt{6}}&b^{+}_{1}&K^{+}_{1B}\\ b^{-}_{1}&-\frac{b^{0}_{1}}{\sqrt{2}}+\frac{h_{1}}{\sqrt{3}}+\frac{h_{8}}{\sqrt{6}}&K^{0}_{1B}\\ K^{-}_{1B}&\overline{K}^{0}_{1B}&\frac{h_{1}}{\sqrt{3}}-\frac{2h_{8}}{\sqrt{6}}\end{array}\right)\,.$ (24)
Mesons $K_{1}(1270)$ and $K_{1}(1400)$ are the mixtures of $K_{1A}$ and $K_{1B}$ by the mixing angle $\theta_{K_{1}}$
$\left(\begin{array}[]{c}K_{1}(1270)\\ K_{1}(1400)\end{array}\right)\,=\left(\begin{array}[]{cc}sin\theta_{K_{1}}&cos\theta_{K_{1}}\\ cos\theta_{K_{1}}&-sin\theta_{K_{1}}\end{array}\right)\,\left(\begin{array}[]{c}K_{1A}\\ K_{1B}\end{array}\right).$ (31)
Mesons $f_{1}(1285)$, $f_{1}(1420)$ ($h_{1}(1170)$, $h_{1}(1415)$) can be expressed by the mixing angle $\theta_{3P_{1}}$ ($\theta_{1P_{1}}$)
$\left(\begin{array}[]{c}f_{1}(1285)\\ f_{1}(1420)\end{array}\right)\,=\left(\begin{array}[]{cc}cos\theta_{3P_{1}}&sin\theta_{3P_{1}}\\ -sin\theta_{3P_{1}}&cos\theta_{3P_{1}}\end{array}\right)\,\left(\begin{array}[]{c}f_{1}\\ f_{8}\end{array}\right),$ (38)
$\left(\begin{array}[]{c}h_{1}(1170)\\ h_{1}(1415)\end{array}\right)\,=\left(\begin{array}[]{cc}cos\theta_{1P_{1}}&sin\theta_{1P_{1}}\\ -sin\theta_{1P_{1}}&cos\theta_{1P_{1}}\end{array}\right)\,\left(\begin{array}[]{c}h_{1}\\ h_{8}\end{array}\right).$ (45)
The value of mixing angle $\theta_{K_{1}}$ has been investigated in various studies, for examples, $33^{\circ}$ and $57^{\circ}$ in Ref. Suzuki:1993yc , $37^{\circ}$ or $58^{\circ}$ Cheng:2003bn , [$35^{\circ},55^{\circ}$] Burakovsky:1997dd , $60^{\circ}$ Tayduganov:2011ui , $(33.6\pm 4.3)^{\circ}$ Divotgey:2013jba , etc. These ranges will not be used in later numerical analysis. We will obtain the ranges of $\theta_{K_{1}}$, $\theta_{1P_{1}}$ and $\theta_{3P_{1}}$ from the $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays.
II.2 Branching ratios
For the semileptonic decays of the $c\rightarrow q\ell^{+}\nu_{\ell}$ transition, the effective Hamiltonian can be expressed as
$\mathcal{H}_{eff}(c\rightarrow q\ell^{+}\nu_{\ell})$ $=$ $\frac{G_{F}}{\sqrt{2}}V_{cq}^{*}\bar{q}\gamma^{\mu}(1-\gamma_{5})c\leavevmode\nobreak\ \bar{\nu}_{\ell}\gamma_{\mu}(1-\gamma_{5})\ell,$ (46)
where $G_{F}$ is the Fermi constant, $V_{cq}$ is the CKM matrix element with $q=d,s$. The decay amplitudes of the $D\to A\ell^{+}\nu_{\ell}$ decays can be written as Wang:2022yyn
$\mathcal{M}(D\rightarrow A\ell^{+}\nu_{\ell})$ $=$ $\frac{G_{F}}{\sqrt{2}}\sum_{m,n}g_{mn}L^{\lambda_{\ell}\lambda_{\nu}}_{m}H_{\lambda_{A}n}.$ (47)
Lepton amplitudes $L^{\lambda_{\ell}\lambda_{\nu}}_{m}$ and hadron amplitudes $H_{\lambda_{A}n}$ are represented as
$L^{\lambda_{\ell}\lambda_{\nu}}_{m}$ $=$ $\epsilon_{\alpha}(m)\bar{\nu_{\ell}}\gamma^{\alpha}(1-\gamma_{5})\ell,$ (48)
$H_{\lambda_{A}n}$ $=$ $V_{cq}^{*}\epsilon_{\beta}^{*}(n)\langle{A}(p,\varepsilon^{*})|\bar{q}\gamma^{\beta}(1-\gamma_{5})c|D(p_{D})\rangle,$ (49)
here $\lambda_{A}=0,\pm 1$ , $\lambda_{\ell}=\pm\frac{1}{2}$ and $\lambda_{\nu}=+\frac{1}{2}$ are the particle helicities, and $\epsilon(m)$ with $m=0,t,\pm 1$ and $\varepsilon^{*}$ are the polarization vectors of the virtual $W$ and axial-vector mesons $A$, respectively.
The hadron amplitudes can be parameterized by form factors of the $D\to A$ transition Cheng:2003sm
$A(p,\varepsilon^{*})\left|q\gamma_{\mu}(1-\gamma_{5})c\right|D(p_{D})\right>$ $=$ $\frac{2iA(q^{2})}{m_{D}-m_{A}}\epsilon_{\mu\nu\alpha\beta}\varepsilon^{*\nu}p^{\alpha}_{D}p^{\beta}$ (50)
$-\left[\varepsilon^{*}_{\mu}(m_{D}-m_{A})V_{1}(q^{2})-(p_{D}+p)_{\mu}(\varepsilon^{*}.p_{D})\frac{V_{2}(q^{2})}{m_{D}-m_{A}}\right]$
$+q_{\mu}(\varepsilon^{*}.p_{D})\frac{2m_{A}}{q^{2}}[V_{3}(q^{2})-V_{0}(q^{2})],$
where $q^{2}\equiv(p_{D}-p)^{2}$, $A(q^{2})$ and $V_{0,1,2,3}(q^{2})$ are the form factors of the $D\to A$ transition, and $m_{A,D,\ell}$ are the masses of $A,D,\ell$. Then hadronic amplitudes can be written as Sun:2011ssd ; Li:2009tx ; Colangelo:2019axi
$H_{\pm}$ $=$ $(m_{D}-m_{A})V_{1}(q^{2})\mp\frac{2m_{D}|\vec{p}_{A}|}{(m_{D}-m_{A})}A(q^{2}),$ (51)
$H_{0}$ $=$ $\frac{1}{2m_{A}\sqrt{q^{2}}}\left[(m_{D}^{2}-m_{A}^{2}-q^{2})(m_{D}-m_{A})V_{1}(q^{2})-\frac{4m_{D}^{2}|\vec{p}_{A}|^{2}}{m_{D}-m_{A}}V_{2}(q^{2})\right],$ (52)
$H_{t}$ $=$ $\frac{2m_{D}|\vec{p}_{A}|}{\sqrt{q^{2}}}V_{0}(q^{2}),$ (53)
in which $|\vec{p}_{A}|\equiv\sqrt{\lambda(m_{D}^{2},m_{A}^{2},q^{2})}/(2m_{D})$ with $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2ab-2ac-2bc$.
The differential branching ratios are Ivanov:2019nqd
$\frac{d\mathcal{B}(D\to A\ell^{+}\nu_{\ell})}{dq^{2}}=\frac{\tau_{D}G_{F}^{2}|V_{cq}|^{2}\lambda^{1/2}(q^{2}-m_{\ell}^{2})^{2}}{24(2\pi)^{3}m_{D}^{3}q^{2}}{\cal H}_{\rm total},$ (54)
with ${\cal H}_{\rm total}\equiv({\cal H}_{U}+{\cal H}_{L})\left(1+\frac{m_{\ell}^{2}}{2q^{2}}\right)+\frac{3m_{\ell}^{2}}{2q^{2}}{\cal H}_{S},$ (55)
and ${\cal H}_{U}=|H_{+}|^{2}+|H_{-}|^{2},\quad{\cal H}_{L}=|H_{0}|^{2},\quad{\cal H}_{P}=|H_{+}|^{2}-|H_{-}|^{2},\quad{\cal H}_{S}=|H_{t}|^{2},\quad{\cal H}_{SL}=\Re(H_{0}H_{t}^{{\dagger}}),$ (56)
where $\lambda\equiv\lambda(m_{D}^{2},m_{A}^{2},q^{2})$, $m_{\ell}^{2}\leq q^{2}\leq(m_{D}-m_{A})^{2}$, and $\tau_{D}$ is the mean life of $D$ meson. Usually, the branching ratios can be obtained by Eq. (54) and the form factors, nevertheless, the results of the form factors are dependent on different approaches.
Branching ratios also can be obtained by using the available data and the SU(3) flavor symmetry/breaking, which is independent of the detailed dynamics, offering us an opportunity to relate different decay modes. In the $\ell=e,\mu$ cases, $m_{\ell}^{2}$ is small, and it can be ignored in ${\cal H}_{\rm total}$ in Eq. (55). Then ${\cal H}_{\rm total}$ only contains the hadronic information, which can be related by the the SU(3) flavor symmetry/breaking.
II.3 Hadronic helicity amplitude relations by SU(3) flavor symmetry/breaking
Similar to the SU(3) flavor symmetry/breaking hadronic helicity amplitudes of $D\to P/V/S\ell^{+}\nu_{\ell}$ decays given in Ref. Wang:2022yyn , the hadronic helicity amplitudes of the $D\to A\ell^{+}\nu_{\ell}$ decays can be parameterized as
$H(D\to M\ell^{+}\nu_{\ell})=c_{0}^{M}D_{i}M^{i}_{j}H^{j}+c_{1}^{M}D_{k}W_{i}^{k}M^{i}_{j}H^{j}+c_{2}^{M}D_{i}M_{k}^{i}W_{j}^{k}H^{j},$ (57)
where $H^{2}\equiv V^{*}_{cd}$ and $H^{3}\equiv V^{*}_{cs}$ are the CKM matrix elements, $W$ is the SU(3) flavor breaking related matrix Xu:2013dta ; He:2014xha , $c_{0,1,2}^{A}$ and $c_{0,1,2}^{B}$ are the non-perturbation coefficients corresponding to $M=A$ and $M=B$, respectively, $c_{0}^{A,B}$ are the nonperturbative coefficients under the SU(3) flavor symmetry, and $c_{1,2}^{A,B}$ are the nonperturbative SU(3) flavor breaking coefficients. Noted that the Okubo-Zweig-Iizuka suppressed processes, for an example, $D^{+}_{s}\to b_{1}(1235)^{0}\ell^{+}\nu_{\ell}$ decays studied in Ref. BESIII:2023clm , are not considered in this work. In terms of Eq. (57), the hadronic helicity amplitude relations for the $D\to A\ell^{+}\nu_{\ell}$ decays are summarized in Tab. 1 of Appendix.
From Tab. 1, one can easily obtain the amplitude relations between different decay processes. The CKM matrix elements $V_{cs}$ and $V_{cd}$ are also listed in Tab. 1 for comparing conveniently. Since the light axial-vector matrix is distinguished as $A$ and $B$ in the meson multiplets, there are eight nonperturbative parameters, $A_{1}$, $A_{2}$, $A_{3}$, $A_{4}$ and $B_{1}$, $B_{2}$, $B_{3}$, $B_{4}$ as defined in the caption of Tab. 1. If neglecting the SU(3) flavor breaking effects of $c^{A}_{1}$, $c^{A}_{2}$, $c^{B}_{1}$ and $c^{B}_{2}$, there are only two nonperturbative parameters $c^{A}_{0}$ and $c^{B}_{0}$. The hadronic amplitudes of various decay processes are connected to the nonperturbative parameters $c^{A}_{0}$ and $c^{B}_{0}$, as well as the mixing angles $\theta_{K_{1}}$, $\theta_{1P_{1}}$ and $\theta_{3P_{1}}$.
II.4 Numerical results for the $D\to A\ell^{+}\nu_{\ell}$ decays
In our numerical analysis, the theoretical input parameters and experimental data within $1\sigma$ error and within $2\sigma$ errors are taken from the PDG PDG2022 . In the $D\to A\ell^{+}\nu_{\ell}$ decays, only $D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e}$ and $D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e}$ decays have been measured, and the branching ratio of $D^{+}_{s}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e}$ has been upper limited. Their experimental data with $1\sigma$ error are PDG2022 ; BESIII:2023clm
$\mathcal{B}(D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e})$ $=$ $(1.01\pm 0.18)\times 10^{-3},$ (58)
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e})$ $=$ $(2.30^{+0.40}_{-0.42})\times 10^{-3},$ (59)
[…]
Only two experimental data can not constrain $c^{A}_{0}$ and $c^{B}_{0}$, $\theta_{K_{1}}$, $\theta_{1P_{1}}$ and $\theta_{3P_{1}}$ five parameters. Three mixing angles $\theta_{K_{1}}$, $\theta_{1P_{1}}$ and $\theta_{3_{P_{1}}}$ can also be constrained by the $A\to PV$ non-leptonic decays, which will be discussed in the next section. $\mathcal{B}(D\to a^{-}_{1}/a^{0}_{1}\ell^{+}\nu_{\ell})$ and $\mathcal{B}(D\to b^{-}_{1}/b^{0}_{1}\ell^{+}\nu_{\ell})$ are proportional to $|c^{A}_{0}|^{2}$ and $|c^{B}_{0}|^{2}$, respectively. As given in Refs. Momeni:2019uag ; Momeni:2022gqb , the theoretical predictions of $\mathcal{B}(D\to a^{-}_{1}/a^{0}_{1}\ell^{+}\nu_{l})$ and $\mathcal{B}(D\to b^{-}_{1}/b^{0}_{1}\ell^{+}\nu_{l})$ are quite depended on different approaches. For the sake of conservatism, we only use $\Big{|}\frac{c^{A}_{0}}{c^{B}_{0}}\Big{|}=\sqrt{\frac{\mathcal{B}(D\to a^{-}_{1}/a^{0}_{1}\ell^{+}\nu_{l})}{\mathcal{B}(D\to b^{-}_{1}/b^{0}_{1}\ell^{+}\nu_{l})}}$ that lies in $[0.88,1.97]$ from Refs. Momeni:2019uag ; Momeni:2022gqb .
Now using the experimental data of $\mathcal{B}(D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e})$, $\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e})$ and $\mathcal{B}(D^{+}_{s}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e})$ in Eqs. (58-60), the ratio $\Big{|}\frac{c^{A}_{0}}{c^{B}_{0}}\Big{|}\in[0.88,1.97]$ from Refs. Momeni:2019uag ; Momeni:2022gqb , and the relevant constraints from later $\mathcal{B}(A\to PV)$ in Eqs. (66-67) and $\mathcal{B}(D\to Ae^{+}\nu_{e},A\to PV)$ in Eqs. (73-74), in terms of the hadronic amplitude relations listed in Tab. 1, one can constrain relevant parameters and then give the predictions of $\mathcal{B}(D\to A\ell^{+}\nu_{\ell})$, which have not been measured yet.
We find that the range of $\Big{|}\frac{c^{A}_{0}}{c^{B}_{0}}\Big{|}$ are not reduced by present experimental data. As for three mixing angles, their constraints mainly come from the $A\to PV$ decays, and everyone of them are constrained from $[-180^{\circ},180^{\circ}]$ to two allowed ranges as shown in Fig. 1 of Appendix. An angle in one allowed range can be found a supplementary angle in another allowed range, and the branching ratios have similar sensitiveness to the angles of two ranges. So we only discuss one of two ranges, which are about larger than $0^{\circ}$, and the allowed ranges are
$c^{A}_{0}\in[1.44,3.00],\leavevmode\nobreak\ c^{B}_{0}\in[0.98,2.27],\leavevmode\nobreak\ \theta_{K_{1}}\in[26^{\circ},70^{\circ}],\leavevmode\nobreak\ \theta_{1P_{1}}\in[-15^{\circ},75^{\circ}],\leavevmode\nobreak\ \theta_{3P_{1}}\in[32^{\circ},149^{\circ}]\leavevmode\nobreak\ \mbox{within}\leavevmode\nobreak\ 2\sigma\leavevmode\nobreak\ \mbox{errors},$ (61)
$c^{A}_{0}\in[1.69,2.42],\leavevmode\nobreak\ c^{B}_{0}\in[1.12,2.08],\leavevmode\nobreak\ \theta_{K_{1}}\in[52^{\circ},65^{\circ}],\leavevmode\nobreak\ \theta_{1P_{1}}\in[5^{\circ},56^{\circ}],\leavevmode\nobreak\ \theta_{3P_{1}}\in[56^{\circ},125^{\circ}]\leavevmode\nobreak\ \mbox{within}\leavevmode\nobreak\ 1\sigma\leavevmode\nobreak\ \mbox{error}.\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$ (62)
As previously mentioned, there are several previous estimations on the mixing angle $\theta_{K_{1}}$. The allowed $\theta_{K_{1}}$ is $[35^{\circ},55^{\circ}]$ in terms of two parameters: the mass difference of the $a_{1}(1260)$ and $b_{1}(1235)$ mesons and the ratio of the constituent quark masses Burakovsky:1997dd , $33^{\circ}$ and $57^{\circ}$ by $\tau$ decay Suzuki:1993yc , $(34\pm 13)^{\circ}$ by $B\to K_{1}\gamma$ Hatanaka:2008xj , $33^{\circ}$ is much more favored than $57^{\circ}$ in Ref. Cheng:2013cwa , etc. The allowed ranges of $\theta_{1P_{1}}$ and $\theta_{3P_{1}}$ are obtain, $\theta_{1P_{1}}=(23.3\pm 1)^{\circ}$ or $(38.3\pm 1)^{\circ}$ Cheng:2011pb , $\theta_{3P_{1}}=(4.3\pm 2)^{\circ}$ or $(66.3\pm 2)^{\circ}$ Cheng:2011pb , $\theta_{3_{P_{1}}}=\big{(}19.4^{+4.5}_{-4.6}\big{)}^{\circ}$ or $\big{(}51.1^{+4.5}_{-4.6}\big{)}^{\circ}$ Yang:2010ah . Our bounds within $1\sigma$ error bar favor the large values of $\theta_{K_{1}}$ and $\theta_{3_{P_{1}}}$ in Refs. Cheng:2011pb ; Yang:2010ah .
[…]
In addition, we show the sensitivities of the branching ratios to the nonperturbative parameters $A$ and $B$ as well as the mixing angles $\theta_{K_{1}}$, $\theta_{3P_{1}}$ and $\theta_{1P_{1}}$ in Figs. 2-3. We give the remarks as follow.
  • As given in Tab. 1, the hadronic amplitudes of the $D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e}$, $D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e}$ and $D^{+}_{s}\to{K}_{1}(1270)^{0}e^{+}\nu_{e}$ decays are proportionate to $sin\theta_{K_{1}}c^{A}_{0}+cos\theta_{K_{1}}c^{B}_{0}$, and the hadronic amplitudes of the $D^{0}\to K_{1}(1400)^{-}e^{+}\nu_{e}$, $D^{+}\to\overline{K}_{1}(1400)^{0}e^{+}\nu_{e}$ and $D^{+}_{s}\to{K}_{1}(1400)^{0}e^{+}\nu_{e}$ decays are proportionate to $cos\theta_{K_{1}}c^{A}_{0}-sin\theta_{K_{1}}c^{B}_{0}$. So as shown in Fig. 2, three $D\to K_{1}(1270)e^{+}\nu_{e}$ decays or three $D\to{K}_{1}(1400)e^{+}\nu_{e}$ decays have the similar variation trend to the three parameters $c^{A}_{0}$, $c^{B}_{0}$ or $\theta_{K_{1}}$. From Fig. 2 ($a_{1,2,3}$) and ($b_{1,2,3}$), one can see the experimental lower limit of $\mathcal{B}(D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e})$ and the experimental upper limit of $\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e})$ give effective constraints on our predictions. Three $\mathcal{B}(D\to K_{1}(1270)e^{+}\nu_{e})$ in Fig. 2 ($a_{1,2,3}$), ($b_{1,2,3}$) and ($c_{1,2,3}$) have some sensitive to $A$ and $B$. As shown in Fig. 2 ($d_{1,2,3}$), ($e_{1,2,3}$) and ($f_{1,2,3}$), three $\mathcal{B}(D\to{K}_{1}(1400)e^{+}\nu_{e})$ are very sensitive to three parameters $A$, $B$ and $\theta_{K_{1}}$.
  • As displayed in Fig. 3 ($a_{1-4}$), $\mathcal{B}(D\to a_{1}(1260)e^{+}\nu_{e})$ and $D\to b_{1}(1235)e^{+}\nu_{e})$ are only related to $A$ and $B$, respectively, and they are very sensitive to $A$ or $B$. Fig. 3 ($b_{1-4}$) and ($c_{1-4}$) show the $\mathcal{B}(D_{(s)}\to f_{1}(1285)/f_{1}(1420)e^{+}\nu_{e})$ which are related to $A$ and $\theta_{3P_{1}}$. Four $\mathcal{B}(D_{(s)}\to f_{1}(1285)/f_{1}(1420)e^{+}\nu_{e})$ have similar sensitive to $A$, but the change trends of four branch ratios to $\theta_{3P_{1}}$ are different. Fig. 3 ($d_{1-4}$) and ($e_{1-4}$) show the $\mathcal{B}(D_{(s)}\to h_{1}(1170)/h_{1}(1415)e^{+}\nu_{e})$, which are related to $B$ and $\theta_{1P_{1}}$. $\mathcal{B}(D^{+}\to h_{1}(1170)e^{+}\nu_{e})$ and $\mathcal{B}(D^{+}_{s}\to h_{1}(1415)e^{+}\nu_{e})$ are sensitive to parameter $B$. $\mathcal{B}(D^{+}_{s}\to h_{1}(1170)e^{+}\nu_{e})$ and $\mathcal{B}(D^{+}\to h_{1}(1415)e^{+}\nu_{e})$ are sensitive to $\theta_{1P_{1}}$, and Fig. 3 ($d_{2}$) and ($e_{4}$) show that the branching ratios are first decreasing and then increasing with $\theta_{1P_{1}}$, there are minimum values when $\theta_{1P_{1}}$ is about $36^{\circ}$.
III Non-leptonic $A\to VP$ decays
III.1 Decay branching ratios of the $A\to VP$ decays
For the non-leptonic decays of axial-vector mesons $A\to VP$, the decay branching ratios can be written as Roca:2003uk ; Parui:2022zfm
$\mathcal{B}(A\to VP)=\frac{\tau_{A}|\lambda_{AVP}|^{2}}{2\pi m_{A}^{2}}q^{\prime}\left(1+\frac{2}{3}\frac{{q^{\prime}}^{2}}{m_{V}^{2}}\right).$ (63)
where $q^{\prime}\equiv\frac{1}{2m_{A}}\lambda^{1/2}(m_{A}^{2},m_{V}^{2},m_{P}^{2})$ is the momentum of the final state particles in the rest frame of the axial-vector meson, and $\lambda_{AVP}$ are the coefficients of the $AVP$ vertex. $\lambda_{AVP}$ can be obtained from the Lagrangians in Eq. (1) of Ref. Roca:2003uk or the SU(3) flavor symmetry/breaking, we will use the later to obtain $\lambda_{AVP}$ in this work.
There are little phase spaces for some axial-vector meson decays, such as $K_{1}(1270)\to\rho K/\omega K/\phi K/K^{*}\eta$, $K_{1}(1400)\to\phi K$, $a_{1}(1260)\to K^{*}\overline{K}/\overline{K}^{*}K$, $b_{1}(1235)\to K^{*}\overline{K}/\overline{K}^{*}K/\rho\eta$, $f_{1}(1285)\to K^{*}\overline{K}/\overline{K}^{*}K$, $h_{1}(1170)/h_{1}(1415)\to K^{*}\overline{K}/\overline{K}^{*}K/\omega\eta$. The mass distributions of the resonances have been considered in Refs. Wang:2022vga ; Roca:2021bxk , and the mass distributions of the resonances and vector mesons have been considered in Ref. Roca:2003uk . We will follow Ref. Roca:2003uk to obtain our numerical results. The decay branching ratios with the mass distributions of the resonances and vector mesons can be written as Roca:2003uk
[…]
×$\times\mathcal{B}_{AVP}(\sqrt{s_{A}},\sqrt{s_{V}})\Theta(\sqrt{s_{A}}-\sqrt{s_{V}}-m_{P})$
[…]
Now we related the $AVP$ vertex coefficients $\lambda_{AVP}$ for different decay modes by the SU(3) flavor symmetry/breaking. For $J^{PC}=1^{++}$ axial-vector mesons decays, the $AVP$ vertex coefficients can be parameterized as
$\lambda_{AVP}=c^{\prime A}_{0}A_{j}^{i}M^{j}_{k}M^{\prime k}_{i}+c^{\prime A}_{1}A_{j}^{a}W^{i}_{a}M^{j}_{k}M^{\prime k}_{i}+c^{\prime A}_{2}A_{j}^{i}M^{a}_{k}W^{j}_{a}M^{\prime k}_{i}+c^{\prime A}_{3}A_{j}^{i}M^{j}_{k}M^{\prime a}_{i}W^{k}_{a},$ (65)
where $M^{(^{\prime})}=V$ or $P$, $c^{\prime A}_{0}$ are the nonperturbative coefficients under the SU(3) flavor symmetry, and $c^{\prime A}_{1,2,3}$ are the nonperturbative SU(3) flavor breaking coefficients. The vertex coefficients for the $J^{PC}=1^{+-}$ axial-vector mesons decays can be obtained by replacing $A$ to $B$ in Eq. (65). To keep the invariant results under the conjugate change, $c^{\prime A/B}_{1}=c^{\prime A/B}_{2}$ are used in our analysis. In terms of Eq. (65), the vertex coefficient relations for the $A/B\to VP$ decays are summarized in Tabs. 3-5. In Tabs. 3-5, we redefine $F_{1,2,3,4,5}$ and $D_{1,2,3,4,5,6}$ by $c^{\prime A}_{0,1,3}$ and $c^{\prime B}_{0,1,3}$, respectively. If ignoring the SU(3) flavor breaking effects, then we have $F_{1}=F_{2}=F_{3}=F_{4}=F_{5}=F$ and $D_{1}=D_{2}=D_{3}=D_{4}=D_{5}=D_{6}=D$, where $F=c^{A}_{0}$ and $D=c^{B}_{0}$. After ignoring the SU(3) flavor breaking effects, most our vertex coefficients in Tabs. 3-5 are in agreement with ones in Ref. Roca:2003uk except some with different meson definitions such as $\eta,\eta^{\prime},h_{1},f_{1}$ mesons. Noted that there is overall minus sign difference for some vertex coefficients due to the different definition of $D$ and $F$, and it does not affect the branching ratio results.
III.2 Numerical results for the $A\to VP$ decays
In the $A\to VP$ decays, only a few branching ratios of $K_{1}(1270)/K_{1}(1400)\to VP$ have been measured PDG2022
$\mathcal{B}(K_{1}(1270)\to\rho K)=(38\pm 13)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(K_{1}(1270)\to\omega K)=(11\pm 2)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(K_{1}(1270)\to K^{*}\pi)=(21\pm 10)\%,$ (66)
$\mathcal{B}(K_{1}(1400)\to\omega K)=(1\pm 1)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(K_{1}(1400)\to\rho K)=(3\pm 3)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(K_{1}(1400)\to K^{*}\pi)=(94\pm 6)\%.$ (67)
And the experimental errors of $\mathcal{B}(K_{1}(1400)\to\rho K)$, $\mathcal{B}(K_{1}(1400)\to\omega K)$ and $\mathcal{B}(K_{1}(1270)\to K^{*}(892)\pi)$ are quite large. In consideration of no enough experimental data and consistency with our study of semi-leptonic decay $D\to A\ell^{+}\nu_{\ell}$, we only consider the SU(3) flavor symmetry for the $A\to VP$ numerical results.
Some values of $D$ and $F$ have been obtained from $A\to VP$ decays Roca:2003uk , considering the results in Ref. Roca:2003uk , we conservatively choose $F\in[0.6,2.0]$ and $D\in[-1.6,-0.4]$ as the original ranges. Then using the experimental data of $\mathcal{B}(K_{1}\to VP)$ in Eqs. (66-67), $\mathcal{B}(a_{1}(1260)/b_{1}(1235)/h_{1}(1170)/$ $h_{1}(1415)/f_{1}(1285)/f_{1}(1420)\to VP)\leq 1$, and also considering extra constraint from $\mathcal{B}(D\to K_{1}(1270)e^{+}\nu_{e})$ in Eqs. (58-60) and $\mathcal{B}(D\to Ae^{+}\nu_{e},A\to PV)$ in Eqs. (73-74), one can constrain relevant parameters $F$, $D$ and the mixing angles $\theta_{K_{1}}$, $\theta_{3P_{1}}$ and $\theta_{1P_{1}}$. Three constrained mixing angles are given in Eqs. (61-62), and the constrained $F$ and $D$ are
[…]
Using relevant constrained parameters, we give the predictions of $\mathcal{B}(A\to VP)$, which have not been measured yet. The numerical results are given in Tabs. 6-8. Most of our branching ratio predictions have large errors, since we consider the errors of the decay widths and the masses, conservatively choose the parameters $D$, $F$ as well as three mixing angles, and consider the present large experimental errors of $K_{1}(1270)/K_{1}(1400)\to VP$. If the result is shown as $a\pm a$, that means the allowed lower limit is so small, and only upper limit $2a$ is obtained. There is no enough phase space for $K_{1}(1270)\to\phi K$ decay, its amplitude is given in Tab. 3 in order to obtain the amplitude of $K_{1}(1400)\to\phi K$.
Previous fitted results have been obtained by three possible solutions in Ref. Roca:2003uk . Except measured decays of $K_{1}(1270)$ and $K_{1}(1400)$ mesons given in Eqs. (66-67), the average values of the three fitted central values are Roca:2003uk
$\mathcal{B}(a_{1}(1260)\to\rho\pi)=62.75\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(a_{1}(1260)\to K^{*}K)=7.53\%,$ (70)
$\mathcal{B}(b_{1}(1235)\to\omega\pi)=73.47\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(b_{1}(1235)\to K^{*}K)=5.92\%.$ (71)
Our prediction of $\mathcal{B}(b_{1}(1235)\to K^{*}K)$ is smaller than ones in Ref. Roca:2003uk , nevertheless, other three branching ratios are consistent with each other.
The sensitivities of $\mathcal{B}(A\to VP)$ to the constrained parameters are shown in Figs. 4-6. As shown in Figs. 4-5, the upper limits of $K_{1}(1270)\to K^{*}\pi,\rho K$ and $K_{1}(1400)\to\rho K$ as well as the lower limits of $K_{1}(1270)\to K^{*}\pi,\omega K$ and $K_{1}(1400)\to K^{*}\pi,\rho K,\omega K$ give the effective constraints on the relevant nonperturbative parameters. $\mathcal{B}(K_{1}(1270)\to K^{*}\pi)$ and $\mathcal{B}(K_{1}(1270)\to K^{*}\eta)$ are sensitive to $F$ and $\theta_{K1}$. $\mathcal{B}(K_{1}(1400)\to\rho K)$, $\mathcal{B}(K_{1}(1400)\to\omega K)$ and $\mathcal{B}(K_{1}(1400)\to K^{*}\eta)$ are sensitive to $D$ and $\theta_{K1}$. $\mathcal{B}(K_{1}(1400)\to\rho K)$ and $\mathcal{B}(K_{1}(1400)\to\omega K)$ have similar change trends, and the inflection point values appear when $D\approx-1.0$ and $\theta_{K_{1}}\approx 56^{\circ}$. As displayed in Fig. 6, many $\mathcal{B}(a_{1}/b_{1}/f_{1}/h_{1}\to VP)$ are very sensitive to $F$, $D$, $\theta_{3_{P1}}$ and $\theta_{1_{P1}}$. Any future measurement of these decays can obviously constrain on the relevant nonperturbative parameters.
IV $D\to A(A\to VP)\ell^{+}\nu$ decays
For the four-body semileptonic decays with resonances, in the case of the decay widths of the resonances are very narrow, the resonant decay branching ratios respect a simple factorization relation
$\mathcal{B}(D\to R\ell^{+}\nu_{\ell},R\to VP)=\mathcal{B}(D\to R\ell^{+}\nu_{\ell})\times\mathcal{B}(R\to VP),$ (72)
where $R$ denotes the resonance. And this relation is also a good approximation for wider resonances.
At present, only experimental upper limits of $\mathcal{B}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\omega\pi^{0})$ and $\mathcal{B}(D^{0}\to b_{1}(1235)^{-}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{-}\to\omega\pi^{-})$ have been reported PDG2022 ; BESIII:2020jan
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\omega\pi^{0})\leq 1.75\times 10^{-4},$ (73)
$\mathcal{B}(D^{0}\to b_{1}(1235)^{-}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{-}\to\omega\pi^{-})\leq 1.12\times 10^{-4}.$ (74)
Branching ratios of $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays can be obtained by using $\mathcal{B}(D\rightarrow A\ell^{+}\nu_{\ell})$ given in Sec. II and $\mathcal{B}(A\rightarrow VP)$ given in Sec. III as well as considering the further constraints in Eqs. (73-74), and our predictions are listed in Tabs. 9-11.
[…]
V Summary
Four-body semileptonic decays $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ have been studied by the SU(3) flavor symmetry/breaking approach. The branching ratios of $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ is obtained by ones of the $D\to A\ell^{+}\nu_{\ell}$ and $A\to VP$ decays. In terms of the SU(3) flavor symmetry/breaking, the decay amplitudes (vertex coefficients) of the $D\to A\ell^{+}\nu_{\ell}$ ($A\to VP$) decays have been related. Then the allowed spaces of relevant nonperturbative parameters have been obtained by using the present experimental data. Finally, the branching ratio predictions of the $D\to A\ell^{+}\nu_{\ell}$, $A\to VP$ and $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays have been obtained. In addition, the sensitivities of the branching ratios of $D\to A\ell^{+}\nu_{\ell}$ and $A\to VP$ decays to the nonperturbative parameters have been explored. Our main results can be summarized as follows.
  • Decays $D\to A\ell^{+}\nu_{\ell}$: $\mathcal{B}(D\to K_{1}(1270)\ell^{+}\nu_{\ell})$ and $\mathcal{B}(D_{s}\to f_{1}/h_{1}\ell^{+}\nu_{\ell})$ are on the order of $\mathcal{O}(10^{-4}-10^{-3})$, which might be measured in near future experiments. $\mathcal{B}(D\to K_{1}(1400)\ell^{+}\nu_{\ell})$ are strongly suppressed by the mixing angle $\theta_{K_{1}}$ and $\mathcal{B}(D\to a_{1}/b_{1}/f_{1}/h_{1}\ell^{+}\nu_{\ell})$ and $\mathcal{B}(D_{s}\to K_{1}(1270)/K_{1}(1400)\ell^{+}\nu_{\ell})$ are strongly suppressed by the CKM matrix element $V_{cd}$, and they are on the order of $\mathcal{O}(10^{-6}-10^{-5})$. And we have found some decay branching ratios are very sensitive to the nonperturbative parameters $A$ and $B$, or the mixing angles $\theta_{K_{1}}$, $\theta_{3P_{1}}$ or $\theta_{1P_{1}}$.
  • Decays $A\to VP$: The experimental data of the nonleptonic two-body $A\to VP$ decays give relatively strong constrained on the nonperturbative parameters and the mixing angles. Many branching ratios have been obtained with the large errors since our conservative choice of the parameters and the experimental errors are lager. We have found that $\mathcal{B}(K_{1}(1270)\to K^{*}\pi,K^{*}\eta)$ are sensitive to $F$ and $\theta_{K1}$. $\mathcal{B}(K_{1}(1400)\to\rho K,\omega K,K^{*}\eta)$ are sensitive to $D$ and $\theta_{K1}$, $\mathcal{B}(K_{1}(1400)\to\rho K,\omega K)$ have the inflection point values when $D\approx-1.0$ and $\theta_{K_{1}}\approx 56^{\circ}$, and many $\mathcal{B}(a_{1}/b_{1}/f_{1}/h_{1}\to VP)$ are very sensitive to $F$, $D$, $\theta_{3_{P1}}$ or $\theta_{1_{P1}}$. Any future measurement of these decays can obviously constrain on the relevant parameters.
  • Decays $D\to A(A\to VP)\ell^{+}\nu_{\ell}$: Our predictions of $\mathcal{B}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\omega\pi^{0})$ and $\mathcal{B}(D^{0}\to b_{1}(1235)^{-}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{-}\to\omega\pi^{-})$ are much smaller than their present experimental upper limits. We have found that many $\mathcal{B}(D\to A\ell^{+}\nu_{\ell},A\to VP)$ are on the order of $\mathcal{O}(10^{-5}-10^{-4})$, and they are expected to be measured in the near future.
Although SU(3) flavor symmetry is approximate, it can still provide very useful information about these decays. According to our rough predictions, some decays of $D\to A\ell^{+}\nu_{\ell}$, $A\to VP$, and $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ might be measured in the near future experiments, and they are useful to test and understand the nonperturbative effects in the semileptonic decays. Nevertheless, some decay branching ratios are really small and could not be measured recently.
Appendix
All figures and tables are listed in this Appendix.
Table 1: The hadronic amplitudes of the $D\to A\ell^{+}\nu_{\ell}$ decays. $A_{1}\equiv c^{A}_{0}+c^{A}_{1}-2c^{A}_{2}$, $A_{2}\equiv c^{A}_{0}-2c^{A}_{1}-2c^{A}_{2}$, $A_{3}\equiv c^{A}_{0}+c^{A}_{1}+c^{A}_{2}$, $A_{4}\equiv c^{A}_{0}-2c^{A}_{1}+c^{A}_{2}$. $B_{1}\equiv c^{B}_{0}+c^{B}_{1}-2c^{B}_{2}$, $B_{2}\equiv c^{B}_{0}-2c^{B}_{1}-2c^{B}_{2}$, $B_{3}\equiv c^{B}_{0}+c^{B}_{1}+c^{B}_{2}$, $B_{4}\equiv c^{B}_{0}-2c^{B}_{1}+c^{B}_{2}$. $A_{1}=A_{2}=A_{3}=A_{4}=c^{A}_{0}$ and $B_{1}=B_{2}=B_{3}=B_{4}=c^{B}_{0}$ if neglecting the SU(3) flavor breaking $c^{A}_{1}$, $c^{A}_{2}$ and $c^{B}_{1}$, $c^{B}_{2}$ terms.
    Decay modes SU(3) hadronic amplitudes     Decay modes SU(3) hadronic amplitudes
$D^{0}\to K_{1}(1270)^{-}\ell^{+}\nu_{\ell}$ $(sin\theta_{K_{1}}A_{1}+cos\theta_{K_{1}}B_{1})\leavevmode\nobreak\ V^{*}_{cs}$ $D^{0}\to K_{1}(1400)^{-}\ell^{+}\nu_{\ell}$ $(cos\theta_{K_{1}}A_{1}-sin\theta_{K_{1}}B_{1})\leavevmode\nobreak\ V^{*}_{cs}$
$D^{+}\to\overline{K}_{1}(1270)^{0}\ell^{+}\nu_{\ell}$ $(sin\theta_{K_{1}}A_{1}+cos\theta_{K_{1}}B_{1})\leavevmode\nobreak\ V^{*}_{cs}$ $D^{+}\to\overline{K}_{1}(1400)^{0}\ell^{+}\nu_{\ell}$ $(cos\theta_{K_{1}}A_{1}-sin\theta_{K_{1}}B_{1})\leavevmode\nobreak\ V^{*}_{cs}$
$D^{+}_{s}\to f_{1}(1285)\ell^{+}\nu_{\ell}$ $(\frac{1}{\sqrt{3}}cos\theta_{3P_{1}}-\sqrt{\frac{2}{3}}sin\theta_{3P_{1}})\leavevmode\nobreak\ A_{2}\leavevmode\nobreak\ V^{*}_{cs}$ $D^{+}_{s}\to f_{1}(1420)\ell^{+}\nu_{\ell}$ $(-\frac{1}{\sqrt{3}}sin\theta_{3P_{1}}-\sqrt{\frac{2}{3}}cos\theta_{3P_{1}})\leavevmode\nobreak\ A_{2}\leavevmode\nobreak\ V^{*}_{cs}$
$D^{+}_{s}\to h_{1}(1170)\ell^{+}\nu_{\ell}$ $(\frac{1}{\sqrt{3}}cos\theta_{1P_{1}}-\sqrt{\frac{2}{3}}sin\theta_{1P_{1}})\leavevmode\nobreak\ B_{2}\leavevmode\nobreak\ V^{*}_{cs}$ $D^{+}_{s}\to h_{1}(1415)\ell^{+}\nu_{\ell}$ $(-\frac{1}{\sqrt{3}}sin\theta_{1P_{1}}-\sqrt{\frac{2}{3}}cos\theta_{1P_{1}})\leavevmode\nobreak\ B_{2}\leavevmode\nobreak\ V^{*}_{cs}$
$D^{0}\to a_{1}(1260)^{-}\ell^{+}\nu_{\ell}$ $A_{3}\leavevmode\nobreak\ V^{*}_{cd}$ $D^{0}\to b_{1}(1235)^{-}\ell^{+}\nu_{\ell}$ $B_{3}\leavevmode\nobreak\ V^{*}_{cd}$
$D^{+}\to a_{1}(1260)^{0}\ell^{+}\nu_{\ell}$ $-\frac{1}{\sqrt{2}}\leavevmode\nobreak\ A_{3}\leavevmode\nobreak\ V^{*}_{cd}$ $D^{+}\to b_{1}(1235)^{0}\ell^{+}\nu_{\ell}$ $-\frac{1}{\sqrt{2}}\leavevmode\nobreak\ B_{3}\leavevmode\nobreak\ V^{*}_{cd}$
$D^{+}\to f_{1}(1285)\ell^{+}\nu_{\ell}$ $(\frac{1}{\sqrt{3}}cos\theta_{3P_{1}}+\frac{1}{\sqrt{6}}sin\theta_{3P_{1}})\leavevmode\nobreak\ A_{3}\leavevmode\nobreak\ V^{*}_{cd}$ $D^{+}\to f_{1}(1420)\ell^{+}\nu_{\ell}$ $(-\frac{1}{\sqrt{3}}sin\theta_{3P_{1}}+\frac{1}{\sqrt{6}}cos\theta_{3P_{1}})\leavevmode\nobreak\ A_{3}\leavevmode\nobreak\ V^{*}_{cd}$
$D^{+}\to h_{1}(1170)\ell^{+}\nu_{\ell}$ $(\frac{1}{\sqrt{3}}cos\theta_{1P_{1}}+\frac{1}{\sqrt{6}}sin\theta_{1P_{1}})\leavevmode\nobreak\ B_{3}\leavevmode\nobreak\ V^{*}_{cd}$ $D^{+}\to h_{1}(1415)\ell^{+}\nu_{\ell}$ $(-\frac{1}{\sqrt{3}}sin\theta_{1P_{1}}+\frac{1}{\sqrt{6}}cos\theta_{1P_{1}})\leavevmode\nobreak\ B_{3}\leavevmode\nobreak\ V^{*}_{cd}$
$D^{+}_{s}\to K_{1}(1270)^{0}\ell^{+}\nu_{\ell})$ $(sin\theta_{K_{1}}A_{4}+cos\theta_{K_{1}}B_{4})\leavevmode\nobreak\ V^{*}_{cd}$ $D^{+}_{s}\to K_{1}(1400)^{0}\ell^{+}\nu_{\ell})$ $(cos\theta_{K_{1}}A_{4}-sin\theta_{K_{1}}B_{4})\leavevmode\nobreak\ V^{*}_{cd}$
[…]
Table 2: Predictions of the branching ratios of the $D\to A\ell^{+}\nu_{\ell}$ decays.
Ones within $1\sigma$ error Ones within $2\sigma$ error Previous results
$\mathcal{B}(D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e})(\times 10^{-3})$ $0.93\pm 0.10$ $0.92\pm 0.27$cdots
$\mathcal{B}(D^{0}\to K_{1}(1400)^{-}e^{+}\nu_{e})(\times 10^{-3})$ $0.02\pm 0.02$ $0.08\pm 0.08$cdots
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e})(\times 10^{-3})$ $2.43\pm 0.27$ $2.39\pm 0.71$ $3.2\pm 0.40$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1400)^{0}e^{+}\nu_{e})(\times 10^{-3})$ $0.05\pm 0.05$ $0.20\pm 0.20$$\{0.005,0.02\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to f_{1}(1285)e^{+}\nu_{e})(\times 10^{-3})$ $0.86\pm 0.73$ $1.16\pm 1.16$$\{0.06,0.36\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to f_{1}(1420)e^{+}\nu_{e})(\times 10^{-3})$ $0.21\pm 0.21$ $0.39\pm 0.39$ $0.25\pm 0.05$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to h_{1}(1170)e^{+}\nu_{e})(\times 10^{-3})$ $0.26\pm 0.26$ $0.43\pm 0.43$$\{0,0.197\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to h_{1}(1415)e^{+}\nu_{e})(\times 10^{-3})$ $0.28\pm 0.16$ $0.31\pm 0.24$ $0.64\pm 0.07$ Cheng:2017pcq
$\mathcal{B}(D^{0}\to a_{1}(1260)^{-}e^{+}\nu_{e})(\times 10^{-5})$ $4.46\pm 2.32$ $7.72\pm 6.54$ $6.90$ Huang:2021owr , $5.261^{+0.745}_{-0.639}$ Hu:2021lkl
$\mathcal{B}(D^{0}\to b_{1}(1235)^{-}e^{+}\nu_{e})(\times 10^{-5})$ $2.61\pm 1.44$ $2.95\pm 2.06$ $4.85$ Huang:2021owr
$\mathcal{B}(D^{+}\to a_{1}(1260)^{0}e^{+}\nu_{e})(\times 10^{-5})$ $5.79\pm 3.00$ $9.95\pm 8.41$ $9.38$ Huang:2021owr , $6.673^{+0.947}_{-0.811}$ Hu:2021lkl
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e})(\times 10^{-5})$ $3.41\pm 1.88$ $3.84\pm 2.71$ $7.4\pm 0.70$ Cheng:2017pcq , $6.58$ Huang:2021owr
$\mathcal{B}(D^{+}\to f_{1}(1285)e^{+}\nu_{e})(\times 10^{-5})$ $1.88\pm 1.88$ $33.33\pm 3$ 3.33 $3.7\pm 0.80$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to f_{1}(1420)e^{+}\nu_{e})(\times 10^{-5})$ $0.64\pm 0.54$ $0.84\pm 0.84$$\{0.02,0.14\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to h_{1}(1170)e^{+}\nu_{e})(\times 10^{-5})$ $5.28\pm 3.00$ $5.59\pm 4.50$ $14\pm 1.50$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to h_{1}(1415)e^{+}\nu_{e})(\times 10^{-5})$ $0.11\pm 0.11$ $0.18\pm 0.18$$\{0,0.02\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to K_{1}(1270)^{0}e^{+}\nu_{e})(\times 10^{-5})$ $11.77\pm 1.39$ $11.70\pm 3.73$ $17\pm 2.00$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to K_{1}(1400)^{0}e^{+}\nu_{e})(\times 10^{-5})$ $0.32\pm 0.32$ $1.17\pm 1.17$$\{0.05,0.14\}$ Cheng:2017pcq
$\mathcal{B}(D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu})(\times 10^{-3})$ $0.81\pm 0.09$ $0.80\pm 0.24$cdots
$\mathcal{B}(D^{0}\to K_{1}(1400)^{-}\mu^{+}\nu_{\mu})(\times 10^{-3})$ $0.02\pm 0.02$ $0.06\pm 0.06$cdots
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu})(\times 10^{-3})$ $2.10\pm 0.23$ $2.07\pm 0.62$ $2.6\pm 0.30$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1400)^{0}\mu^{+}\nu_{\mu})(\times 10^{-3})$ $0.04\pm 0.04$ $0.16\pm 0.16$$\{0.004,0.017\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to f_{1}(1285)\mu^{+}\nu_{\mu})(\times 10^{-3})$ $0.76\pm 0.65$ $1.03\pm 1.03$$\{0.052,0.306\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to f_{1}(1420)\mu^{+}\nu_{\mu})(\times 10^{-3})$ $0.18\pm 0.18$ $0.32\pm 0.32$ $0.21\pm 0.05$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to h_{1}(1170)\mu^{+}\nu_{\mu})(\times 10^{-3})$ $0.24\pm 0.24$ $0.39\pm 0.39$$\{0,0.174\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to h_{1}(1415)\mu^{+}\nu_{\mu})(\times 10^{-3})$ $0.24\pm 0.14$ $0.26\pm 0.20$ $0.54\pm 0.06$ Cheng:2017pcq
$\mathcal{B}(D^{0}\to a_{1}(1260)^{-}\mu^{+}\nu_{\mu})(\times 10^{-5})$ $3.93\pm 2.08$ $6.91\pm 5.90$ $6.27$ Huang:2021owr , $4.732^{+0.685}_{-0.590}$ Hu:2021lkl
$\mathcal{B}(D^{0}\to b_{1}(1235)^{-}\mu^{+}\nu_{\mu})(\times 10^{-5})$ $2.29\pm 1.26$ $2.58\pm 1.81$ $4.40$ Huang:2021owr
$\mathcal{B}(D^{+}\to a_{1}(1260)^{0}\mu^{+}\nu_{\mu})(\times 10^{-5})$ $5.11\pm 2.70$ $8.92\pm 7.61$ $8.52$ Huang:2021owr , $6.002^{+0.796}_{-0.748}$ Hu:2021lkl
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}\mu^{+}\nu_{\mu})(\times 10^{-5})$ $2.99\pm 1.65$ $3.37\pm 2.38$ $6.4\pm 0.6$ Cheng:2017pcq , $6.00$ Huang:2021owr
$\mathcal{B}(D^{+}\to f_{1}(1285)\mu^{+}\nu_{\mu})(\times 10^{-5})$ $1.61\pm 1.61$ $2.84\pm 2.84$ $3.2\pm 0.6$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to f_{1}(1420)\mu^{+}\nu_{\mu})(\times 10^{-5})$ $0.48\pm 0.41$ $0.64\pm 0.64$$\{0.02,0.12\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to h_{1}(1170)\mu^{+}\nu_{\mu})(\times 10^{-5})$ $4.73\pm 2.69$ $5.01\pm 4.04$ $12.2\pm 1.3$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to h_{1}(1415)\mu^{+}\nu_{\mu})(\times 10^{-5})$ $0.08\pm 0.08$ $0.14\pm 0.14$$\{0,0.02\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to K_{1}(1270)^{0}\mu^{+}\nu_{\mu})(\times 10^{-5})$ $10.57\pm 1.25$ $10.51\pm 3.33$ $15\pm 2$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to K_{1}(1400)^{0}\mu^{+}\nu_{\mu})(\times 10^{-5})$ $0.27\pm 0.27$ $1.00\pm 1.00$$\{0.05,0.12\}$ Cheng:2017pcq
[…]
Figure 3: Branching ratios of the $D_{(s)}\to a_{1}/b_{1}/f_{1}/h_{1}e^{+}\nu_{e}$ decays.
Table 3: The vertex coefficients of $K_{1}(1270)\to VP$ decays. $F_{1}=c^{\prime A}_{0}-c^{\prime A}_{1}+c^{\prime A}_{3},\leavevmode\nobreak\ D_{1}=c^{\prime B}_{0}-c^{\prime B}_{1}+c^{\prime B}_{3}$, $F_{2}=c^{\prime A}_{0}-c^{\prime A}_{1}-2c^{\prime A}_{3},\leavevmode\nobreak\ D_{2}=c^{\prime B}_{0}-c^{\prime B}_{1}-2c^{\prime B}_{3}$. Ones of $K_{1}(1400)\to VP$ decays could be obtained from $K_{1}(1270)\to VP$ decays by replacing $sin\theta_{K_{1}}\to cos\theta_{K_{1}}$ and $cos\theta_{K_{1}}\to-sin\theta_{K_{1}}$.
   Decay modes Vertex coefficients $\lambda_{AVP}$
$K_{1}(1270)^{-}\to K^{*-}\pi^{0}$ $\frac{1}{\sqrt{2}}\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$K_{1}(1270)^{-}\to K^{*-}\eta$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}\big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\big{)}+\big{(}F_{2}sin\theta_{K_{1}}-D_{2}cos\theta_{K_{1}}\big{)}\big{(}\frac{2cos\theta_{P}}{\sqrt{6}}+\frac{sin\theta_{P}}{\sqrt{3}}\big{)}$
$K_{1}(1270)^{-}\to K^{*-}\eta^{\prime}$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}\big{(}\frac{sin\theta_{P}}{\sqrt{6}}+\frac{cos\theta_{P}}{\sqrt{3}}\big{)}+\big{(}F_{2}sin\theta_{K_{1}}-D_{2}cos\theta_{K_{1}}\big{)}\big{(}\frac{2sin\theta_{P}}{\sqrt{6}}-\frac{cos\theta_{P}}{\sqrt{3}}\big{)}$
$K_{1}(1270)^{-}\to\rho^{0}K^{-}$ $\frac{1}{\sqrt{2}}\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$K_{1}(1270)^{-}\to\omega K^{-}$ $\frac{1}{\sqrt{2}}\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$K_{1}(1270)^{-}\to K^{*0}\pi^{-}$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$K_{1}(1270)^{-}\to\rho^{-}\overline{K}^{0}$ $\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$K_{1}(1270)^{-}\to\phi K^{-}$ $\big{(}F_{2}sin\theta_{K_{1}}+D_{2}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to K^{*-}\pi^{+}$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\rho^{+}K^{-}$ $\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\pi^{0}$ $-\frac{1}{\sqrt{2}}\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\eta$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}\big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\big{)}+\big{(}F_{2}sin\theta_{K_{1}}-D_{2}cos\theta_{K_{1}}\big{)}\big{(}\frac{2cos\theta_{P}}{\sqrt{6}}+\frac{sin\theta_{P}}{\sqrt{3}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\eta^{\prime}$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}\big{(}\frac{sin\theta_{P}}{\sqrt{6}}+\frac{cos\theta_{P}}{\sqrt{3}}\big{)}+\big{(}F_{2}sin\theta_{K_{1}}-D_{2}cos\theta_{K_{1}}\big{)}\big{(}\frac{2sin\theta_{P}}{\sqrt{6}}-\frac{cos\theta_{P}}{\sqrt{3}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\rho^{0}\overline{K}^{0}$ $-\frac{1}{\sqrt{2}}\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\omega\overline{K}^{0}$ $\frac{1}{\sqrt{2}}\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\phi\overline{K}^{0}$ $\big{(}F_{2}sin\theta_{K_{1}}+D_{2}cos\theta_{K_{1}}\big{)}$
Table 4: The vertex coefficients of the $a_{1}(1260)/b_{1}(1235)\to VP$ decays. $F_{3}=c^{\prime A}_{0}+2c^{\prime A}_{1}+c^{\prime A}_{3}$, $F_{4}=c^{\prime A}_{0}+2c^{\prime A}_{1}-2c^{\prime A}_{3}$, $D_{3}=c^{\prime B}_{0}+2c^{\prime B}_{1}+c^{\prime B}_{3}$, $D_{4}=c^{\prime B}_{0}+2c^{\prime B}_{1}-2c^{\prime B}_{3}$.
Table 5: The vertex coefficients of the $h_{1}(1170)/f_{1}(1285)\to VP$ decays. $F_{5}=c^{\prime A}_{0}-4c^{\prime A}_{1}+c^{\prime A}_{3}$, $D_{5}=c^{\prime B}_{0}-4c^{\prime B}_{1}+c^{\prime B}_{3}$, $D_{6}=c^{\prime B}_{0}-4c^{\prime B}_{1}-2c^{\prime B}_{3}$. Ones of the $h_{1}(1415)\to VP$ decays could be obtained from the $h_{1}(1170)\to VP$ decays by replacing $sin\theta_{1P_{1}}\to cos\theta_{1P_{1}}$ and $cos\theta_{1P_{1}}\to-sin\theta_{1P_{1}}$. Ones of the $f_{1}(1420)\to VP$ decays could be obtained from the $f_{1}(1285)\to VP$ decays by replacing $sin\theta_{3P_{1}}\to cos\theta_{3P_{1}}$ and $cos\theta_{3P_{1}}\to-sin\theta_{3P_{1}}$.
   Decay modes Vertex coefficients $\lambda_{AVP}$    Decay modes Vertex coefficients $\lambda_{AVP}$
$a_{1}(1260)^{0}\to\rho^{+}\pi^{-}$ $\sqrt{2}F_{3}$ $b_{1}(1235)^{0}\to\rho^{0}\eta$ $2D_{3}\big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\big{)}$
$a_{1}(1260)^{0}\to\rho^{-}\pi^{+}$ $-\sqrt{2}F_{3}$ $b_{1}(1235)^{0}\to\rho^{0}\eta^{\prime}$ $2D_{3}\big{(}\frac{sin\theta_{P}}{\sqrt{6}}+\frac{cos\theta_{P}}{\sqrt{3}}\big{)}$
$a_{1}(1260)^{0}\to K^{*0}\overline{K}^{0}$ $-\frac{1}{\sqrt{2}}F_{4}$ $b_{1}(1235)^{0}\to K^{*0}\overline{K}^{0}$ $-\frac{1}{\sqrt{2}}D_{4}$
$a_{1}(1260)^{0}\to\overline{K}^{*0}K^{0}$ $\frac{1}{\sqrt{2}}F_{4}$ $b_{1}(1235)^{0}\to\overline{K}^{*0}K^{0}$ $-\frac{1}{\sqrt{2}}D_{4}$
$a_{1}(1260)^{0}\to K^{*+}K^{-}$ $\frac{1}{\sqrt{2}}F_{4}$ $b_{1}(1235)^{0}\to K^{*+}K^{-}$ $\frac{1}{\sqrt{2}}D_{4}$
$a_{1}(1260)^{0}\to K^{*-}K^{+}$ $-\frac{1}{\sqrt{2}}F_{4}$ $b_{1}(1235)^{0}\to K^{*-}K^{+}$ $\frac{1}{\sqrt{2}}D_{4}$
$b_{1}(1235)^{0}\to\omega\pi^{0}$ $\sqrt{2}D_{3}$
$a_{1}(1260)^{-}\to\rho^{-}\pi^{0}$ $\sqrt{2}F_{3}$ $b_{1}(1235)^{-}\to\rho^{-}\eta$ $2D_{3}\big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\big{)}$
$a_{1}(1260)^{-}\to\rho^{0}\pi^{-}$ $-\sqrt{2}F_{3}$ $b_{1}(1235)^{-}\to\rho^{-}\eta^{\prime}$ $2D_{3}\big{(}\frac{sin\theta_{P}}{\sqrt{6}}+\frac{cos\theta_{P}}{\sqrt{3}}\big{)}$
$a_{1}(1260)^{-}\to K^{*0}K^{-}$ $F_{4}$ $b_{1}(1235)^{-}\to K^{*0}K^{-}$ $D_{4}$
$a_{1}(1260)^{-}\to K^{*-}K^{0}$ $-F_{4}$ $b_{1}(1235)^{-}\to K^{*-}K^{0}$ $D_{4}$
$b_{1}(1235)^{-}\to\omega\pi^{-}$ $\sqrt{2}D_{3}$
   Decay modes Vertex coefficients $\lambda_{AVP}$
$h_{1}(1170)\to\rho^{0}\pi^{0}$ $\frac{2}{\sqrt{6}}D_{3}\big{(}\sqrt{2}cos\theta_{1P_{1}}+sin\theta_{1P_{1}}\big{)}$
$h_{1}(1170)\to\omega\eta$ $\frac{\sqrt{2}}{3}D_{3}\big{(}\sqrt{2}cos\theta_{1P_{1}}+sin\theta_{1P_{1}}\big{)}\big{(}cos\theta_{P}-\sqrt{2}sin\theta_{P}\big{)}$
$h_{1}(1170)\to\omega\eta^{\prime}$ $\frac{\sqrt{2}}{3}D_{3}\big{(}\sqrt{2}cos\theta_{1P_{1}}+sin\theta_{1P_{1}}\big{)}\big{(}sin\theta_{P}+\sqrt{2}cos\theta_{P}\big{)}$
$h_{1}(1170)\to\rho^{+}\pi^{-}$ $\frac{2}{\sqrt{6}}D_{3}\big{(}\sqrt{2}cos\theta_{1P_{1}}+sin\theta_{1P_{1}}\big{)}$
$h_{1}(1170)\to\rho^{-}\pi^{+}$ $\frac{2}{\sqrt{6}}D_{3}\big{(}\sqrt{2}cos\theta_{1P_{1}}+sin\theta_{1P_{1}}\big{)}$
$h_{1}(1170)\to K^{*+}K^{-}$ $\frac{cos\theta_{1P_{1}}}{\sqrt{3}}\big{(}D_{4}+D_{5}\big{)}+\frac{sin\theta_{1P_{1}}}{\sqrt{6}}\big{(}D_{4}-2D_{5}\big{)}$
$h_{1}(1170)\to K^{*-}K^{+}$ $\frac{cos\theta_{1P_{1}}}{\sqrt{3}}\big{(}D_{4}+D_{5}\big{)}+\frac{sin\theta_{1P_{1}}}{\sqrt{6}}\big{(}D_{4}-2D_{5}\big{)}$
$h_{1}(1170)\to K^{*0}\overline{K}^{0}$ $\frac{cos\theta_{1P_{1}}}{\sqrt{3}}\big{(}D_{4}+D_{5}\big{)}+\frac{sin\theta_{1P_{1}}}{\sqrt{6}}\big{(}D_{4}-2D_{5}\big{)}$
$h_{1}(1170)\to\overline{K}^{*0}K^{0}$ $\frac{cos\theta_{1P_{1}}}{\sqrt{3}}\big{(}D_{4}+D_{5}\big{)}+\frac{sin\theta_{1P_{1}}}{\sqrt{6}}\big{(}D_{4}-2D_{5}\big{)}$
$h_{1}(1170)\to\phi\eta$ $\frac{2}{3}D_{6}\big{(}-cos\theta_{1P_{1}}+\sqrt{2}sin\theta_{1P_{1}}\big{)}\big{(}\sqrt{2}cos\theta_{P}+sin\theta_{P}\big{)}$
$h_{1}(1170)\to\phi\eta^{\prime}$ $\frac{2}{3}D_{6}\big{(}-cos\theta_{1P_{1}}+\sqrt{2}sin\theta_{1P_{1}}\big{)}\big{(}\sqrt{2}sin\theta_{P}-cos\theta_{P}\big{)}$
$f_{1}(1285)\to K^{*+}K^{-}$ $\frac{cos\theta_{3P_{1}}}{\sqrt{3}}\big{(}F_{4}-F_{5}\big{)}+\frac{sin\theta_{3P_{1}}}{\sqrt{6}}\big{(}F_{4}+2F_{5}\big{)}$
$f_{1}(1285)\to K^{*-}K^{+}$ $-\frac{cos\theta_{3P_{1}}}{\sqrt{3}}\big{(}F_{4}-F_{5}\big{)}-\frac{sin\theta_{3P_{1}}}{\sqrt{6}}\big{(}F_{4}+2F_{5}\big{)}$
$f_{1}(1285)\to K^{*0}\overline{K}^{0}$ $\frac{cos\theta_{3P_{1}}}{\sqrt{3}}\big{(}F_{4}-F_{5}\big{)}+\frac{sin\theta_{3P_{1}}}{\sqrt{6}}\big{(}F_{4}+2F_{5}\big{)}$
$f_{1}(1285)\to\overline{K}^{*0}K^{0}$ $-\frac{cos\theta_{3P_{1}}}{\sqrt{3}}\big{(}F_{4}-F_{5}\big{)}-\frac{sin\theta_{3P_{1}}}{\sqrt{6}}\big{(}F_{4}+2F_{5}\big{)}$
Table 5: The vertex coefficients of the $h_{1}(1170)/f_{1}(1285)\to VP$ decays. $F_{5}=c^{\prime A}_{0}-4c^{\prime A}_{1}+c^{\prime A}_{3}$, $D_{5}=c^{\prime B}_{0}-4c^{\prime B}_{1}+c^{\prime B}_{3}$, $D_{6}=c^{\prime B}_{0}-4c^{\prime B}_{1}-2c^{\prime B}_{3}$. Ones of the $h_{1}(1415)\to VP$ decays could be obtained from the $h_{1}(1170)\to VP$ decays by replacing $sin\theta_{1P_{1}}\to cos\theta_{1P_{1}}$ and $cos\theta_{1P_{1}}\to-sin\theta_{1P_{1}}$. Ones of the $f_{1}(1420)\to VP$ decays could be obtained from the $f_{1}(1285)\to VP$ decays by replacing $sin\theta_{3P_{1}}\to cos\theta_{3P_{1}}$ and $cos\theta_{3P_{1}}\to-sin\theta_{3P_{1}}$.
Table 6: Predictions of the branching ratios of the $K_{1}\to VP$ decays by the SU(3) flavor symmetry. The unit is $10^{-2}$ for all branching ratios.
Branching ratios Predictions with $1\sigma$ error Predictions with $2\sigma$ errors
$\mathcal{B}(K_{1}(1270)^{-}\to K^{*-}\pi^{0})$ $7.24\pm 3.17$ $7.07\pm 6.73$
$\mathcal{B}(K_{1}(1270)^{-}\to K^{*-}\eta)$ $0.50\pm 0.16$ $0.71\pm 0.71$
$\mathcal{B}(K_{1}(1270)^{-}\to\rho^{0}K^{-})$ $16.82\pm 0.49$ $16.49\pm 5.29$
$\mathcal{B}(K_{1}(1270)^{-}\to\omega K^{-})$ $9.52\pm 0.21$ $9.55\pm 2.34$
$\mathcal{B}(K_{1}(1270)^{-}\to\overline{K}^{*0}\pi^{-})$ $14.20\pm 6.22$ $13.84\pm 13.18$
$\mathcal{B}(K_{1}(1270)^{-}\to\rho^{-}\overline{K}^{0})$ $32.71\pm 0.82$ $32.03\pm 10.08$
$\mathcal{B}(\overline{K}_{1}(1270)^{0}\to K^{*-}\pi^{+})$ $14.39\pm 6.31$ $14.03\pm 13.36$
$\mathcal{B}(\overline{K}_{1}(1270)^{0}\to\rho^{+}K^{-})$ $33.91\pm 0.86$ $33.27\pm 10.59$
$\mathcal{B}(\overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\pi^{0})$ $7.15\pm 3.13$ $6.96\pm 6.63$
$\mathcal{B}(\overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\eta)$ $0.41\pm 0.14$ $0.63\pm 0.63$
$\mathcal{B}(\overline{K}_{1}(1270)^{0}\to\rho^{0}\overline{K}^{0})$ $16.22\pm 0.47$ $15.86\pm 5.02$
$\mathcal{B}(\overline{K}_{1}(1270)^{0}\to\omega\overline{K}^{0})$ $8.87\pm 0.20$ $8.76\pm 2.19$
$\mathcal{B}(K_{1}(1270)\to K^{*}\pi)$ $21.48\pm 9.42$ $20.94\pm 19.93$
$\mathcal{B}(K_{1}(1270)\to\rho K)$ $49.78\pm 1.21$ $48.66\pm 15.32$
$\mathcal{B}(K_{1}(1270)\to\omega K)$ $9.20\pm 0.20$ $9.21\pm 2.21$
$\mathcal{B}(K_{1}(1270)\to K^{*}\eta)$ $0.45\pm 0.15$ $0.67\pm 0.67$
$\mathcal{B}(K_{1}(1400)^{-}\to K^{*-}\pi^{0})$ $30.54\pm 1.01$ $29.65\pm 2.15$
$\mathcal{B}(K_{1}(1400)^{-}\to K^{*-}\eta)$ $2.81\pm 1.84$ $3.29\pm 2.79$
$\mathcal{B}(K_{1}(1400)^{-}\to\rho^{0}K^{-})$ $0.75\pm 0.75$ $1.49\pm 1.49$
$\mathcal{B}(K_{1}(1400)^{-}\to\omega K^{-})$ $0.73\pm 0.73$ $1.44\pm 1.44$
$\mathcal{B}(K_{1}(1400)^{-}\to\overline{K}^{*0}\pi^{-})$ $60.32\pm 1.97$ $58.52\pm 4.22$
$\mathcal{B}(K_{1}(1400)^{-}\to\rho^{-}\overline{K}^{0})$ $1.48\pm 1.48$ $2.93\pm 2.93$
$\mathcal{B}(K_{1}(1400)^{-}\to\phi K^{-})$ $3.81\pm 0.65$ $3.61\pm 1.43$
$\mathcal{B}(\overline{K}_{1}(1400)^{0}\to K^{*-}\pi^{+})$ $60.88\pm 2.01$ $59.10\pm 4.29$
$\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\rho^{+}K^{-})$ $1.51\pm 1.51$ $2.98\pm 2.98$
$\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\overline{K}^{*0}\pi^{0})$ $30.27\pm 0.99$ $29.36\pm 2.12$
$\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\overline{K}^{*0}\eta)$ $2.67\pm 1.75$ $3.13\pm 2.66$
$\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\rho^{0}\overline{K}^{0})$ $0.74\pm 0.74$ $1.47\pm 1.47$
$\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\omega\overline{K}^{0})$ $0.71\pm 0.71$ $1.41\pm 1.41$
$\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\phi\overline{K}^{0})$ $3.64\pm 0.63$ $3.45\pm 1.39$
$\mathcal{B}(K_{1}(1400)\to K^{*}\pi)$ $91.01\pm 2.99$ $88.35\pm 6.35$
$\mathcal{B}(K_{1}(1400)\to\rho K)$ $2.24\pm 2.24$ $44.44\pm 4$ 4.44
$\mathcal{B}(K_{1}(1400)\to\omega K)$ $0.72\pm 0.72$ $1.42\pm 1.42$
$\mathcal{B}(K_{1}(1400)\to K^{*}\eta)$ $2.74\pm 1.79$ $3.21\pm 2.73$
$\mathcal{B}(K_{1}(1400)\to\phi K)$ $3.73\pm 0.64$ $3.53\pm 1.41$
Table 7: Predictions of the branching ratios of the $a_{1}/b_{1}\to VP$ decays by the SU(3) flavor symmetry. The unit is $10^{-2}$ for all branching ratios.
Branching ratios Predictions with $1\sigma$ error Predictions with $2\sigma$ error
$\mathcal{B}(a_{1}(1260)^{0}\to\rho^{+}\pi^{-})$ $38.15\pm 8.46$ $30.94\pm 17.66$
$\mathcal{B}(a_{1}(1260)^{0}\to\rho^{-}\pi^{+})$ $38.15\pm 8.46$ $30.94\pm 17.66$
$\mathcal{B}(a_{1}(1260)^{0}\to K^{*0}\overline{K}^{0})$ $1.82\pm 0.52$ $1.60\pm 1.15$
$\mathcal{B}(a_{1}(1260)^{0}\to\overline{K}^{*0}K^{0})$ $1.82\pm 0.52$ $1.60\pm 1.15$
$\mathcal{B}(a_{1}(1260)^{0}\to K^{*+}K^{-})$ $1.89\pm 0.53$ $1.67\pm 1.18$
$\mathcal{B}(a_{1}(1260)^{0}\to K^{*-}K^{+})$ $1.89\pm 0.53$ $1.67\pm 1.18$
$\mathcal{B}(a_{1}(1260)^{-}\to\rho^{-}\pi^{0})$ $38.33\pm 8.51$ $31.12\pm 17.78$
$\mathcal{B}(a_{1}(1260)^{-}\to\rho^{0}\pi^{-})$ $38.13\pm 8.45$ $30.94\pm 17.68$
$\mathcal{B}(a_{1}(1260)^{-}\to K^{*0}K^{-})$ $3.70\pm 1.04$ $3.26\pm 2.33$
$\mathcal{B}(a_{1}(1260)^{-}\to K^{*-}K^{0})$ $3.72\pm 1.04$ $3.28\pm 2.34$
$\mathcal{B}(a_{1}(1260)\to\rho\pi)$ $76.38\pm 16.93$ $61.97\pm 35.39$
$\mathcal{B}(a_{1}(1260)\to K^{*}K)$ $7.42\pm 2.09$ $6.53\pm 4.67$
$\mathcal{B}(b_{1}(1235)^{0}\to\rho^{0}\eta)$ $5.07\pm 2.14$ $4.91\pm 3.14$
$\mathcal{B}(b_{1}(1235)^{0}\to\omega\pi^{0})$ $65.46\pm 22.46$ $58.27\pm 34.31$
$\mathcal{B}(b_{1}(1235)^{0}\to K^{*0}\overline{K}^{0})$ $0.37\pm 0.15$ $0.40\pm 0.28$
$\mathcal{B}(b_{1}(1235)^{0}\to\overline{K}^{*0}K^{0})$ $0.37\pm 0.15$ $0.40\pm 0.28$
$\mathcal{B}(b_{1}(1235)^{0}\to K^{*+}K^{-})$ $0.42\pm 0.17$ $0.46\pm 0.32$
$\mathcal{B}(b_{1}(1235)^{0}\to K^{*-}K^{+})$ $0.42\pm 0.17$ $0.46\pm 0.32$
$\mathcal{B}(b_{1}(1235)^{-}\to\rho^{-}\eta)$ $5.13\pm 2.16$ $5.00\pm 3.21$
$\mathcal{B}(b_{1}(1235)^{-}\to\omega\pi^{-})$ $65.17\pm 22.36$ $58.00\pm 34.15$
$\mathcal{B}(b_{1}(1235)^{-}\to K^{*0}K^{-})$ $0.78\pm 0.32$ $0.85\pm 0.60$
$\mathcal{B}(b_{1}(1235)^{-}\to K^{*-}K^{0})$ $0.79\pm 0.33$ $0.86\pm 0.60$
$\mathcal{B}(b_{1}(1235)\to\rho\eta)$ $5.10\pm 2.15$ $4.96\pm 3.18$
$\mathcal{B}(b_{1}(1235)\to\omega\pi)$ $65.31\pm 22.41$ $58.14\pm 34.23$
$\mathcal{B}(b_{1}(1235)\to K^{*}K)$ $1.58\pm 0.65$ $1.71\pm 1.20$
Table 8: Predictions of the branching ratios of the $f_{1}/h_{1}\to VP$ decays by the SU(3) flavor symmetry. ${}^{\sharp}$$n=3$ is used for $\mathcal{B}(h_{1}(1415)\to\phi\eta)$. The unit is $10^{-2}$ for all branching ratios.
Branching ratios Predictions with $1\sigma$ error Predictions with $2\sigma$ error
$\mathcal{B}(f_{1}(1285)\to K^{*+}K^{-})$ $0.62\pm 0.24$ $0.52\pm 0.46$
$\mathcal{B}(f_{1}(1285)\to K^{*-}K^{+})$ $0.62\pm 0.24$ $0.52\pm 0.46$
$\mathcal{B}(f_{1}(1285)\to K^{*0}\overline{K}^{0})$ $0.13\pm 0.05$ $0.13\pm 0.11$
$\mathcal{B}(f_{1}(1285)\to\overline{K}^{*0}K^{0})$ $0.13\pm 0.05$ $0.13\pm 0.11$
$\mathcal{B}(f_{1}(1285)\to{K}K^{*})$ $1.49\pm 0.59$ $1.23\pm 1.07$
$\mathcal{B}(f_{1}(1420)\to K^{*+}K^{-})$ $13.09\pm 13.09$ $13.21\pm 13.21$
$\mathcal{B}(f_{1}(1420)\to K^{*-}K^{+})$ $13.09\pm 13.09$ $13.21\pm 13.21$
$\mathcal{B}(f_{1}(1420)\to K^{*0}\overline{K}^{0})$ $11.75\pm 11.75$ $11.84\pm 11.84$
$\mathcal{B}(f_{1}(1420)\to\overline{K}^{*0}K^{0})$ $11.75\pm 11.75$ $11.84\pm 11.84$
$\mathcal{B}(f_{1}(1420)\to K^{*}K)$ $49.64\pm 49.64$ $49.96\pm 49.96$
$\mathcal{B}(h_{1}(1170)\to\rho^{0}\pi^{0})$ $21.84\pm 9.33$ $18.70\pm 13.51$
$\mathcal{B}(h_{1}(1170)\to\omega\eta)$ $2.01\pm 1.04$ $1.99\pm 1.65$
$\mathcal{B}(h_{1}(1170)\to\rho^{+}\pi^{-})$ $21.74\pm 9.29$ $18.63\pm 13.47$
$\mathcal{B}(h_{1}(1170)\to\rho^{-}\pi^{+})$ $21.74\pm 9.29$ $18.63\pm 13.47$
$\mathcal{B}(h_{1}(1170)\to K^{*+}K^{-})$ $0.85\pm 0.76$ $1.10\pm 1.10$
$\mathcal{B}(h_{1}(1170)\to K^{*-}K^{+})$ $0.85\pm 0.76$ $1.10\pm 1.10$
$\mathcal{B}(h_{1}(1170)\to K^{*0}\overline{K}^{0})$ $0.81\pm 0.73$ $1.04\pm 1.04$
$\mathcal{B}(h_{1}(1170)\to\overline{K}^{*0}K^{0})$ $0.81\pm 0.73$ $1.04\pm 1.04$
$\mathcal{B}(h_{1}(1170)\to\phi\eta)$ $0.08\pm 0.08$ $0.13\pm 0.13$
$\mathcal{B}(h_{1}(1170)\to\rho\pi)$ $65.31\pm 27.91$ $55.97\pm 40.46$
$\mathcal{B}(h_{1}(1170)\to K^{*}K)$ $3.33\pm 2.98$ $4.29\pm 4.29$
$\mathcal{B}(h_{1}(1415)\to\rho^{0}\pi^{0})$ $13.20\pm 13.20$ $14.86\pm 14.86$
$\mathcal{B}(h_{1}(1415)\to\omega\eta)$ $3.37\pm 3.37$ $3.86\pm 3.86$
$\mathcal{B}(h_{1}(1415)\to\rho^{+}\pi^{-})$ $13.17\pm 13.17$ $14.81\pm 14.81$
$\mathcal{B}(h_{1}(1415)\to\rho^{-}\pi^{+})$ $13.17\pm 13.17$ $14.81\pm 14.81$
$\mathcal{B}(h_{1}(1415)\to K^{*+}K^{-})$ $12.15\pm 9.79$ $13.01\pm 12.96$
$\mathcal{B}(h_{1}(1415)\to K^{*-}K^{+})$ $12.15\pm 9.79$ $13.01\pm 12.96$
$\mathcal{B}(h_{1}(1415)\to K^{*0}\overline{K}^{0})$ $10.42\pm 8.50$ $11.80\pm 11.75$
$\mathcal{B}(h_{1}(1415)\to\overline{K}^{*0}K^{0})$ $10.42\pm 8.50$ $11.80\pm 11.75$
$\mathcal{B}(h_{1}(1415)\to\phi\eta)$ $1.19\pm 0.84^{\sharp}$ $1.34\pm 1.29^{\sharp}$
$\mathcal{B}(h_{1}(1415)\to\rho\pi)$ $39.54\pm 39.53$ $44.48\pm 44.48$
$\mathcal{B}(h_{1}(1415)\to K^{*}K)$ $45.13\pm 36.58$ $49.62\pm 49.42$
Figure 4: The constrained effects of the nonperturbative parameters in the $K_{1}(1270)\to VP$ decays. Branching ratios are in units of $10^{-2}$.
Figure 5: The constrained effects of the nonperturbative parameters in the $K_{1}(1400)\to VP$ decays. Branching ratios are in units of $10^{-2}$.
Figure 6: The constrained effects of the nonperturbative parameters in the $a_{1}/b_{1}/f_{1}/h_{1}\to VP$ decays. Branching ratios are in units of $10^{-2}$.
Table 9: Branching ratios of the $D\to VP\ell^{+}\nu_{\ell}$ decays with the $K_{1}$ resonances.
[…]
$\mathcal{B}(D^{0}\to K_{1}(1270)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{-}\to K^{*-}\pi^{0})(\times 10^{-5})$ $7.29\pm 3.48$ $8.09\pm 7.85$ $6.29\pm 3.00$ $7.04\pm 6.84$
$\mathcal{B}(D^{0}\to K_{1}(1270)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{-}\to\overline{K}^{*0}\pi^{-})(\times 10^{-4})$ $1.43\pm 0.68$ $1.59\pm 1.54$ $1.23\pm 0.59$ $1.38\pm 1.34$
$\mathcal{B}(D^{0}\to K_{1}(1270)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{-}\to\rho^{0}K^{-})(\times 10^{-4})$ $1.58\pm 0.20$ $1.67\pm 0.92$ $1.36\pm 0.17$ $1.45\pm 0.80$
$\mathcal{B}(D^{0}\to K_{1}(1270)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{-}\to\rho^{-}\overline{K}^{0})(\times 10^{-4})$ $3.06\pm 0.39$ $3.24\pm 1.75$ $2.64\pm 0.34$ $2.80\pm 1.52$
$\mathcal{B}(D^{0}\to K_{1}(1270)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{-}\to\omega K^{-})(\times 10^{-5})$ $8.93\pm 1.12$ $9.03\pm 4.26$ $7.70\pm 0.96$ $7.78\pm 3.65$
$\mathcal{B}(D^{0}\to K_{1}(1270)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{-}\to K^{*-}\eta)(\times 10^{-6})$ $4.79\pm 1.95$ $7.97\pm 7.93$ $4.13\pm 1.68$ $6.85\pm 6.82$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1270)^{0}\to K^{*-}\pi^{+})(\times 10^{-4})$ $3.77\pm 1.79$ $4.20\pm 4.08$ $3.26\pm 1.55$ $3.66\pm 3.56$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\pi^{0})(\times 10^{-4})$ $1.87\pm 0.89$ $2.09\pm 2.03$ $1.62\pm 0.77$ $1.82\pm 1.77$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1270)^{0}\to\rho^{+}K^{-})(\times 10^{-4})$ $8.25\pm 1.06$ $8.75\pm 4.74$ $7.13\pm 0.92$ $7.58\pm 4.12$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1270)^{0}\to\rho^{0}\overline{K}^{0})(\times 10^{-4})$ $3.95\pm 0.50$ $4.17\pm 2.25$ $3.42\pm 0.44$ $3.61\pm 1.96$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1270)^{0}\to\omega\overline{K}^{0})(\times 10^{-4})$ $2.17\pm 0.28$ $2.18\pm 1.04$ $1.87\pm 0.24$ $1.88\pm 0.89$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\eta)(\times 10^{-5})$ $1.03\pm 0.44$ $1.85\pm 1.84$ $0.89\pm 0.38$ $1.59\pm 1.59$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{0}\to K^{*-}\pi^{+})(\times 10^{-5})$ $1.82\pm 0.87$ $2.03\pm 1.97$ $1.64\pm 0.78$ $1.83\pm 1.78$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{0}\to\overline{K}^{*0}\pi^{0})(\times 10^{-5})$ $0.91\pm 0.43$ $1.01\pm 0.98$ $0.81\pm 0.39$ $0.91\pm 0.88$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{0}\to\rho^{+}K^{-})(\times 10^{-5})$ $4.00\pm 0.54$ $4.23\pm 2.33$ $3.59\pm 0.48$ $3.80\pm 2.09$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{0}\to\rho^{0}\overline{K}^{0})(\times 10^{-5})$ $1.92\pm 0.25$ $2.02\pm 1.10$ $1.72\pm 0.23$ $1.82\pm 0.99$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{0}\to\omega\overline{K}^{0}))(\times 10^{-5})$ $1.04\pm 0.14$ $1.05\pm 0.51$ $0.94\pm 0.12$ $0.94\pm 0.45$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1270)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1270)^{0}\to\overline{K}^{*0}\eta)(\times 10^{-7})$ $5.04\pm 2.15$ $9.06\pm 9.05$ $4.52\pm 1.93$ $8.11\pm 8.10$
$\mathcal{B}(D^{0}\to K_{1}(1400)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{-}\to K^{*-}\pi^{0})(\times 10^{-5})$ $0.60\pm 0.60$ $2.20\pm 2.20$ $0.47\pm 0.47$ $1.73\pm 1.73$
$\mathcal{B}(D^{0}\to K_{1}(1400)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{-}\to\overline{K}^{*0}\pi^{-})(\times 10^{-5})$ $1.19\pm 1.19$ $4.34\pm 4.34$ $0.93\pm 0.93$ $3.41\pm 3.41$
$\mathcal{B}(D^{0}\to K_{1}(1400)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{-}\to\rho^{0}K^{-})(\times 10^{-6})$ $0.31\pm 0.31$ $1.70\pm 1.70$ $0.24\pm 0.24$ $1.33\pm 1.33$
$\mathcal{B}(D^{0}\to K_{1}(1400)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{-}\to\rho^{-}\overline{K}^{0})(\times 10^{-6})$ $0.60\pm 0.60$ $3.34\pm 3.34$ $0.47\pm 0.47$ $2.62\pm 2.62$
$\mathcal{B}(D^{0}\to K_{1}(1400)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{-}\to\omega K^{-})(\times 10^{-6})$ $0.29\pm 0.29$ $1.63\pm 1.63$ $0.23\pm 0.23$ $1.28\pm 1.28$
$\mathcal{B}(D^{0}\to K_{1}(1400)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{-}\to K^{*-}\eta)(\times 10^{-6})$ $0.35\pm 0.35$ $3.51\pm 3.51$ $0.28\pm 0.28$ $2.76\pm 2.76$
$\mathcal{B}(D^{0}\to K_{1}(1400)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{-}\to\phi K^{-})(\times 10^{-6})$ $0.72\pm 0.72$ $2.32\pm 2.32$ $0.57\pm 0.57$ $1.80\pm 1.80$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1400)^{0}\to K^{*-}\pi^{+}(\times 10^{-4})$ $0.32\pm 0.32$ $1.16\pm 1.16$ $0.25\pm 0.25$ $0.91\pm 0.91$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1400)^{0}\to\overline{K}^{*0}\pi^{0})(\times 10^{-5})$ $1.58\pm 1.58$ $5.75\pm 5.75$ $1.24\pm 1.24$ $4.53\pm 4.53$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1400)^{0}\to\rho^{+}K^{-})(\times 10^{-6})$ $1.62\pm 1.62$ $8.95\pm 8.95$ $1.28\pm 1.28$ $7.07\pm 7.07$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1400)^{0}\to\rho^{0}\overline{K}^{0})(\times 10^{-6})$ $0.80\pm 0.80$ $4.40\pm 4.40$ $0.63\pm 0.63$ $3.47\pm 3.47$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1400)^{0}\to\omega\overline{K}^{0})(\times 10^{-6})$ $0.77\pm 0.77$ $4.23\pm 4.23$ $0.60\pm 0.60$ $33.33\pm 3$ 3.33
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1400)^{0}\to\overline{K}^{*0}\eta)(\times 10^{-6})$ $0.88\pm 0.88$ $8.72\pm 8.72$ $0.70\pm 0.70$ $6.89\pm 6.89$
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ \overline{K}_{1}(1400)^{0}\to\phi\overline{K}^{0})(\times 10^{-6})$ $1.84\pm 1.84$ $5.92\pm 5.92$ $1.44\pm 1.44$ $4.63\pm 4.63$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{0}\to K^{*-}\pi^{+})(\times 10^{-6})$ $1.89\pm 1.89$ $6.63\pm 6.63$ $1.60\pm 1.60$ $5.63\pm 5.63$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{0}\to\overline{K}^{*0}\pi^{0})(\times 10^{-6})$ $0.94\pm 0.94$ $3.30\pm 3.30$ $0.80\pm 0.80$ $2.80\pm 2.80$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{0}\to\rho^{+}K^{-})(\times 10^{-7})$ $0.97\pm 0.97$ $5.14\pm 5.14$ $0.82\pm 0.82$ $4.36\pm 4.36$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{0}\to\rho^{0}\overline{K}^{0})(\times 10^{-7})$ $0.47\pm 0.47$ $2.52\pm 2.52$ $0.40\pm 0.40$ $2.14\pm 2.14$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{0}\to\omega\overline{K}^{0})(\times 10^{-7})$ $0.46\pm 0.46$ $2.42\pm 2.42$ $0.39\pm 0.39$ $2.06\pm 2.06$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{0}\to\overline{K}^{*0}\eta)(\times 10^{-7})$ $0.52\pm 0.52$ $5.01\pm 5.01$ $0.44\pm 0.44$ $4.25\pm 4.25$
$\mathcal{B}(D^{+}_{s}\to K_{1}(1400)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ K_{1}(1400)^{0}\to\phi\overline{K}^{0}))(\times 10^{-7})$ $1.09\pm 1.09$ $3.46\pm 3.46$ $0.93\pm 0.93$ $2.92\pm 2.92$
Table 10: Branching ratios of the $D\to VP\ell^{+}\nu_{\ell}$ decays with the $a_{1}/b_{1}/f_{1}$ resonances.
[…]
$\mathcal{B}(D^{0}\to a_{1}(1260)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ a_{1}(1260)^{-}\to\rho^{-}\pi^{0})(\times 10^{-5})$ $1.87\pm 1.09$ $2.89\pm 2.59$ $1.64\pm 0.97$ $2.58\pm 2.33$
$\mathcal{B}(D^{0}\to a_{1}(1260)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ a_{1}(1260)^{-}\to\rho^{0}\pi^{-})(\times 10^{-5})$ $1.87\pm 1.08$ $2.87\pm 2.58$ $1.64\pm 0.96$ $2.56\pm 2.31$
$\mathcal{B}(D^{0}\to a_{1}(1260)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ a_{1}(1260)^{-}\to K^{*0}K^{-})(\times 10^{-6})$ $1.48\pm 0.69$ $1.96\pm 1.65$ $1.30\pm 0.62$ $1.76\pm 1.48$
$\mathcal{B}(D^{0}\to a_{1}(1260)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ a_{1}(1260)^{-}\to K^{*-}K^{0})(\times 10^{-6})$ $1.48\pm 0.70$ $1.97\pm 1.65$ $1.30\pm 0.62$ $1.76\pm 1.49$
$\mathcal{B}(D^{+}\to a_{1}(1260)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ a_{1}(1260)^{0}\to\rho^{+}\pi^{-})(\times 10^{-5})$ $2.41\pm 1.39$ $3.72\pm 3.33$ $2.12\pm 1.24$ $3.31\pm 2.99$
$\mathcal{B}(D^{+}\to a_{1}(1260)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ a_{1}(1260)^{0}\to\rho^{-}\pi^{+})(\times 10^{-5})$ $2.41\pm 1.39$ $3.72\pm 3.33$ $2.12\pm 1.24$ $3.31\pm 2.99$
$\mathcal{B}(D^{+}\to a_{1}(1260)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ a_{1}(1260)^{0}\to K^{*0}\overline{K}^{0})(\times 10^{-6})$ $0.94\pm 0.44$ $1.25\pm 1.05$ $0.83\pm 0.40$ $1.12\pm 0.94$
$\mathcal{B}(D^{+}\to a_{1}(1260)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ a_{1}(1260)^{0}\to\overline{K}^{*0}K^{0})(\times 10^{-6})$ $0.94\pm 0.44$ $1.25\pm 1.05$ $0.83\pm 0.40$ $1.12\pm 0.94$
$\mathcal{B}(D^{+}\to a_{1}(1260)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ a_{1}(1260)^{0}\to K^{*+}K^{-})(\times 10^{-6})$ $0.98\pm 0.45$ $1.29\pm 1.08$ $0.86\pm 0.41$ $1.16\pm 0.98$
$\mathcal{B}(D^{+}\to a_{1}(1260)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ a_{1}(1260)^{0}\to K^{*-}K^{+})(\times 10^{-6})$ $0.98\pm 0.45$ $1.29\pm 1.08$ $0.86\pm 0.41$ $1.16\pm 0.98$
$\mathcal{B}(D^{0}\to b_{1}(1235)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ b_{1}(1235)^{-}\to\omega\pi^{-})(\times 10^{-5})$ $1.89\pm 1.26$ $2.31\pm 2.07$ $1.65\pm 1.11$ $2.02\pm 1.81$
$\mathcal{B}(D^{0}\to b_{1}(1235)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ b_{1}(1235)^{-}\to\rho^{-}\eta)(\times 10^{-6})$ $1.55\pm 1.04$ $1.86\pm 1.66$ $1.36\pm 0.91$ $1.63\pm 1.45$
$\mathcal{B}(D^{0}\to b_{1}(1235)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ b_{1}(1235)^{-}\to K^{*0}K^{-})(\times 10^{-7})$ $2.38\pm 1.70$ $3.07\pm 2.69$ $2.08\pm 1.49$ $2.68\pm 2.35$
$\mathcal{B}(D^{0}\to b_{1}(1235)^{-}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ b_{1}(1235)^{-}\to K^{*-}K^{0})(\times 10^{-7})$ $2.43\pm 1.73$ $3.15\pm 2.76$ $2.12\pm 1.52$ $2.75\pm 2.41$
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\omega\pi^{0})(\times 10^{-5})$ $2.47\pm 1.66$ $3.03\pm 2.72$ $2.17\pm 1.45$ $2.66\pm 2.39$
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\rho^{0}\eta)(\times 10^{-6})$ $1.99\pm 1.33$ $2.38\pm 2.12$ $1.75\pm 1.17$ $2.09\pm 1.87$
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ b_{1}(1235)^{0}\to K^{*0}\overline{K}^{0})(\times 10^{-7})$ $1.46\pm 1.05$ $1.89\pm 1.65$ $1.28\pm 0.92$ $1.65\pm 1.45$
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\overline{K}^{*0}K^{0})(\times 10^{-7})$ $1.46\pm 1.05$ $1.89\pm 1.65$ $1.28\pm 0.92$ $1.65\pm 1.45$
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ b_{1}(1235)^{0}\to K^{*+}K^{-})(\times 10^{-7})$ $1.69\pm 1.21$ $2.16\pm 1.89$ $1.48\pm 1.06$ $1.89\pm 1.65$
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}\ell^{+}\nu_{\ell},\leavevmode\nobreak\ b_{1}(1235)^{0}\to K^{*-}K^{+})(\times 10^{-7})$ $1.69\pm 1.21$ $2.16\pm 1.89$ $1.48\pm 1.06$ $1.89\pm 1.65$
$\mathcal{B}(D^{+}_{s}\to f_{1}(1285)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1285)\to K^{*+}K^{-})(\times 10^{-6})$ $5.13\pm 4.56$ $6.18\pm 6.18$ $4.57\pm 4.06$ $5.50\pm 5.50$
$\mathcal{B}(D^{+}_{s}\to f_{1}(1285)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1285)\to K^{*-}K^{+})(\times 10^{-6})$ $5.13\pm 4.56$ $6.18\pm 6.18$ $4.57\pm 4.06$ $5.50\pm 5.50$
$\mathcal{B}(D^{+}_{s}\to f_{1}(1285)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1285)\to K^{*0}\overline{K}^{0})(\times 10^{-6})$ $1.12\pm 0.99$ $1.62\pm 1.62$ $1.00\pm 0.88$ $1.44\pm 1.44$
$\mathcal{B}(D^{+}_{s}\to f_{1}(1285)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1285)\to\overline{K}^{*0}K^{0})(\times 10^{-6})$ $1.12\pm 0.99$ $1.62\pm 1.62$ $1.00\pm 0.88$ $1.44\pm 1.44$
$\mathcal{B}(D^{+}\to f_{1}(1285)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1285)\to K^{*+}K^{-})(\times 10^{-7})$ $1.00\pm 1.00$ $1.27\pm 1.27$ $0.85\pm 0.85$ $1.09\pm 1.09$
$\mathcal{B}(D^{+}\to f_{1}(1285)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1285)\to K^{*-}K^{+})(\times 10^{-7})$ $1.00\pm 1.00$ $1.27\pm 1.27$ $0.85\pm 0.85$ $1.09\pm 1.09$
$\mathcal{B}(D^{+}\to f_{1}(1285)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1285)\to K^{*0}\overline{K}^{0})(\times 10^{-8})$ $2.30\pm 2.30$ $3.39\pm 3.39$ $1.97\pm 1.97$ $2.90\pm 2.90$
$\mathcal{B}(D^{+}\to f_{1}(1285)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1285)\to\overline{K}^{*0}K^{0})(\times 10^{-8})$ $2.30\pm 2.30$ $3.39\pm 3.39$ $1.97\pm 1.97$ $2.90\pm 2.90$
$\mathcal{B}(D^{+}_{s}\to f_{1}(1420)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1420)\to K^{*+}K^{-})(\times 10^{-5})$ $5.24\pm 5.24$ $8.23\pm 8.23$ $4.36\pm 4.36$ $6.85\pm 6.85$
$\mathcal{B}(D^{+}_{s}\to f_{1}(1420)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1420)\to K^{*-}K^{+})(\times 10^{-5})$ $5.24\pm 5.24$ $8.23\pm 8.23$ $4.36\pm 4.36$ $6.85\pm 6.85$
$\mathcal{B}(D^{+}_{s}\to f_{1}(1420)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1420)\to K^{*0}\overline{K}^{0})(\times 10^{-5})$ $4.71\pm 4.71$ $7.36\pm 7.36$ $3.92\pm 3.92$ $6.13\pm 6.13$
$\mathcal{B}(D^{+}_{s}\to f_{1}(1420)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1420)\to\overline{K}^{*0}K^{0})(\times 10^{-5})$ $4.71\pm 4.71$ $7.36\pm 7.36$ $3.92\pm 3.92$ $6.13\pm 6.13$
$\mathcal{B}(D^{+}\to f_{1}(1420)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1420)\to K^{*+}K^{-})(\times 10^{-6})$ $1.46\pm 1.46$ $1.93\pm 1.93$ $11.11\pm 1$ 1.11 $1.47\pm 1.47$
$\mathcal{B}(D^{+}\to f_{1}(1420)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1420)\to K^{*-}K^{+})(\times 10^{-6})$ $1.46\pm 1.46$ $1.93\pm 1.93$ $11.11\pm 1$ 1.11 $1.47\pm 1.47$
$\mathcal{B}(D^{+}\to f_{1}(1420)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1420)\to K^{*0}\overline{K}^{0})(\times 10^{-6})$ $1.31\pm 1.31$ $1.71\pm 1.71$ $1.00\pm 1.00$ $1.31\pm 1.31$
$\mathcal{B}(D^{+}\to f_{1}(1420)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ f_{1}(1420)\to\overline{K}^{*0}K^{0})(\times 10^{-6})$ $1.31\pm 1.31$ $1.71\pm 1.71$ $1.00\pm 1.00$ $1.31\pm 1.31$
Table 11: Branching ratios of the $D\to VP\ell^{+}\nu_{\ell}$ decays with the $h_{1}$ resonances.
[…]
$\mathcal{B}(D^{+}_{s}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to\rho^{0}\pi^{0})(\times 10^{-5})$ $4.23\pm 4.23$ $5.98\pm 5.98$ $3.89\pm 3.88$ $5.48\pm 5.48$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to\rho^{+}\pi^{-})(\times 10^{-5})$ $4.21\pm 4.21$ $5.95\pm 5.95$ $3.86\pm 3.86$ $5.45\pm 5.45$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to\rho^{-}\pi^{+})(\times 10^{-5})$ $4.21\pm 4.21$ $5.95\pm 5.95$ $3.86\pm 3.86$ $5.45\pm 5.45$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to K^{*+}K^{-})(\times 10^{-6})$ $3.71\pm 3.71$ $6.34\pm 6.34$ $3.41\pm 3.41$ $5.82\pm 5.82$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to K^{*-}K^{+})(\times 10^{-6})$ $3.71\pm 3.71$ $6.34\pm 6.34$ $3.41\pm 3.41$ $5.82\pm 5.82$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to K^{*0}\overline{K}^{0})(\times 10^{-6})$ $3.52\pm 3.52$ $6.05\pm 6.05$ $3.23\pm 3.23$ $55.55\pm 5$ 5.55
$\mathcal{B}(D^{+}_{s}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to\overline{K}^{*0}K^{0})(\times 10^{-6})$ $3.52\pm 3.52$ $6.05\pm 6.05$ $3.23\pm 3.23$ $55.55\pm 5$ 5.55
$\mathcal{B}(D^{+}_{s}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to\omega\eta)(\times 10^{-6})$ $4.07\pm 4.07$ $6.62\pm 6.62$ $3.74\pm 3.74$ $6.07\pm 6.07$
$\mathcal{B}(D^{+}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to\rho^{0}\pi^{0})(\times 10^{-5})$ $1.40\pm 1.06$ $1.62\pm 1.56$ $1.26\pm 0.95$ $1.45\pm 1.40$
$\mathcal{B}(D^{+}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to\rho^{+}\pi^{-})(\times 10^{-5})$ $1.40\pm 1.05$ $1.61\pm 1.55$ $1.25\pm 0.94$ $1.44\pm 1.39$
$\mathcal{B}(D^{+}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to\rho^{-}\pi^{+})(\times 10^{-5})$ $1.40\pm 1.05$ $1.61\pm 1.55$ $1.25\pm 0.94$ $1.44\pm 1.39$
$\mathcal{B}(D^{+}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to K^{*+}K^{-})(\times 10^{-7})$ $5.12\pm 4.82$ $6.92\pm 6.92$ $4.58\pm 4.31$ $6.22\pm 6.22$
$\mathcal{B}(D^{+}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to K^{*-}K^{+})(\times 10^{-7})$ $5.12\pm 4.82$ $6.92\pm 6.92$ $4.58\pm 4.31$ $6.22\pm 6.22$
$\mathcal{B}(D^{+}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to K^{*0}\overline{K}^{0})(\times 10^{-7})$ $4.87\pm 4.59$ $6.57\pm 6.57$ $4.36\pm 4.10$ $5.90\pm 5.90$
$\mathcal{B}(D^{+}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to\overline{K}^{*0}K^{0})(\times 10^{-7})$ $4.87\pm 4.59$ $6.57\pm 6.57$ $4.36\pm 4.10$ $5.90\pm 5.90$
$\mathcal{B}(D^{+}\to h_{1}(1170)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1170)\to\omega\eta)(\times 10^{-6})$ $1.21\pm 0.96$ $1.46\pm 1.41$ $1.09\pm 0.86$ $1.31\pm 1.27$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to\rho^{0}\pi^{0})(\times 10^{-5})$ $3.48\pm 3.48$ $4.94\pm 4.94$ $2.93\pm 2.93$ $4.17\pm 4.17$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to\rho^{+}\pi^{-})(\times 10^{-5})$ $3.47\pm 3.47$ $4.93\pm 4.93$ $2.92\pm 2.92$ $4.16\pm 4.16$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to\rho^{-}\pi^{+})(\times 10^{-5})$ $3.47\pm 3.47$ $4.93\pm 4.93$ $2.92\pm 2.92$ $4.16\pm 4.16$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to K^{*+}K^{-})(\times 10^{-5})$ $4.04\pm 3.70$ $5.08\pm 5.08$ $3.39\pm 3.11$ $4.27\pm 4.27$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to K^{*-}K^{+})(\times 10^{-5})$ $4.04\pm 3.70$ $5.08\pm 5.08$ $3.39\pm 3.11$ $4.27\pm 4.27$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to K^{*0}\overline{K}^{0})(\times 10^{-5})$ $3.43\pm 3.16$ $4.28\pm 4.27$ $2.88\pm 2.65$ $3.59\pm 3.59$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to\overline{K}^{*0}K^{0})(\times 10^{-5})$ $3.43\pm 3.16$ $4.28\pm 4.27$ $2.88\pm 2.65$ $3.59\pm 3.59$
$\mathcal{B}(D^{+}_{s}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to\omega\eta)(\times 10^{-5})$ $0.89\pm 0.89$ $1.29\pm 1.29$ $0.75\pm 0.75$ $1.09\pm 1.09$
$\mathcal{B}(D^{+}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to\rho^{0}\pi^{0})(\times 10^{-7})$ $2.79\pm 2.79$ $5.23\pm 5.23$ $2.16\pm 2.16$ $4.09\pm 4.09$
$\mathcal{B}(D^{+}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to\rho^{+}\pi^{-})(\times 10^{-7})$ $2.78\pm 2.78$ $5.22\pm 5.22$ $2.16\pm 2.16$ $4.08\pm 4.08$
$\mathcal{B}(D^{+}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to\rho^{-}\pi^{+})(\times 10^{-7})$ $2.78\pm 2.78$ $5.22\pm 5.22$ $2.16\pm 2.16$ $4.08\pm 4.08$
$\mathcal{B}(D^{+}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to K^{*+}K^{-})(\times 10^{-7})$ $0.74\pm 0.74$ $1.46\pm 1.46$ $0.57\pm 057$ $1.14\pm 1.14$
$\mathcal{B}(D^{+}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to K^{*-}K^{+})(\times 10^{-7})$ $0.74\pm 0.74$ $1.46\pm 1.46$ $0.57\pm 0.57$ $1.14\pm 1.14$
$\mathcal{B}(D^{+}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to K^{*0}\overline{K}^{0})(\times 10^{-7})$ $0.60\pm 0.60$ $1.02\pm 1.02$ $0.46\pm 0.46$ $0.80\pm 0.80$
$\mathcal{B}(D^{+}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to\overline{K}^{*0}K^{0})(\times 10^{-7})$ $0.60\pm 0.60$ $1.02\pm 1.02$ $0.46\pm 0.46$ $0.80\pm 0.80$
$\mathcal{B}(D^{+}\to h_{1}(1415)\ell^{+}\nu_{\ell},\leavevmode\nobreak\ h_{1}(1415)\to\omega\eta)(\times 10^{-7})$ $0.71\pm 0.71$ $1.09\pm 1.09$ $0.55\pm 0.55$ $0.85\pm 0.85$
Table 12: Total branching ratios of the $D\rightarrow VP\ell^{+}\nu_{\ell}$ decays with multiple axial-vector resonances. $\mathcal{B}_{[R_{1},R_{2}]}$ denotes the branching ratios with the $R_{1}$ and $R_{2}$ resonant contributions.
[…]
$\mathcal{B}(D^{0}\rightarrow K^{\ast-}\pi^{0}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-4})$ $\leq 1.11$ $\leq 1.71$ $\leq 0.95$ $\leq 1.46$
$\mathcal{B}(D^{0}\rightarrow\overline{K}^{\ast 0}\pi^{-}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-4})$ $\leq 2.17$ $\leq 3.35$ $\leq 1.86$ $\leq 2.87$
$\mathcal{B}(D^{0}\rightarrow\rho^{0}K^{-}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-4})$ $\leq 1.77$ $\leq 2.59$ $\leq 1.53$ $\leq 2.24$
$\mathcal{B}(D^{0}\rightarrow\rho^{-}\overline{K}^{0}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-4})$ $\leq 3.45$ $\leq 5.00$ $\leq 2.98$ $\leq 4.32$
$\mathcal{B}(D^{0}\rightarrow\omega K^{-}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-4})$ $\leq 1.01$ $\leq 1.33$ $\leq 0.87$ $\leq 1.14$
$\mathcal{B}(D^{0}\rightarrow K^{\ast-}\eta\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-5})$ $\leq 0.68$ $\leq 1.62$ $\leq 0.59$ $\leq 1.39$
$\mathcal{B}(D^{0}\rightarrow\rho^{-}\pi^{0}\ell^{+}\nu_{\ell})_{[a_{1}(1260)]}(\times 10^{-5})$ $\leq 2.96$ $\leq 5.48$ $\leq 2.61$ $\leq 4.91$
$\mathcal{B}(D^{0}\rightarrow\rho^{0}\pi^{-}\ell^{+}\nu_{\ell})_{[a_{1}(1260)]}(\times 10^{-5})$ $\leq 2.95$ $\leq 5.45$ $\leq 2.60$ $\leq 4.88$
$\mathcal{B}(D^{0}\rightarrow K^{\ast 0}K^{-}\ell^{+}\nu_{\ell})_{[a_{1}(1260),b_{1}(1235)]}(\times 10^{-6})$ $\leq 2.28$ $\leq 3.70$ $\leq 2.01$ $\leq 3.32$
$\mathcal{B}(D^{0}\rightarrow K^{\ast-}K^{0}\ell^{+}\nu_{\ell})_{[a_{1}(1260),b_{1}(1235)]}(\times 10^{-6})$ $\leq 2.29$ $\leq 3.71$ $\leq 2.02$ $\leq 3.33$
$\mathcal{B}(D^{0}\rightarrow\omega\pi^{-}\ell^{+}\nu_{\ell})_{[b_{1}(1235)]}(\times 10^{-5})$ $\leq 3.15$ $\leq 4.38$ $\leq 2.76$ $\leq 3.83$
$\mathcal{B}(D^{0}\rightarrow\rho^{-}\eta\ell^{+}\nu_{\ell})_{[b_{1}(1235)]}(\times 10^{-6})$ $\leq 2.59$ $\leq 3.52$ $\leq 2.26$ $\leq 3.08$
$\mathcal{B}(D^{+}\rightarrow K^{\ast-}\pi^{+}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-4})$ $\leq 5.72$ $\leq 8.90$ $\leq 4.91$ $\leq 7.63$
$\mathcal{B}(D^{+}\rightarrow\overline{K}^{\ast 0}\pi^{0}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-4})$ $\leq 2.84$ $\leq 4.43$ $\leq 2.44$ $\leq 3.80$
$\mathcal{B}(D^{+}\rightarrow\rho^{+}K^{-}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-3})$ $\leq 0.93$ $\leq 1.35$ $\leq 0.80$ $\leq 1.17$
$\mathcal{B}(D^{+}\rightarrow\rho^{0}\overline{K}^{0}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-4})$ $\leq 4.46$ $\leq 6.42$ $\leq 3.86$ $\leq 5.57$
$\mathcal{B}(D^{+}\rightarrow\omega\overline{K}^{0}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-4})$ $\leq 2.45$ $\leq 3.21$ $\leq 2.12$ $\leq 2.77$
$\mathcal{B}(D^{+}\rightarrow\overline{K}^{\ast 0}\eta\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-5})$ $\leq 1.49$ $\leq 3.77$ $\leq 1.29$ $\leq 3.24$
$\mathcal{B}(D^{+}\rightarrow\rho^{+}\pi^{-}\ell^{+}\nu_{\ell})_{[a_{1}(1260),h_{1}(1170),h_{1}(1415)]}(\times 10^{-5})$ $\leq 4.46$ $\leq 8.20$ $\leq 3.96$ $\leq 7.30$
$\mathcal{B}(D^{+}\rightarrow\rho^{-}\pi^{+}\ell^{+}\nu_{\ell})_{[a_{1}(1260),h_{1}(1170),h_{1}(1415)]}(\times 10^{-5})$ $\leq 4.46$ $\leq 8.20$ $\leq 3.96$ $\leq 7.30$
$\mathcal{B}(D^{+}\rightarrow K^{\ast 0}\overline{K}^{0}\ell^{+}\nu_{\ell})_{\big{[}^{a_{1}(1260),b_{1}(1235),f_{1}(1285)}_{f_{1}(1420),h_{1}(1170),h_{1}(1415)}\big{]}}(\times 10^{-6})$ $\leq 4.00$ $\leq 5.38$ $\leq 3.26$ $\leq 4.35$
$\mathcal{B}(D^{+}\rightarrow\overline{K}^{\ast 0}K^{0}\ell^{+}\nu_{\ell})_{\big{[}^{a_{1}(1260),b_{1}(1235),f_{1}(1285)}_{f_{1}(1420),h_{1}(1170),h_{1}(1415)}\big{]}}(\times 10^{-6})$ $\leq 4.00$ $\leq 5.38$ $\leq 3.26$ $\leq 4.35$
$\mathcal{B}(D^{+}\rightarrow K^{\ast+}K^{-}\ell^{+}\nu_{\ell})_{\big{[}^{a_{1}(1260),b_{1}(1235),f_{1}(1285)}_{f_{1}(1420),h_{1}(1170),h_{1}(1415)}\big{]}}(\times 10^{-6})$ $\leq 4.36$ $\leq 5.88$ $\leq 3.55$ $\leq 4.74$
$\mathcal{B}(D^{+}\rightarrow K^{\ast-}K^{+}\ell^{+}\nu_{\ell})_{\big{[}^{a_{1}(1260),b_{1}(1235),f_{1}(1285)}_{f_{1}(1420),h_{1}(1170),h_{1}(1415)}\big{]}}(\times 10^{-6})$ $\leq 4.36$ $\leq 5.88$ $\leq 3.55$ $\leq 4.74$
$\mathcal{B}(D^{+}\rightarrow\omega\pi^{0}\ell^{+}\nu_{\ell})_{[b_{1}(1235)]}(\times 10^{-5})$ $\leq 4.13$ $\leq 5.75$ $\leq 3.62$ $\leq 5.05$
$\mathcal{B}(D^{+}\rightarrow\rho^{0}\eta\ell^{+}\nu_{\ell})_{[b_{1}(1235)]}(\times 10^{-6})$ $\leq 3.33$ $\leq 4.50$ $\leq 2.92$ $\leq 3.95$
$\mathcal{B}(D^{+}\rightarrow\rho^{0}\pi^{0}\ell^{+}\nu_{\ell})_{[h_{1}(1170),h_{1}(1415)]}(\times 10^{-5})$ $\leq 2.46$ $\leq 3.18$ $\leq 2.21$ $\leq 2.85$
$\mathcal{B}(D^{+}\rightarrow\omega\eta\ell^{+}\nu_{\ell})_{[h_{1}(1170),h_{1}(1415)]}(\times 10^{-6})$ $\leq 2.17$ $\leq 2.88$ $\leq 1.94$ $\leq 2.59$
$\mathcal{B}(D^{+}\rightarrow\phi\eta\ell^{+}\nu_{\ell})_{[h_{1}^{0}(1170),h_{1}(1415)]}(\times 10^{-7})$ $\leq 1.05$ $\leq 1.65$ $\leq 0.92$ $\leq 1.44$
$\mathcal{B}(D_{s}^{+}\rightarrow K^{\ast+}K^{-}\ell^{+}\nu_{\ell})_{[f_{1}(1285),f_{1}(1420),h_{1}(1170),h_{1}(1415)]}(\times 10^{-4})$ $\leq 1.28$ $\leq 2.12$ $\leq 1.06$ $\leq 1.77$
$\mathcal{B}(D_{s}^{+}\rightarrow K^{\ast-}K^{+}\ell^{+}\nu_{\ell})_{[f_{1}(1285),f_{1}(1420),h_{1}(1170),h_{1}(1415)]}(\times 10^{-4})$ $\leq 1.28$ $\leq 2.12$ $\leq 1.06$ $\leq 1.77$
$\mathcal{B}(D_{s}^{+}\rightarrow K^{\ast 0}\overline{K}^{0}\ell^{+}\nu_{\ell})_{[f_{1}(1285),f_{1}(1420),h_{1}(1170),h_{1}(1415)]}(\times 10^{-4})$ $\leq 1.13$ $\leq 1.82$ $\leq 0.94$ $\leq 1.52$
$\mathcal{B}(D_{s}^{+}\rightarrow\overline{K}^{\ast 0}K^{0}\ell^{+}\nu_{\ell})_{[f_{1}(1285),f_{1}(1420),h_{1}(1170),h_{1}(1415)]}(\times 10^{-4})$ $\leq 1.13$ $\leq 1.82$ $\leq 0.94$ $\leq 1.52$
$\mathcal{B}(D_{s}^{+}\rightarrow\rho^{0}\pi^{0}\ell^{+}\nu_{\ell})_{[f_{1}(1285),f_{1}(1420),h_{1}(1170),h_{1}(1415)]}(\times 10^{-4})$ $\leq 1.54$ $\leq 2.08$ $\leq 1.36$ $\leq 1.84$
$\mathcal{B}(D_{s}^{+}\rightarrow\rho^{+}\pi^{-}\ell^{+}\nu_{\ell})_{[f_{1}(1285),f_{1}(1420),h_{1}(1170),h_{1}(1415)]}(\times 10^{-4})$ $\leq 1.54$ $\leq 2.07$ $\leq 1.36$ $\leq 1.83$
$\mathcal{B}(D_{s}^{+}\rightarrow\rho^{-}\pi^{+}\ell^{+}\nu_{\ell})_{[f_{1}(1285),f_{1}(1420),h_{1}(1170),h_{1}(1415)]}(\times 10^{-4})$ $\leq 1.54$ $\leq 2.07$ $\leq 1.36$ $\leq 1.83$
$\mathcal{B}(D_{s}^{+}\rightarrow\omega\eta\ell^{+}\nu_{\ell})_{[f_{1}(1285),f_{1}(1420),h_{1}(1170),h_{1}(1415)]}(\times 10^{-5})$ $\leq 2.59$ $\leq 3.72$ $\leq 2.24$ $\leq 3.23$
$\mathcal{B}(D_{s}^{+}\rightarrow\phi\eta\ell^{+}\nu_{\ell})_{[f_{1}(1285),f_{1}(1420),h_{1}(1170),h_{1}(1415)]}(\times 10^{-5})$ $\leq 0.80$ $\leq 1.14$ $\leq 0.67$ $\leq 0.95$
$\mathcal{B}(D_{s}^{+}\rightarrow K^{\ast-}\pi^{+}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-5})$ $\leq 2.83$ $\leq 4.40$ $\leq 2.52$ $\leq 3.92$
$\mathcal{B}(D_{s}^{+}\rightarrow\overline{K}^{\ast 0}\pi^{0}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-5})$ $\leq 1.40$ $\leq 2.19$ $\leq 1.25$ $\leq 1.95$
$\mathcal{B}(D_{s}^{+}\rightarrow\rho^{+}K^{-}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-5})$ $\leq 4.53$ $\leq 6.55$ $\leq 4.07$ $\leq 5.89$
$\mathcal{B}(D_{s}^{+}\rightarrow\rho^{0}\overline{K}^{0}\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-5})$ $\leq 2.17$ $\leq 3.12$ $\leq 1.95$ $\leq 2.81$
$\mathcal{B}(D_{s}^{+}\rightarrow\omega\overline{K}^{0})\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-5})$ $\leq 1.18$ $\leq 1.56$ $\leq 1.06$ $\leq 1.40$
$\mathcal{B}(D_{s}^{+}\rightarrow\overline{K}^{\ast 0}\eta\ell^{+}\nu_{\ell})_{[K_{1}(1270),K_{1}(1400)]}(\times 10^{-6})$ $\leq 0.73$ $\leq 1.86$ $\leq 0.66$ $\leq 1.66$
ACKNOWLEDGEMENTS
We thank Hai-Bo Li, Mao-Zhi Yang and Xian-Wei Kang for useful discussions. The work was supported by the National Natural Science Foundation of China (No. 12175088 and No. 12365014).