### v2

# Study of the axial-vector and tensor resonant contributions to the $D\to VP\ell^{+}\nu_{\ell}$ decays based on SU(3) flavor analysis

Abstract
Semileptonic three-body $D\to M\ell^{+}\nu_{\ell}$ decays, non-leptonic $M\to VP$ decays, and semileptonic four-body $D\to M(M\to VP)\ell^{+}\nu_{\ell}$ decays are analyzed using the SU(3) flavor symmetry/breaking approach, where $\ell=e/\mu$, $M=A/T$, and $A/T/V/P$ denote the axial-vector/tensor/vector/pseudoscalar mesons, respectively. In terms of SU(3) flavor symmetry/breaking, the decay amplitudes of the $D\to M\ell^{+}\nu_{\ell}$ decays and the vertex coefficients of the $M\to VP$ decays are related. The relevant non-perturbative parameters of the $D\to A\ell^{+}\nu_{\ell}$, $A\to VP$ and $T\to VP$ decays are constrained by the present experimental data, and the non-perturbative parameters of $D\to T\ell^{+}\nu_{\ell}$ decays are taken from the results in the light-front quark model since no experimental data are available at present. The branching ratios of the $D\to M\ell^{+}\nu_{\ell}$, $M\to VP$, and $D\to M(M\to VP)\ell^{+}\nu_{\ell}$ decays are then predicted. We find that some processes receive both tensor and axial-vector resonant contributions, while other processes receive only axial-vector resonant contributions. In cases where both kinds of resonant contributions exist, the axial-vector contributions are dominant. Some branching ratios with axial-vector resonance states are large, and they may be measured experimentally in the near future. In addition, the sensitivities of the branching ratios of $D\to A\ell^{+}\nu_{\ell}$ and $A\to VP$ decays to the parameters are also investigated, and some decay branching ratios are found to be sensitive to the non-perturbative parameters.
I INTRODUCTION
Semileptonic decays of the charmed mesons are very useful for determining the quark mixing parameters and values of the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements, and in searching for new physics beyond the standard model. The intermediate resonances in these semileptonic decays provide a clean environment in which to explore meson spectroscopy, as there is no interference from other particles Ke:2023qzc . Among the semileptonic four-body decays $D\to VP\ell^{+}\nu_{\ell}$, only the branching ratios of $D^{+}\to b^{0}_{1}(b_{1}^{0}\to\omega\pi^{0})e^{+}\nu_{e}$ and $D^{0}\to b^{-}_{1}(b_{1}^{-}\to\omega\pi^{-})e^{+}\nu_{e}$ decays have experimental upper limits PDG2022 . Nevertheless, two semileptonic three-body decays $D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e}$ and $D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e}$ have been found by the CLEO Collaboration CLEO:2007oer , and they have been much improved by the BESIII Collaboration BESIII:2019eao ; BESIII:2021uqr . The six non-leptonic decays $K_{1}\to VP$ were measured by the ACCMOR Collaboration ACCMOR:1981yww . One can use them to study the $D\to VP\ell^{+}\nu_{\ell}$ semileptonic decays with the axial-vector resonance states. As for the $D\to VP\ell^{+}\nu_{\ell}$ decays with the tensor resonance states, only four $T\to VP$ modes have been measured PDG2022 ; as there are no experimental data in either $D\to T\ell^{+}\nu_{\ell}$ or $D\to T(T\to VP)\ell^{+}\nu_{\ell}$ decays, they can be studied by combining theoretical calculations.
For semileptonic decays, since the leptons do not participate in the strong interaction, the weak and strong dynamics can be separated, and the theoretical description of the semileptonic decays is relatively simple. The hadronic transition form factors contain all the strong dynamics in the initial and final hadrons, and they are crucial for testing the theoretical calculations of the involved strong interaction. The structure of the $D\to VP$ form factors is more complicated than for the $D\to V$ or $D\to P$ form factors; nevertheless, we have not found their systematic calculations up to now. In the absence of reliable calculations, symmetry analysis can provide very useful information about the decays. Because SU(3) flavor symmetry is independent of the detailed dynamics, it can provide us an opportunity to relate different decay modes. Even though SU(3) flavor symmetry is only an approximate symmetry, it still provides some valuable information about the decays. SU(3) flavor symmetry has been widely used to study hadron decays, including b-hadron decays He:1998rq ; He:2000ys ; Fu:2003fy ; Hsiao:2015iiu ; He:2015fwa ; He:2015fsa ; Deshpande:1994ii ; Gronau:1994rj ; Gronau:1995hm ; Shivashankara:2015cta ; Zhou:2016jkv ; Cheng:2014rfa ; Wang:2021uzi ; Wang:2020wxn and c-hadron decays Wang:2021uzi ; Wang:2020wxn ; Grossman:2012ry ; Pirtskhalava:2011va ; Cheng:2012xb ; Savage:1989qr ; Savage:1991wu ; Altarelli:1975ye ; Lu:2016ogy ; Geng:2017esc ; Geng:2018plk ; Geng:2017mxn ; Geng:2019bfz ; Wang:2017azm ; Wang:2019dls ; Wang:2017gxe ; Muller:2015lua .
Semileptonic $D\to M\ell^{+}\nu_{\ell}$ decays have been studied, for example, by the quantum chromodynamics (QCD) sum rules approach Momeni:2020zrb ; Momeni:2019uag ; Zuo:2016msr ; Hu:2021lkl ; Khosravi:2008jw , light-cone sum rules Huang:2021owr , and quark models Isgur:1988gb ; Scora:1995ty ; Cheng:2017pcq . The non-leptonic $M\to VP$ decays have been studied by many methods, for examples, the chiral unitary approach Geng:2006yb ; Cabrera:2009ep , quark models Ackleh:1996yt ; Bloch:1999vka ; Du:2021zdg , and the effective Lagrangian of the strong interaction Lichard:2010zza ; Li:1995aw ; Li:1995tv ; Wess:1967jq ; Roca:2003uk ; Vojik:2010ua . The four-body semileptonic $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays have also been studied by the unitary extensions of chiral perturbation theory Wang:2020pyy . In this work, we will analyze the $D\to M(M\to VP)\ell^{+}\nu_{\ell}$ four-body semileptonic decays using SU(3) flavor symmetry/breaking approach. We will first obtain the decay amplitude relations of the $D\to M\ell^{+}\nu_{\ell}$ or the vertex coefficient relations of $M\to VP$ by using SU(3) flavor symmetry/breaking. Then, after ignoring the SU(3) flavor breaking effects due to poor experimental data of $D\to M\ell^{+}\nu_{\ell}$, $M\to VP$ and $D\to M(M\to VP)\ell^{+}\nu_{\ell}$, the relevant non-perturbative coefficients and mixing angles will be constrained by the present experimental data. Finally, the branching ratios which have not yet been measured will be predicted by the SU(3) flavor symmetry. In addition, the sensitivity of $\mathcal{B}(D\to A\ell^{+}\nu_{\ell})$ and $\mathcal{B}(A\to VP)$ decays to the non-perturbative parameters will be displayed.
This paper is organized as follows. In Sect. II, the $D\rightarrow A\ell^{+}\nu_{\ell}$, $A\to VP$ and $D\rightarrow A(A\to PV)\ell^{+}\nu_{\ell}$ decays are presented. In Sect. III, the $D\rightarrow T\ell^{+}\nu_{\ell}$, $T\to VP$ and $D\rightarrow T(T\to PV)\ell^{+}\nu_{\ell}$ decays are discussed. We summarize our results in Sect. IV. Finally, the meson multiplets of the SU(3) flavor group are listed in the Appendix.
II Semileptonic $D\rightarrow A(A\to PV)\ell^{+}\nu_{\ell}$ decays
II.1 Semileptonic $D\rightarrow A\ell^{+}\nu_{\ell}$ decays
II.1.1 Theoretical framework for the $D\to A\ell^{+}\nu_{\ell}$ decays
For the semileptonic decays of the $c\rightarrow q\ell^{+}\nu_{\ell}$ transition, the effective Hamiltonian can be expressed as
$\mathcal{H}_{eff}(c\rightarrow q\ell^{+}\nu_{\ell})$ $=$ $\frac{G_{F}}{\sqrt{2}}V_{cq}^{*}\bar{q}\gamma^{\mu}(1-\gamma_{5})c\leavevmode\nobreak\ \bar{\nu}_{\ell}\gamma_{\mu}(1-\gamma_{5})\ell,$ (1)
where $G_{F}$ is the Fermi constant, $V_{cq}$ is the CKM matrix element with $q=d,s$. The decay amplitudes of the $D\to A\ell^{+}\nu_{\ell}$ decays can be written as Wang:2022yyn
$\mathcal{M}(D\rightarrow A\ell^{+}\nu_{\ell})$ $=$ $\frac{G_{F}}{\sqrt{2}}\sum_{m,n}g_{mn}L^{\lambda_{\ell}\lambda_{\nu}}_{m}H^{\lambda_{A}}_{n}.$ (2)
Leptonic amplitudes $L^{\lambda_{\ell}\lambda_{\nu}}_{m}$ and hadronic amplitudes $H^{\lambda_{A}}_{n}$ are represented as
$L^{\lambda_{\ell}\lambda_{\nu}}_{m}$ $=$ $\epsilon_{\alpha}(m)\bar{\nu_{\ell}}\gamma^{\alpha}(1-\gamma_{5})\ell,$ (3)
$H^{\lambda_{A}}_{n}$ $=$ $V_{cq}^{*}\epsilon_{\beta}^{*}(n)\langle{A}(p,\varepsilon^{*})|\bar{q}\gamma^{\beta}(1-\gamma_{5})c|D(p_{D})\rangle,$ (4)
where $\lambda_{A}=0,\pm 1$, $\lambda_{\ell}=\pm\frac{1}{2}$, and $\lambda_{\nu}=+\frac{1}{2}$ are the particle helicity, and $\epsilon(m)$ with $m=0,t,\pm 1$ and $\varepsilon^{*}$ are the polarization vectors of the virtual $W$ and axial-vector mesons $A$, respectively.
The hadron amplitudes can be parameterized by form factors of the $D\to A$ transition Cheng:2003sm
$A(p,\varepsilon^{*})\left|\bar{q}\gamma_{\mu}(1-\gamma_{5})c\right|D(p_{D})\right>$ $=$ $\frac{2iA(q^{2})}{m_{D}-m_{A}}\epsilon_{\mu\nu\alpha\beta}\varepsilon^{*\nu}p^{\alpha}_{D}p^{\beta}$ (5)
$-\left[\varepsilon^{*}_{\mu}(m_{D}-m_{A})V_{1}(q^{2})-(p_{D}+p)_{\mu}(\varepsilon^{*}.p_{D})\frac{V_{2}(q^{2})}{m_{D}-m_{A}}\right]$
$+q_{\mu}(\varepsilon^{*}.p_{D})\frac{2m_{A}}{q^{2}}[V_{3}(q^{2})-V_{0}(q^{2})],$
where $q^{2}\equiv(p_{D}-p)^{2}$, $A(q^{2})$ and $V_{0,1,2,3}(q^{2})$ are the form factors of the $D\to A$ transition, and $m_{A,D,\ell}$ are the masses of $A,D,\ell$. Ten hadronic amplitudes can then be written as Sun:2011ssd ; Li:2009tx ; Colangelo:2019axi
$H_{\pm}$ $=$ $(m_{D}-m_{A})V_{1}(q^{2})\mp\frac{2m_{D}|\vec{p}_{A}|}{(m_{D}-m_{A})}A(q^{2}),$ (6)
$H_{0}$ $=$ $\frac{1}{2m_{A}\sqrt{q^{2}}}\left[(m_{D}^{2}-m_{A}^{2}-q^{2})(m_{D}-m_{A})V_{1}(q^{2})-\frac{4m_{D}^{2}|\vec{p}_{A}|^{2}}{m_{D}-m_{A}}V_{2}(q^{2})\right],$ (7)
$H_{t}$ $=$ $\frac{2m_{D}|\vec{p}_{A}|}{\sqrt{q^{2}}}V_{0}(q^{2}),$ (8)
in which $|\vec{p}_{A}|\equiv\sqrt{\lambda(m_{D}^{2},m_{A}^{2},q^{2})}/(2m_{D})$ with $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2ab-2ac-2bc$.
The differential branching ratios are Ivanov:2019nqd
$\frac{d\mathcal{B}(D\to A\ell^{+}\nu_{\ell})}{dq^{2}}=\frac{\tau_{D}G_{F}^{2}|V_{cq}|^{2}\lambda^{1/2}(q^{2}-m_{\ell}^{2})^{2}}{24(2\pi)^{3}m_{D}^{3}q^{2}}{\cal H}_{\rm total},$ (9)
with ${\cal H}_{\rm total}\equiv({\cal H}_{U}+{\cal H}_{L})\left(1+\frac{m_{\ell}^{2}}{2q^{2}}\right)+\frac{3m_{\ell}^{2}}{2q^{2}}{\cal H}_{S},$ (10)
and ${\cal H}_{U}=|H_{+}|^{2}+|H_{-}|^{2},\quad{\cal H}_{L}=|H_{0}|^{2},\quad{\cal H}_{P}=|H_{+}|^{2}-|H_{-}|^{2},\quad{\cal H}_{S}=|H_{t}|^{2},\quad{\cal H}_{SL}=\Re(H_{0}H_{t}^{{\dagger}}),$ (11)
where $\lambda\equiv\lambda(m_{D}^{2},m_{A}^{2},q^{2})$, $m_{\ell}^{2}\leq q^{2}\leq(m_{D}-m_{A})^{2}$, and $\tau_{D}$ are the mean lifetime of $D$ meson. The branching ratios are usually obtained by Eq. (9) and the form factors; however, the results of the form factors are dependent on different approaches.
Branching ratios also can be obtained by using the available data and the SU(3) flavor symmetry/breaking, which is independent of the detailed dynamics, offering us an opportunity to relate different decay modes. In the $\ell=e,\mu$ cases, $m_{\ell}^{2}$ is small, and it can be ignored in ${\cal H}_{\rm total}$ in Eq. (10). Then ${\cal H}_{\rm total}$ only contains the hadronic information, which can be related by the the SU(3) flavor symmetry/breaking.
The SU(3) flavor symmetry analysis can provide the hadronic helicity amplitude relationships among various decay models. The representations for meson multiplets of the SU(3) flavor group are collected in the Appendix. Similar to the SU(3) flavor symmetry/breaking hadronic helicity amplitudes of $D\to P/V/S\ell^{+}\nu_{\ell}$ decays given in Ref. Wang:2022yyn , the hadronic helicity amplitudes of the $D\to A\ell^{+}\nu_{\ell}$ decays can be parameterized as
$H(D\to M\ell^{+}\nu_{\ell})=c_{0}^{M}D_{i}M^{i}_{j}H^{j}+c_{1}^{M}D_{k}W_{i}^{k}M^{i}_{j}H^{j}+c_{2}^{M}D_{i}M_{k}^{i}W_{j}^{k}H^{j},$ (12)
where $H^{2}\equiv V^{*}_{cd}$ and $H^{3}\equiv V^{*}_{cs}$ are the CKM matrix elements, $W$ is the SU(3) flavor breaking related matrix Xu:2013dta ; He:2014xha , $c_{0,1,2}^{A}$ and $c_{0,1,2}^{B}$ are the non-perturbative coefficients corresponding to $M=A$ and $M=B$, respectively, $c_{0}^{A,B}$ are those under the SU(3) flavor symmetry, and $c_{1,2}^{A,B}$ are the SU(3) flavor breaking ones. Note that the Okubo-Zweig-Iizuka suppressed processes, such as $D^{+}_{s}\to b_{1}(1235)^{0}\ell^{+}\nu_{\ell}$ decays studied in Ref. BESIII:2023clm , are not considered in this work. In terms of Eq. (12), the hadronic helicity amplitude relations for the $D\to A\ell^{+}\nu_{\ell}$ decays are summarized in Tab. 1.
From Tab. 1, one can easily obtain the amplitude relations between different decay processes. The CKM matrix elements $V_{cs}$ and $V_{cd}$ are also listed in Tab. 1 for convenient comparison. Since the light axial-vector matrix is distinguished as $A$ and $B$ in the meson multiplets, there are eight non-perturbative parameters $A_{1}$, $A_{2}$, $A_{3}$, $A_{4}$ and $B_{1}$, $B_{2}$, $B_{3}$, $B_{4}$ as defined in the caption of Tab. 1. If we neglect the SU(3) flavor breaking effects of $c^{A}_{1}$, $c^{A}_{2}$, $c^{B}_{1}$, and $c^{B}_{2}$, there are only two non-perturbative parameters $c^{A}_{0}$ and $c^{B}_{0}$. The hadronic amplitudes of various decay processes are connected to the non-perturbative parameters $c^{A}_{0}$ and $c^{B}_{0}$, as well as the mixing angles $\theta_{K_{1}}$, $\theta_{1P_{1}}$ and $\theta_{3P_{1}}$.
Table 1: The hadronic amplitudes of the $D\to A\ell^{+}\nu_{\ell}$ decays. $A_{1}\equiv c^{A}_{0}+c^{A}_{1}-2c^{A}_{2}$, $A_{2}\equiv c^{A}_{0}-2c^{A}_{1}-2c^{A}_{2}$, $A_{3}\equiv c^{A}_{0}+c^{A}_{1}+c^{A}_{2}$, $A_{4}\equiv c^{A}_{0}-2c^{A}_{1}+c^{A}_{2}$. $B_{1}\equiv c^{B}_{0}+c^{B}_{1}-2c^{B}_{2}$, $B_{2}\equiv c^{B}_{0}-2c^{B}_{1}-2c^{B}_{2}$, $B_{3}\equiv c^{B}_{0}+c^{B}_{1}+c^{B}_{2}$, $B_{4}\equiv c^{B}_{0}-2c^{B}_{1}+c^{B}_{2}$. $A_{1}=A_{2}=A_{3}=A_{4}=c^{A}_{0}$ and $B_{1}=B_{2}=B_{3}=B_{4}=c^{B}_{0}$ if neglecting the SU(3) flavor breaking $c^{A}_{1}$, $c^{A}_{2}$, and $c^{B}_{1}$, $c^{B}_{2}$ terms.
    Decay modes SU(3) hadronic amplitudes     Decay modes SU(3) hadronic amplitudes
$D^{0}\to K_{1}(1270)^{-}\ell^{+}\nu_{\ell}$ $(sin\theta_{K_{1}}A_{1}+cos\theta_{K_{1}}B_{1})\leavevmode\nobreak\ V^{*}_{cs}$ $D^{0}\to K_{1}(1400)^{-}\ell^{+}\nu_{\ell}$ $(cos\theta_{K_{1}}A_{1}-sin\theta_{K_{1}}B_{1})\leavevmode\nobreak\ V^{*}_{cs}$
$D^{+}\to\overline{K}_{1}(1270)^{0}\ell^{+}\nu_{\ell}$ $(sin\theta_{K_{1}}A_{1}+cos\theta_{K_{1}}B_{1})\leavevmode\nobreak\ V^{*}_{cs}$ $D^{+}\to\overline{K}_{1}(1400)^{0}\ell^{+}\nu_{\ell}$ $(cos\theta_{K_{1}}A_{1}-sin\theta_{K_{1}}B_{1})\leavevmode\nobreak\ V^{*}_{cs}$
$D^{+}_{s}\to f_{1}(1285)\ell^{+}\nu_{\ell}$ $(\frac{1}{\sqrt{3}}cos\theta_{3P_{1}}-\sqrt{\frac{2}{3}}sin\theta_{3P_{1}})\leavevmode\nobreak\ A_{2}\leavevmode\nobreak\ V^{*}_{cs}$ $D^{+}_{s}\to f_{1}(1420)\ell^{+}\nu_{\ell}$ $(-\frac{1}{\sqrt{3}}sin\theta_{3P_{1}}-\sqrt{\frac{2}{3}}cos\theta_{3P_{1}})\leavevmode\nobreak\ A_{2}\leavevmode\nobreak\ V^{*}_{cs}$
$D^{+}_{s}\to h_{1}(1170)\ell^{+}\nu_{\ell}$ $(\frac{1}{\sqrt{3}}cos\theta_{1P_{1}}-\sqrt{\frac{2}{3}}sin\theta_{1P_{1}})\leavevmode\nobreak\ B_{2}\leavevmode\nobreak\ V^{*}_{cs}$ $D^{+}_{s}\to h_{1}(1415)\ell^{+}\nu_{\ell}$ $(-\frac{1}{\sqrt{3}}sin\theta_{1P_{1}}-\sqrt{\frac{2}{3}}cos\theta_{1P_{1}})\leavevmode\nobreak\ B_{2}\leavevmode\nobreak\ V^{*}_{cs}$
$D^{0}\to a_{1}(1260)^{-}\ell^{+}\nu_{\ell}$ $A_{3}\leavevmode\nobreak\ V^{*}_{cd}$ $D^{0}\to b_{1}(1235)^{-}\ell^{+}\nu_{\ell}$ $B_{3}\leavevmode\nobreak\ V^{*}_{cd}$
$D^{+}\to a_{1}(1260)^{0}\ell^{+}\nu_{\ell}$ $-\frac{1}{\sqrt{2}}\leavevmode\nobreak\ A_{3}\leavevmode\nobreak\ V^{*}_{cd}$ $D^{+}\to b_{1}(1235)^{0}\ell^{+}\nu_{\ell}$ $-\frac{1}{\sqrt{2}}\leavevmode\nobreak\ B_{3}\leavevmode\nobreak\ V^{*}_{cd}$
$D^{+}\to f_{1}(1285)\ell^{+}\nu_{\ell}$ $(\frac{1}{\sqrt{3}}cos\theta_{3P_{1}}+\frac{1}{\sqrt{6}}sin\theta_{3P_{1}})\leavevmode\nobreak\ A_{3}\leavevmode\nobreak\ V^{*}_{cd}$ $D^{+}\to f_{1}(1420)\ell^{+}\nu_{\ell}$ $(-\frac{1}{\sqrt{3}}sin\theta_{3P_{1}}+\frac{1}{\sqrt{6}}cos\theta_{3P_{1}})\leavevmode\nobreak\ A_{3}\leavevmode\nobreak\ V^{*}_{cd}$
$D^{+}\to h_{1}(1170)\ell^{+}\nu_{\ell}$ $(\frac{1}{\sqrt{3}}cos\theta_{1P_{1}}+\frac{1}{\sqrt{6}}sin\theta_{1P_{1}})\leavevmode\nobreak\ B_{3}\leavevmode\nobreak\ V^{*}_{cd}$ $D^{+}\to h_{1}(1415)\ell^{+}\nu_{\ell}$ $(-\frac{1}{\sqrt{3}}sin\theta_{1P_{1}}+\frac{1}{\sqrt{6}}cos\theta_{1P_{1}})\leavevmode\nobreak\ B_{3}\leavevmode\nobreak\ V^{*}_{cd}$
$D^{+}_{s}\to K_{1}(1270)^{0}\ell^{+}\nu_{\ell})$ $(sin\theta_{K_{1}}A_{4}+cos\theta_{K_{1}}B_{4})\leavevmode\nobreak\ V^{*}_{cd}$ $D^{+}_{s}\to K_{1}(1400)^{0}\ell^{+}\nu_{\ell})$ $(cos\theta_{K_{1}}A_{4}-sin\theta_{K_{1}}B_{4})\leavevmode\nobreak\ V^{*}_{cd}$
II.1.2 Numerical results for the $D\to A\ell^{+}\nu_{\ell}$ decays
In our numerical analysis, the theoretical input parameters and experimental data within $1\sigma$ error and within $2\sigma$ errors are taken from the PDG PDG2022 . In the $D\to A\ell^{+}\nu_{\ell}$ decays, only $D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e}$ and $D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e}$ decays have been measured, and the branching ratio of $D^{+}_{s}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e}$ is constrained by an upper limit. Their experimental data with $1\sigma$ error are PDG2022 ; BESIII:2023clm
$\mathcal{B}(D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e})$ $=$ $(1.01\pm 0.18)\times 10^{-3},$ (13)
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e})$ $=$ $(2.30^{+0.40}_{-0.42})\times 10^{-3},$ (14)
[…]
Only two experimental data can not constrain the five parameters $c^{A}_{0}$ and $c^{B}_{0}$, $\theta_{K_{1}}$, $\theta_{1P_{1}}$, and $\theta_{3P_{1}}$. Three mixing angles $\theta_{K_{1}}$, $\theta_{1P_{1}}$, and $\theta_{3_{P_{1}}}$ can also be constrained by the $A\to PV$ non-leptonic decays, which will be discussed in the next section. $\mathcal{B}(D\to a^{-}_{1}/a^{0}_{1}\ell^{+}\nu_{\ell})$ and $\mathcal{B}(D\to b^{-}_{1}/b^{0}_{1}\ell^{+}\nu_{\ell})$ are proportional to $|c^{A}_{0}|^{2}$ and $|c^{B}_{0}|^{2}$, respectively. As given in Refs. Momeni:2019uag ; Momeni:2022gqb , the theoretical predictions of $\mathcal{B}(D\to a^{-}_{1}/a^{0}_{1}\ell^{+}\nu_{l})$ and $\mathcal{B}(D\to b^{-}_{1}/b^{0}_{1}\ell^{+}\nu_{l})$ are quite dependent on different approaches. For the sake of conservatism, we only use $\Big{|}\frac{c^{A}_{0}}{c^{B}_{0}}\Big{|}=\sqrt{\frac{\mathcal{B}(D\to a^{-}_{1}/a^{0}_{1}\ell^{+}\nu_{l})}{\mathcal{B}(D\to b^{-}_{1}/b^{0}_{1}\ell^{+}\nu_{l})}}$ from Refs. Momeni:2019uag ; Momeni:2022gqb , which lies within $[0.88,1.97]$.
Now, using the experimental data for $\mathcal{B}(D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e})$, $\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e})$, and $\mathcal{B}(D^{+}_{s}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e})$ in Eqs. (13-15), the ratio $\Big{|}\frac{c^{A}_{0}}{c^{B}_{0}}\Big{|}\in[0.88,1.97]$ from Refs. Momeni:2019uag ; Momeni:2022gqb , and the relevant constraints from later $\mathcal{B}(A\to PV)$ in Eqs. (21-22) and $\mathcal{B}(D\to Ae^{+}\nu_{e},A\to PV)$ in Eqs. (28-29), in terms of the hadronic amplitude relations listed in Tab. 1, one can constrain relevant parameters and then provide the predictions for $\mathcal{B}(D\to A\ell^{+}\nu_{\ell})$, which have not yet been measured.
We find that the range of $\Big{|}\frac{c^{A}_{0}}{c^{B}_{0}}\Big{|}$ is not reduced by the present experimental data. As for the three mixing angles, their constraints mainly come from the $A\to PV$ decays, and they are all constrained from $[-180^{\circ},180^{\circ}]$ to two allowed ranges as shown in Fig. 1.
[…]
An angle in one allowed range can be found as a supplementary angle in another allowed range, and the branching ratios have similar sensitivity to the angles of the two ranges. Therefore, we only discuss one of two ranges, which are about larger than $0^{\circ}$, and the allowed ranges are
$c^{A}_{0}\in[1.44,3.00],\leavevmode\nobreak\ c^{B}_{0}\in[0.98,2.27],\leavevmode\nobreak\ \theta_{K_{1}}\in[26^{\circ},70^{\circ}],\leavevmode\nobreak\ \theta_{1P_{1}}\in[-15^{\circ},75^{\circ}],\leavevmode\nobreak\ \theta_{3P_{1}}\in[32^{\circ},149^{\circ}]\leavevmode\nobreak\ \mbox{within}\leavevmode\nobreak\ 2\sigma\leavevmode\nobreak\ \mbox{errors},$ (16)
$c^{A}_{0}\in[1.69,2.42],\leavevmode\nobreak\ c^{B}_{0}\in[1.12,2.08],\leavevmode\nobreak\ \theta_{K_{1}}\in[52^{\circ},65^{\circ}],\leavevmode\nobreak\ \theta_{1P_{1}}\in[5^{\circ},56^{\circ}],\leavevmode\nobreak\ \theta_{3P_{1}}\in[56^{\circ},125^{\circ}]\leavevmode\nobreak\ \mbox{within}\leavevmode\nobreak\ 1\sigma\leavevmode\nobreak\ \mbox{error}.\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$ (17)
As previously mentioned, there are several earlier estimations on the mixing angle $\theta_{K_{1}}$. The allowed $\theta_{K_{1}}$ is $[35^{\circ},55^{\circ}]$ in terms of two parameters, the mass difference of the $a_{1}(1260)$ and $b_{1}(1235)$ mesons and the ratio of the constituent quark masses Burakovsky:1997dd ; $33^{\circ}$ and $57^{\circ}$ by $\tau$ decay Suzuki:1993yc , $(34\pm 13)^{\circ}$ by $B\to K_{1}\gamma$ Hatanaka:2008xj ; $33^{\circ}$ is much more favored than $57^{\circ}$ in Ref. Cheng:2013cwa , etc. The allowed ranges of $\theta_{1P_{1}}$ and $\theta_{3P_{1}}$ are obtained, $\theta_{1P_{1}}=(23.3\pm 1)^{\circ}$ or $(38.3\pm 1)^{\circ}$ Cheng:2011pb , $\theta_{3P_{1}}=(4.3\pm 2)^{\circ}$ or $(66.3\pm 2)^{\circ}$ Cheng:2011pb , $\theta_{3_{P_{1}}}=\big{(}19.4^{+4.5}_{-4.6}\big{)}^{\circ}$ or $\big{(}51.1^{+4.5}_{-4.6}\big{)}^{\circ}$ Yang:2010ah . Our bounds within $1\sigma$ error bar favor the large values of $\theta_{K_{1}}$ and $\theta_{3_{P_{1}}}$ in Refs. Cheng:2011pb ; Yang:2010ah .
Table 2: The branching ratio predictions of the $D\to A\ell^{+}\nu_{\ell}$ decays within $1\sigma$ error, and Edenotes experimental data.
Our results $(\ell=e)$ Previous results $(\ell=e)$ Our results $(\ell=\mu)$ Previous results $(\ell=\mu)$
$\mathcal{B}(D^{0}\to K_{1}(1270)^{-}\ell^{+}\nu_{\ell})(\times 10^{-3})$ ${}^{0.93\pm 0.10}_{1.01\pm 0.18^{E}}$cdots $0.81\pm 0.09$cdots
$\mathcal{B}(D^{0}\to K_{1}(1400)^{-}\ell^{+}\nu_{\ell})(\times 10^{-3})$ $0.02\pm 0.02$cdots $0.02\pm 0.02$cdots
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}\ell^{+}\nu_{\ell})(\times 10^{-3})$ ${}^{2.43\pm 0.27}_{2.30^{+0.40\leavevmode\nobreak\ E}_{-0.42}}$ $3.2\pm 0.40$ Cheng:2017pcq $2.10\pm 0.23$ $2.6\pm 0.30$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to\overline{K}_{1}(1400)^{0}\ell^{+}\nu_{\ell})(\times 10^{-3})$ $0.05\pm 0.05$$\{0.005,0.02\}$ Cheng:2017pcq $0.04\pm 0.04$$\{0.004,0.017\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to f_{1}(1285)\ell^{+}\nu_{\ell})(\times 10^{-3})$ $0.86\pm 0.73$$\{0.06,0.36\}$ Cheng:2017pcq $0.76\pm 0.65$$\{0.052,0.306\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to f_{1}(1420)\ell^{+}\nu_{\ell})(\times 10^{-3})$ $0.21\pm 0.21$ $0.25\pm 0.05$ Cheng:2017pcq $0.18\pm 0.18$ $0.21\pm 0.05$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to h_{1}(1170)\ell^{+}\nu_{\ell})(\times 10^{-3})$ $0.26\pm 0.26$$\{0,0.197\}$ Cheng:2017pcq $0.24\pm 0.24$$\{0,0.174\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to h_{1}(1415)\ell^{+}\nu_{\ell})(\times 10^{-3})$ $0.28\pm 0.16$ $0.64\pm 0.07$ Cheng:2017pcq $0.24\pm 0.14$ $0.54\pm 0.06$ Cheng:2017pcq
$\mathcal{B}(D^{0}\to a_{1}(1260)^{-}\ell^{+}\nu_{\ell})(\times 10^{-5})$ $4.46\pm 2.32$ $6.90$ Huang:2021owr , $5.261^{+0.745}_{-0.639}$ Hu:2021lkl $3.93\pm 2.08$ $6.27$ Huang:2021owr , $4.732^{+0.685}_{-0.590}$ Hu:2021lkl
$\mathcal{B}(D^{0}\to b_{1}(1235)^{-}\ell^{+}\nu_{\ell})(\times 10^{-5})$ $2.61\pm 1.44$ $4.85$ Huang:2021owr $2.29\pm 1.26$ $4.40$ Huang:2021owr
$\mathcal{B}(D^{+}\to a_{1}(1260)^{0}\ell^{+}\nu_{\ell})(\times 10^{-5})$ $5.79\pm 3.00$ $9.38$ Huang:2021owr , $6.673^{+0.947}_{-0.811}$ Hu:2021lkl $5.11\pm 2.70$ $8.52$ Huang:2021owr , $6.002^{+0.796}_{-0.748}$ Hu:2021lkl
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}\ell^{+}\nu_{\ell})(\times 10^{-5})$ $3.41\pm 1.88$ $7.4\pm 0.70$ Cheng:2017pcq , $6.58$ Huang:2021owr $2.99\pm 1.65$ $6.4\pm 0.6$ Cheng:2017pcq , $6.00$ Huang:2021owr
$\mathcal{B}(D^{+}\to f_{1}(1285)\ell^{+}\nu_{\ell})(\times 10^{-5})$ $1.88\pm 1.88$ $3.7\pm 0.80$ Cheng:2017pcq $1.61\pm 1.61$ $3.2\pm 0.6$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to f_{1}(1420)\ell^{+}\nu_{\ell})(\times 10^{-5})$ $0.64\pm 0.54$$\{0.02,0.14\}$ Cheng:2017pcq $0.48\pm 0.41$$\{0.02,0.12\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to h_{1}(1170)\ell^{+}\nu_{\ell})(\times 10^{-5})$ $5.28\pm 3.00$ $14\pm 1.50$ Cheng:2017pcq $4.73\pm 2.69$ $12.2\pm 1.3$ Cheng:2017pcq
$\mathcal{B}(D^{+}\to h_{1}(1415)\ell^{+}\nu_{\ell})(\times 10^{-5})$ $0.11\pm 0.11$$\{0,0.02\}$ Cheng:2017pcq $0.08\pm 0.08$$\{0,0.02\}$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to K_{1}(1270)^{0}\ell^{+}\nu_{\ell})(\times 10^{-5})$ <${}^{11.77\pm 1.39}_{<41^{E}}$ $17\pm 2.00$ Cheng:2017pcq $10.57\pm 1.25$ $15\pm 2$ Cheng:2017pcq
$\mathcal{B}(D^{+}_{s}\to K_{1}(1400)^{0}\ell^{+}\nu_{\ell})(\times 10^{-5})$ $0.32\pm 0.32$$\{0.05,0.14\}$ Cheng:2017pcq $0.27\pm 0.27$$\{0.05,0.12\}$ Cheng:2017pcq
[…]
In addition, we show the sensitivity of the branching ratios to the non-perturbative parameters $A$ and $B$ as well as the mixing angles $\theta_{K_{1}}$, $\theta_{3P_{1}}$, and $\theta_{1P_{1}}$ in Figs. 3 and 3. We give the remarks as follows.
[…]
Figure 3: Branching ratios of the $D_{(s)}\to a_{1}/b_{1}/f_{1}/h_{1}e^{+}\nu_{e}$ decays.
  • As given in Tab. 1, the hadronic amplitudes of the $D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e}$, $D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e}$, and $D^{+}_{s}\to{K}_{1}(1270)^{0}e^{+}\nu_{e}$ decays are proportional to $sin\theta_{K_{1}}c^{A}_{0}+cos\theta_{K_{1}}c^{B}_{0}$, and the hadronic amplitudes of the $D^{0}\to K_{1}(1400)^{-}e^{+}\nu_{e}$, $D^{+}\to\overline{K}_{1}(1400)^{0}e^{+}\nu_{e}$, and $D^{+}_{s}\to{K}_{1}(1400)^{0}e^{+}\nu_{e}$ decays are proportional to $cos\theta_{K_{1}}c^{A}_{0}-sin\theta_{K_{1}}c^{B}_{0}$. Thus as shown in Fig. 3, three $D\to K_{1}(1270)e^{+}\nu_{e}$ decays or three $D\to{K}_{1}(1400)e^{+}\nu_{e}$ decays have similar variation trends as the three parameters $c^{A}_{0}$, $c^{B}_{0}$, or $\theta_{K_{1}}$. From Fig. 3 ($a_{1,2,3}$) and ($b_{1,2,3}$), one can see that the experimental lower limit of $\mathcal{B}(D^{0}\to K_{1}(1270)^{-}e^{+}\nu_{e})$ and the experimental upper limit of $\mathcal{B}(D^{+}\to\overline{K}_{1}(1270)^{0}e^{+}\nu_{e})$ give effective constraints on our predictions. Three $\mathcal{B}(D\to K_{1}(1270)e^{+}\nu_{e})$ in Fig. 3 ($a_{1,2,3}$), ($b_{1,2,3}$), and ($c_{1,2,3}$) have some sensitivity to $A$ and $B$. As shown in Fig. 3 ($d_{1,2,3}$), ($e_{1,2,3}$), and ($f_{1,2,3}$), three $\mathcal{B}(D\to{K}_{1}(1400)e^{+}\nu_{e})$ are very sensitive to three parameters $A$, $B$, and $\theta_{K_{1}}$.
  • As displayed in Fig. 3 ($a_{1-4}$), $\mathcal{B}(D\to a_{1}(1260)e^{+}\nu_{e})$ and $D\to b_{1}(1235)e^{+}\nu_{e})$ are only related to $A$ and $B$, respectively, and they are very sensitive to $A$ or $B$. Fig. 3 ($b_{1-4}$) and ($c_{1-4}$) show the $\mathcal{B}(D_{(s)}\to f_{1}(1285)/f_{1}(1420)e^{+}\nu_{e})$ which are related to $A$ and $\theta_{3P_{1}}$. Four $\mathcal{B}(D_{(s)}\to f_{1}(1285)/f_{1}(1420)e^{+}\nu_{e})$ have similar sensitivity to $A$, but the change trends of four branch ratios to $\theta_{3P_{1}}$ are different. Fig. 3 ($d_{1-4}$) and ($e_{1-4}$) show the $\mathcal{B}(D_{(s)}\to h_{1}(1170)/h_{1}(1415)e^{+}\nu_{e})$, which are related to $B$ and $\theta_{1P_{1}}$. $\mathcal{B}(D^{+}\to h_{1}(1170)e^{+}\nu_{e})$ and $\mathcal{B}(D^{+}_{s}\to h_{1}(1415)e^{+}\nu_{e})$ are sensitive to parameter $B$. $\mathcal{B}(D^{+}_{s}\to h_{1}(1170)e^{+}\nu_{e})$ and $\mathcal{B}(D^{+}\to h_{1}(1415)e^{+}\nu_{e})$ are sensitive to $\theta_{1P_{1}}$, and Fig. 3 ($d_{2}$) and ($e_{4}$) show that the branching ratios first decrease and then increase with $\theta_{1P_{1}}$, reaching minimum values when $\theta_{1P_{1}}$ is about $36^{\circ}$.
II.2 Non-leptonic $A\to VP$ decays
II.2.1 Theoretical framework for the $A\to VP$ decays
For the non-leptonic decays of axial-vector mesons $A\to VP$, the decay branching ratios can be written as Roca:2003uk ; Parui:2022zfm
$\mathcal{B}(A\to VP)=\frac{\tau_{A}|\lambda_{AVP}|^{2}}{2\pi m_{A}^{2}}q^{\prime}\left(1+\frac{2}{3}\frac{{q^{\prime}}^{2}}{m_{V}^{2}}\right).$ (18)
where $q^{\prime}\equiv\frac{1}{2m_{A}}\lambda^{1/2}(m_{A}^{2},m_{V}^{2},m_{P}^{2})$ is the momentum of the final state particles in the rest frame of the axial-vector meson, and $\lambda_{AVP}$ are the coefficients of the $AVP$ vertex. $\lambda_{AVP}$ can be obtained from the effective Hamiltonian in Eq. (1) of Ref. Roca:2003uk or the SU(3) flavor symmetry/breaking; we will use the latter to obtain $\lambda_{AVP}$ in this work.
There are small phase spaces for some axial-vector meson decays, such as $K_{1}(1270)\to\rho K/\omega K/\phi K/K^{*}\eta$, $K_{1}(1400)\to\phi K/\phi\eta$, $a_{1}(1260)\to K^{*}K$, $b_{1}(1235)\to K^{*}K/\rho\eta$, $f_{1}(1285)\to K^{*}K$, $h_{1}(1170)\to K^{*}K/\omega\eta/\phi\eta$, and $h_{1}(1415)\to\phi\eta$. The mass distributions of the resonances have been considered in Refs. Wang:2022vga ; Roca:2021bxk , and the mass distributions of the resonances and vector mesons have been considered in Ref. Roca:2003uk . We will follow Ref. Roca:2003uk to obtain our numerical results. The decay branching ratios with the mass distributions of the resonances and vector mesons can be written as Roca:2003uk
[…]
×$\times\mathcal{B}_{AVP}(\sqrt{s_{A}},\sqrt{s_{V}})\Theta(\sqrt{s_{A}}-\sqrt{s_{V}}-m_{P})$
[…]
Now we relate the $AVP$ vertex coefficients $\lambda_{AVP}$ for different decay modes by the SU(3) flavor symmetry/breaking. For $J^{PC}=1^{++}$ axial-vector mesons decays, the $AVP$ vertex coefficients can be parameterized as
$\lambda_{AVP}=c^{\prime A}_{0}A_{j}^{i}M^{j}_{k}M^{\prime k}_{i}+c^{\prime A}_{1}A_{j}^{a}W^{i}_{a}M^{j}_{k}M^{\prime k}_{i}+c^{\prime A}_{2}A_{j}^{i}M^{a}_{k}W^{j}_{a}M^{\prime k}_{i}+c^{\prime A}_{3}A_{j}^{i}M^{j}_{k}M^{\prime a}_{i}W^{k}_{a},$ (20)
where $M^{(^{\prime})}=V$ or $P$, $c^{\prime A}_{0}$ are the non-perturbative coefficients under the SU(3) flavor symmetry, and $c^{\prime A}_{1,2,3}$ are the non-perturbative SU(3) flavor breaking coefficients. The vertex coefficients for the $J^{PC}=1^{+-}$ axial-vector mesons decays can be obtained by replacing $A$ to $B$ in Eq. (20). To keep the results invariant under the conjugate change, $c^{\prime A/B}_{1}=c^{\prime A/B}_{2}$ are used in our analysis. In terms of Eq. (20), the vertex coefficient relations for the $A/B\to VP$ decays are summarized in Tabs. 4-5. Although there is not enough phase space for the $K_{1}(1270)\to\phi K$ decay, we still give its vertex coefficient in Tab. 4, because we will use it to obtain one for $K_{1}(1400)\to\phi K$. In Tabs. 4-5, we redefine $F_{1,2,3,4,5}$ and $D_{1,2,3,4,5,6}$ by $c^{\prime A}_{0,1,3}$ and $c^{\prime B}_{0,1,3}$, respectively. If we ignore the SU(3) flavor breaking effects, then we have $F_{1}=F_{2}=F_{3}=F_{4}=F_{5}=F$ and $D_{1}=D_{2}=D_{3}=D_{4}=D_{5}=D_{6}=D$, where $F=c^{A}_{0}$ and $D=c^{B}_{0}$. After ignoring the SU(3) flavor breaking effects, most of our vertex coefficients in Tabs. 4-5 are in agreement with those in Ref. Roca:2003uk , except for some with different meson definitions, $i.e.$, $\eta,\eta^{\prime},h_{1},f_{1}$ mesons. Note that there is an overall negative difference for some vertex coefficients due to the different definitions of $D$ and $F$, and this does not affect the branching ratio results.
Table 3: The vertex coefficients of $K_{1}(1270)\to VP$ decays. $F_{1}=c^{\prime A}_{0}-c^{\prime A}_{1}+c^{\prime A}_{3},\leavevmode\nobreak\ D_{1}=c^{\prime B}_{0}-c^{\prime B}_{1}+c^{\prime B}_{3}$, $F_{2}=c^{\prime A}_{0}-c^{\prime A}_{1}-2c^{\prime A}_{3},\leavevmode\nobreak\ D_{2}=c^{\prime B}_{0}-c^{\prime B}_{1}-2c^{\prime B}_{3}$. Ones of $K_{1}(1400)\to VP$ decays could be obtained from $K_{1}(1270)\to VP$ decays by replacing $sin\theta_{K_{1}}\to cos\theta_{K_{1}}$ and $cos\theta_{K_{1}}\to-sin\theta_{K_{1}}$.
   Decay modes Vertex coefficients $\lambda_{AVP}$
$K_{1}(1270)^{-}\to K^{*-}\pi^{0}$ $\frac{1}{\sqrt{2}}\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$K_{1}(1270)^{-}\to K^{*-}\eta$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}\big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\big{)}+\big{(}F_{2}sin\theta_{K_{1}}-D_{2}cos\theta_{K_{1}}\big{)}\big{(}\frac{2cos\theta_{P}}{\sqrt{6}}+\frac{sin\theta_{P}}{\sqrt{3}}\big{)}$
$K_{1}(1270)^{-}\to K^{*-}\eta^{\prime}$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}\big{(}\frac{sin\theta_{P}}{\sqrt{6}}+\frac{cos\theta_{P}}{\sqrt{3}}\big{)}+\big{(}F_{2}sin\theta_{K_{1}}-D_{2}cos\theta_{K_{1}}\big{)}\big{(}\frac{2sin\theta_{P}}{\sqrt{6}}-\frac{cos\theta_{P}}{\sqrt{3}}\big{)}$
$K_{1}(1270)^{-}\to\rho^{0}K^{-}$ $\frac{1}{\sqrt{2}}\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$K_{1}(1270)^{-}\to\omega K^{-}$ $\frac{1}{\sqrt{2}}\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$K_{1}(1270)^{-}\to\overline{K}^{*0}\pi^{-}$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$K_{1}(1270)^{-}\to\rho^{-}\overline{K}^{0}$ $\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$K_{1}(1270)^{-}\to\phi K^{-}$ $\big{(}F_{2}sin\theta_{K_{1}}+D_{2}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to K^{*-}\pi^{+}$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\rho^{+}K^{-}$ $\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\pi^{0}$ $-\frac{1}{\sqrt{2}}\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\eta$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}\big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\big{)}+\big{(}F_{2}sin\theta_{K_{1}}-D_{2}cos\theta_{K_{1}}\big{)}\big{(}\frac{2cos\theta_{P}}{\sqrt{6}}+\frac{sin\theta_{P}}{\sqrt{3}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\eta^{\prime}$ $\big{(}F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}\big{(}\frac{sin\theta_{P}}{\sqrt{6}}+\frac{cos\theta_{P}}{\sqrt{3}}\big{)}+\big{(}F_{2}sin\theta_{K_{1}}-D_{2}cos\theta_{K_{1}}\big{)}\big{(}\frac{2sin\theta_{P}}{\sqrt{6}}-\frac{cos\theta_{P}}{\sqrt{3}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\rho^{0}\overline{K}^{0}$ $-\frac{1}{\sqrt{2}}\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\omega\overline{K}^{0}$ $\frac{1}{\sqrt{2}}\big{(}-F_{1}sin\theta_{K_{1}}+D_{1}cos\theta_{K_{1}}\big{)}$
$\overline{K}_{1}(1270)^{0}\to\phi\overline{K}^{0}$ $\big{(}F_{2}sin\theta_{K_{1}}+D_{2}cos\theta_{K_{1}}\big{)}$
Table 4: The vertex coefficients of the $a_{1}(1260)/b_{1}(1235)\to VP$ decays. $F_{3}=c^{\prime A}_{0}+2c^{\prime A}_{1}+c^{\prime A}_{3}$, $F_{4}=c^{\prime A}_{0}+2c^{\prime A}_{1}-2c^{\prime A}_{3}$, $D_{3}=c^{\prime B}_{0}+2c^{\prime B}_{1}+c^{\prime B}_{3}$, $D_{4}=c^{\prime B}_{0}+2c^{\prime B}_{1}-2c^{\prime B}_{3}$.
   Decay modes Vertex coefficients $\lambda_{AVP}$    Decay modes Vertex coefficients $\lambda_{AVP}$
$a_{1}(1260)^{0}\to\rho^{+}\pi^{-}$ $\sqrt{2}F_{3}$ $b_{1}(1235)^{0}\to\rho^{0}\eta$ $2D_{3}\big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\big{)}$
$a_{1}(1260)^{0}\to\rho^{-}\pi^{+}$ $-\sqrt{2}F_{3}$ $b_{1}(1235)^{0}\to\rho^{0}\eta^{\prime}$ $2D_{3}\big{(}\frac{sin\theta_{P}}{\sqrt{6}}+\frac{cos\theta_{P}}{\sqrt{3}}\big{)}$
$a_{1}(1260)^{0}\to K^{*0}\overline{K}^{0}$ $-\frac{1}{\sqrt{2}}F_{4}$ $b_{1}(1235)^{0}\to K^{*0}\overline{K}^{0}$ $-\frac{1}{\sqrt{2}}D_{4}$
$a_{1}(1260)^{0}\to\overline{K}^{*0}K^{0}$ $\frac{1}{\sqrt{2}}F_{4}$ $b_{1}(1235)^{0}\to\overline{K}^{*0}K^{0}$ $-\frac{1}{\sqrt{2}}D_{4}$
$a_{1}(1260)^{0}\to K^{*+}K^{-}$ $\frac{1}{\sqrt{2}}F_{4}$ $b_{1}(1235)^{0}\to K^{*+}K^{-}$ $\frac{1}{\sqrt{2}}D_{4}$
$a_{1}(1260)^{0}\to K^{*-}K^{+}$ $-\frac{1}{\sqrt{2}}F_{4}$ $b_{1}(1235)^{0}\to K^{*-}K^{+}$ $\frac{1}{\sqrt{2}}D_{4}$
$b_{1}(1235)^{0}\to\omega\pi^{0}$ $\sqrt{2}D_{3}$
$a_{1}(1260)^{-}\to\rho^{-}\pi^{0}$ $\sqrt{2}F_{3}$ $b_{1}(1235)^{-}\to\rho^{-}\eta$ $2D_{3}\big{(}\frac{cos\theta_{P}}{\sqrt{6}}-\frac{sin\theta_{P}}{\sqrt{3}}\big{)}$
$a_{1}(1260)^{-}\to\rho^{0}\pi^{-}$ $-\sqrt{2}F_{3}$ $b_{1}(1235)^{-}\to\rho^{-}\eta^{\prime}$ $2D_{3}\big{(}\frac{sin\theta_{P}}{\sqrt{6}}+\frac{cos\theta_{P}}{\sqrt{3}}\big{)}$
$a_{1}(1260)^{-}\to K^{*0}K^{-}$ $F_{4}$ $b_{1}(1235)^{-}\to K^{*0}K^{-}$ $D_{4}$
$a_{1}(1260)^{-}\to K^{*-}K^{0}$ $-F_{4}$ $b_{1}(1235)^{-}\to K^{*-}K^{0}$ $D_{4}$
$b_{1}(1235)^{-}\to\omega\pi^{-}$ $\sqrt{2}D_{3}$
Table 5: The vertex coefficients of the $h_{1}(1170)/f_{1}(1285)\to VP$ decays. $F_{5}=c^{\prime A}_{0}-4c^{\prime A}_{1}+c^{\prime A}_{3}$, $D_{5}=c^{\prime B}_{0}-4c^{\prime B}_{1}+c^{\prime B}_{3}$, $D_{6}=c^{\prime B}_{0}-4c^{\prime B}_{1}-2c^{\prime B}_{3}$. Ones of the $h_{1}(1415)\to VP$ decays could be obtained from the $h_{1}(1170)\to VP$ decays by replacing $sin\theta_{1P_{1}}\to cos\theta_{1P_{1}}$ and $cos\theta_{1P_{1}}\to-sin\theta_{1P_{1}}$. Ones of the $f_{1}(1420)\to VP$ decays could be obtained from the $f_{1}(1285)\to VP$ decays by replacing $sin\theta_{3P_{1}}\to cos\theta_{3P_{1}}$ and $cos\theta_{3P_{1}}\to-sin\theta_{3P_{1}}$.
          Decay modes        Vertex coefficients $\lambda_{AVP}$
       $h_{1}(1170)\to\rho^{0}\pi^{0}$        $\frac{2}{\sqrt{6}}D_{3}\big{(}\sqrt{2}cos\theta_{1P_{1}}+sin\theta_{1P_{1}}\big{)}$
       $h_{1}(1170)\to\omega\eta$        $\frac{\sqrt{2}}{3}D_{3}\big{(}\sqrt{2}cos\theta_{1P_{1}}+sin\theta_{1P_{1}}\big{)}\big{(}cos\theta_{P}-\sqrt{2}sin\theta_{P}\big{)}$
       $h_{1}(1170)\to\omega\eta^{\prime}$        $\frac{\sqrt{2}}{3}D_{3}\big{(}\sqrt{2}cos\theta_{1P_{1}}+sin\theta_{1P_{1}}\big{)}\big{(}sin\theta_{P}+\sqrt{2}cos\theta_{P}\big{)}$
       $h_{1}(1170)\to\rho^{+}\pi^{-}$        $\frac{2}{\sqrt{6}}D_{3}\big{(}\sqrt{2}cos\theta_{1P_{1}}+sin\theta_{1P_{1}}\big{)}$
       $h_{1}(1170)\to\rho^{-}\pi^{+}$        $\frac{2}{\sqrt{6}}D_{3}\big{(}\sqrt{2}cos\theta_{1P_{1}}+sin\theta_{1P_{1}}\big{)}$
       $h_{1}(1170)\to K^{*+}K^{-}$        $\frac{cos\theta_{1P_{1}}}{\sqrt{3}}\big{(}D_{4}+D_{5}\big{)}+\frac{sin\theta_{1P_{1}}}{\sqrt{6}}\big{(}D_{4}-2D_{5}\big{)}$
       $h_{1}(1170)\to K^{*-}K^{+}$        $\frac{cos\theta_{1P_{1}}}{\sqrt{3}}\big{(}D_{4}+D_{5}\big{)}+\frac{sin\theta_{1P_{1}}}{\sqrt{6}}\big{(}D_{4}-2D_{5}\big{)}$
       $h_{1}(1170)\to K^{*0}\overline{K}^{0}$        $\frac{cos\theta_{1P_{1}}}{\sqrt{3}}\big{(}D_{4}+D_{5}\big{)}+\frac{sin\theta_{1P_{1}}}{\sqrt{6}}\big{(}D_{4}-2D_{5}\big{)}$
       $h_{1}(1170)\to\overline{K}^{*0}K^{0}$        $\frac{cos\theta_{1P_{1}}}{\sqrt{3}}\big{(}D_{4}+D_{5}\big{)}+\frac{sin\theta_{1P_{1}}}{\sqrt{6}}\big{(}D_{4}-2D_{5}\big{)}$
       $h_{1}(1170)\to\phi\eta$        $\frac{2}{3}D_{6}\big{(}-cos\theta_{1P_{1}}+\sqrt{2}sin\theta_{1P_{1}}\big{)}\big{(}\sqrt{2}cos\theta_{P}+sin\theta_{P}\big{)}$
       $h_{1}(1170)\to\phi\eta^{\prime}$        $\frac{2}{3}D_{6}\big{(}-cos\theta_{1P_{1}}+\sqrt{2}sin\theta_{1P_{1}}\big{)}\big{(}\sqrt{2}sin\theta_{P}-cos\theta_{P}\big{)}$
       $f_{1}(1285)\to K^{*+}K^{-}$        $\frac{cos\theta_{3P_{1}}}{\sqrt{3}}\big{(}F_{4}-F_{5}\big{)}+\frac{sin\theta_{3P_{1}}}{\sqrt{6}}\big{(}F_{4}+2F_{5}\big{)}$
       $f_{1}(1285)\to K^{*-}K^{+}$        $-\frac{cos\theta_{3P_{1}}}{\sqrt{3}}\big{(}F_{4}-F_{5}\big{)}-\frac{sin\theta_{3P_{1}}}{\sqrt{6}}\big{(}F_{4}+2F_{5}\big{)}$
       $f_{1}(1285)\to K^{*0}\overline{K}^{0}$        $\frac{cos\theta_{3P_{1}}}{\sqrt{3}}\big{(}F_{4}-F_{5}\big{)}+\frac{sin\theta_{3P_{1}}}{\sqrt{6}}\big{(}F_{4}+2F_{5}\big{)}$
       $f_{1}(1285)\to\overline{K}^{*0}K^{0}$        $-\frac{cos\theta_{3P_{1}}}{\sqrt{3}}\big{(}F_{4}-F_{5}\big{)}-\frac{sin\theta_{3P_{1}}}{\sqrt{6}}\big{(}F_{4}+2F_{5}\big{)}$
II.2.2 Numerical results for the $A\to VP$ decays
In the $A\to VP$ decays, only a few branching ratios of $K_{1}(1270)/K_{1}(1400)\to VP$ have been measured PDG2022
$\mathcal{B}(K_{1}(1270)\to\rho K)=(38\pm 13)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(K_{1}(1270)\to\omega K)=(11\pm 2)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(K_{1}(1270)\to K^{*}\pi)=(21\pm 10)\%,$ (21)
$\mathcal{B}(K_{1}(1400)\to\omega K)=(1\pm 1)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(K_{1}(1400)\to\rho K)=(3\pm 3)\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(K_{1}(1400)\to K^{*}\pi)=(94\pm 6)\%.$ (22)
In addition, the experimental errors of $\mathcal{B}(K_{1}(1400)\to\rho K)$, $\mathcal{B}(K_{1}(1400)\to\omega K)$ and $\mathcal{B}(K_{1}(1270)\to K^{*}(892)\pi)$ are quite large. Given the insufficient experimental data and consistency with our study of semileptonic decay $D\to A\ell^{+}\nu_{\ell}$, we only consider the SU(3) flavor symmetry for the $A\to VP$ numerical results.
Some values of $D$ and $F$ have been obtained from $A\to VP$ decays Roca:2003uk . Considering the results in Ref. Roca:2003uk , we conservatively choose $F\in[0.6,2.0]$ and $D\in[-1.6,-0.4]$ as the original ranges. Then, using the experimental data of $\mathcal{B}(K_{1}\to VP)$ in Eqs. (21-22), $\mathcal{B}(a_{1}(1260)/b_{1}(1235)/h_{1}(1170)/$ $h_{1}(1415)/f_{1}(1285)/f_{1}(1420)\to VP)\leq 1$, and also considering an extra constraint from $\mathcal{B}(D\to K_{1}(1270)e^{+}\nu_{e})$ in Eqs. (13-15) and $\mathcal{B}(D\to Ae^{+}\nu_{e},A\to PV)$ in Eqs. (28-29), one can constrain the relevant parameters $F$, $D$, and the mixing angles $\theta_{K_{1}}$, $\theta_{3P_{1}}$, and $\theta_{1P_{1}}$. Three constrained mixing angles are given in Eqs. (16-17), and the constrained $F$ and $D$ are
[…]
Using the relevant constrained parameters, we give the predictions of $\mathcal{B}(A\to VP)$, which have not been measured to date. The numerical results are given in Tab. 6.
Table 6: The branching ratio predictions of the $K_{1}/a_{1}/b_{1}/f_{1}/h_{1}\to VP$ decays by the SU(3) flavor symmetry with $1\sigma$ error. The unit is $10^{-2}$ for all branching ratios.
      Branching ratios       Our predictions       Our predictions
      $\mathcal{B}(K_{1}(1270)^{-}\to K^{*-}\pi^{0})$       $7.24\pm 3.17$       $\mathcal{B}(K_{1}(1400)^{-}\to K^{*-}\pi^{0})$       $30.54\pm 1.01$
      $\mathcal{B}(K_{1}(1270)^{-}\to K^{*-}\eta)$       $0.50\pm 0.16$       $\mathcal{B}(K_{1}(1400)^{-}\to K^{*-}\eta)$       $2.81\pm 1.84$
      $\mathcal{B}(K_{1}(1270)^{-}\to\rho^{0}K^{-})$       $16.82\pm 0.49$       $\mathcal{B}(K_{1}(1400)^{-}\to\rho^{0}K^{-})$       $0.75\pm 0.75$
      $\mathcal{B}(K_{1}(1270)^{-}\to\omega K^{-})$       $9.52\pm 0.21$       $\mathcal{B}(K_{1}(1400)^{-}\to\omega K^{-})$       $0.73\pm 0.73$
      $\mathcal{B}(K_{1}(1270)^{-}\to\overline{K}^{*0}\pi^{-})$       $14.20\pm 6.22$       $\mathcal{B}(K_{1}(1400)^{-}\to\overline{K}^{*0}\pi^{-})$       $60.32\pm 1.97$
      $\mathcal{B}(K_{1}(1270)^{-}\to\rho^{-}\overline{K}^{0})$       $32.71\pm 0.82$       $\mathcal{B}(K_{1}(1400)^{-}\to\rho^{-}\overline{K}^{0})$       $1.48\pm 1.48$
      $\mathcal{B}(\overline{K}_{1}(1270)^{0}\to K^{*-}\pi^{+})$       $14.39\pm 6.31$       $\mathcal{B}(K_{1}(1400)^{-}\to\phi K^{-})$       $3.81\pm 0.65$
      $\mathcal{B}(\overline{K}_{1}(1270)^{0}\to\rho^{+}K^{-})$       $33.91\pm 0.86$       $\mathcal{B}(\overline{K}_{1}(1400)^{0}\to K^{*-}\pi^{+})$       $60.88\pm 2.01$
      $\mathcal{B}(\overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\pi^{0})$       $7.15\pm 3.13$       $\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\rho^{+}K^{-})$       $1.51\pm 1.51$
      $\mathcal{B}(\overline{K}_{1}(1270)^{0}\to\overline{K}^{*0}\eta)$       $0.41\pm 0.14$       $\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\overline{K}^{*0}\pi^{0})$       $30.27\pm 0.99$
      $\mathcal{B}(\overline{K}_{1}(1270)^{0}\to\rho^{0}\overline{K}^{0})$       $16.22\pm 0.47$       $\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\overline{K}^{*0}\eta)$       $2.67\pm 1.75$
      $\mathcal{B}(\overline{K}_{1}(1270)^{0}\to\omega\overline{K}^{0})$       $8.87\pm 0.20$       $\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\rho^{0}\overline{K}^{0})$       $0.74\pm 0.74$
      $\mathcal{B}(K_{1}(1270)\to K^{*}\pi)$       $21.48\pm 9.42$       $\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\omega\overline{K}^{0})$       $0.71\pm 0.71$
      $\mathcal{B}(K_{1}(1270)\to\rho K)$       $49.78\pm 1.21$       $\mathcal{B}(\overline{K}_{1}(1400)^{0}\to\phi\overline{K}^{0})$       $3.64\pm 0.63$
      $\mathcal{B}(K_{1}(1270)\to\omega K)$       $9.20\pm 0.20$       $\mathcal{B}(K_{1}(1400)\to K^{*}\pi)$       $91.01\pm 2.99$
      $\mathcal{B}(K_{1}(1270)\to K^{*}\eta)$       $0.45\pm 0.15$       $\mathcal{B}(K_{1}(1400)\to\rho K)$       $2.24\pm 2.24$
      $\mathcal{B}(K_{1}(1400)\to\omega K)$       $0.72\pm 0.72$
      $\mathcal{B}(K_{1}(1400)\to K^{*}\eta)$       $2.74\pm 1.79$
      $\mathcal{B}(K_{1}(1400)\to\phi K)$       $3.73\pm 0.64$
      $\mathcal{B}(a_{1}(1260)^{0}\to\rho^{+}\pi^{-})$       $38.15\pm 8.46$       $\mathcal{B}(b_{1}(1235)^{0}\to\rho^{0}\eta)$       $5.07\pm 2.14$
      $\mathcal{B}(a_{1}(1260)^{0}\to\rho^{-}\pi^{+})$       $38.15\pm 8.46$       $\mathcal{B}(b_{1}(1235)^{0}\to\omega\pi^{0})$       $65.46\pm 22.46$
      $\mathcal{B}(a_{1}(1260)^{0}\to K^{*0}\overline{K}^{0})$       $1.82\pm 0.52$       $\mathcal{B}(b_{1}(1235)^{0}\to K^{*0}\overline{K}^{0})$       $0.37\pm 0.15$
      $\mathcal{B}(a_{1}(1260)^{0}\to\overline{K}^{*0}K^{0})$       $1.82\pm 0.52$       $\mathcal{B}(b_{1}(1235)^{0}\to\overline{K}^{*0}K^{0})$       $0.37\pm 0.15$
      $\mathcal{B}(a_{1}(1260)^{0}\to K^{*+}K^{-})$       $1.89\pm 0.53$       $\mathcal{B}(b_{1}(1235)^{0}\to K^{*+}K^{-})$       $0.42\pm 0.17$
      $\mathcal{B}(a_{1}(1260)^{0}\to K^{*-}K^{+})$       $1.89\pm 0.53$       $\mathcal{B}(b_{1}(1235)^{0}\to K^{*-}K^{+})$       $0.42\pm 0.17$
      $\mathcal{B}(a_{1}(1260)^{-}\to\rho^{-}\pi^{0})$       $38.33\pm 8.51$       $\mathcal{B}(b_{1}(1235)^{-}\to\rho^{-}\eta)$       $5.13\pm 2.16$
      $\mathcal{B}(a_{1}(1260)^{-}\to\rho^{0}\pi^{-})$       $38.13\pm 8.45$       $\mathcal{B}(b_{1}(1235)^{-}\to\omega\pi^{-})$       $65.17\pm 22.36$
      $\mathcal{B}(a_{1}(1260)^{-}\to K^{*0}K^{-})$       $3.70\pm 1.04$       $\mathcal{B}(b_{1}(1235)^{-}\to K^{*0}K^{-})$       $0.78\pm 0.32$
      $\mathcal{B}(a_{1}(1260)^{-}\to K^{*-}K^{0})$       $3.72\pm 1.04$       $\mathcal{B}(b_{1}(1235)^{-}\to K^{*-}K^{0})$       $0.79\pm 0.33$
      $\mathcal{B}(a_{1}(1260)\to\rho\pi)$       $76.38\pm 16.93$       $\mathcal{B}(b_{1}(1235)\to\rho\eta)$       $5.10\pm 2.15$
      $\mathcal{B}(a_{1}(1260)\to K^{*}K)$       $7.42\pm 2.09$       $\mathcal{B}(b_{1}(1235)\to\omega\pi)$       $65.31\pm 22.41$
      $\mathcal{B}(b_{1}(1235)\to K^{*}K)$       $1.58\pm 0.65$
      $\mathcal{B}(f_{1}(1285)\to K^{*+}K^{-})$       $0.62\pm 0.24$       $\mathcal{B}(f_{1}(1420)\to K^{*+}K^{-})$       $13.09\pm 13.09$
      $\mathcal{B}(f_{1}(1285)\to K^{*-}K^{+})$       $0.62\pm 0.24$       $\mathcal{B}(f_{1}(1420)\to K^{*-}K^{+})$       $13.09\pm 13.09$
      $\mathcal{B}(f_{1}(1285)\to K^{*0}\overline{K}^{0})$       $0.13\pm 0.05$       $\mathcal{B}(f_{1}(1420)\to K^{*0}\overline{K}^{0})$       $11.75\pm 11.75$
      $\mathcal{B}(f_{1}(1285)\to\overline{K}^{*0}K^{0})$       $0.13\pm 0.05$       $\mathcal{B}(f_{1}(1420)\to\overline{K}^{*0}K^{0})$       $11.75\pm 11.75$
      $\mathcal{B}(f_{1}(1285)\to{K}K^{*})$       $1.49\pm 0.59$       $\mathcal{B}(f_{1}(1420)\to K^{*}K)$       $49.64\pm 49.64$
      $\mathcal{B}(h_{1}(1170)\to\rho^{0}\pi^{0})$       $21.84\pm 9.33$       $\mathcal{B}(h_{1}(1415)\to\rho^{0}\pi^{0})$       $13.20\pm 13.20$
      $\mathcal{B}(h_{1}(1170)\to\omega\eta)$       $2.01\pm 1.04$       $\mathcal{B}(h_{1}(1415)\to\omega\eta)$       $3.37\pm 3.37$
      $\mathcal{B}(h_{1}(1170)\to\rho^{+}\pi^{-})$       $21.74\pm 9.29$       $\mathcal{B}(h_{1}(1415)\to\rho^{+}\pi^{-})$       $13.17\pm 13.17$
      $\mathcal{B}(h_{1}(1170)\to\rho^{-}\pi^{+})$       $21.74\pm 9.29$       $\mathcal{B}(h_{1}(1415)\to\rho^{-}\pi^{+})$       $13.17\pm 13.17$
      $\mathcal{B}(h_{1}(1170)\to K^{*+}K^{-})$       $0.85\pm 0.76$       $\mathcal{B}(h_{1}(1415)\to K^{*+}K^{-})$       $12.15\pm 9.79$
      $\mathcal{B}(h_{1}(1170)\to K^{*-}K^{+})$       $0.85\pm 0.76$       $\mathcal{B}(h_{1}(1415)\to K^{*-}K^{+})$       $12.15\pm 9.79$
      $\mathcal{B}(h_{1}(1170)\to K^{*0}\overline{K}^{0})$       $0.81\pm 0.73$       $\mathcal{B}(h_{1}(1415)\to K^{*0}\overline{K}^{0})$       $10.42\pm 8.50$
      $\mathcal{B}(h_{1}(1170)\to\overline{K}^{*0}K^{0})$       $0.81\pm 0.73$       $\mathcal{B}(h_{1}(1415)\to\overline{K}^{*0}K^{0})$       $10.42\pm 8.50$
      $\mathcal{B}(h_{1}(1170)\to\phi\eta)$       $0.08\pm 0.08$       $\mathcal{B}(h_{1}(1415)\to\phi\eta)$       $1.19\pm 0.84$
      $\mathcal{B}(h_{1}(1170)\to\rho\pi)$       $65.31\pm 27.91$       $\mathcal{B}(h_{1}(1415)\to\rho\pi)$       $39.54\pm 39.53$
      $\mathcal{B}(h_{1}(1170)\to K^{*}K)$       $3.33\pm 2.98$       $\mathcal{B}(h_{1}(1415)\to K^{*}K)$       $45.13\pm 36.58$
Most of our branching ratio predictions have large errors, since we consider the errors of the decay widths and the masses; we conservatively choose the parameters $D$ and $F$ as well as three mixing angles, and consider the present large experimental errors of $K_{1}(1270)/K_{1}(1400)\to VP$. If the result is shown as $a\pm a$, it means that the allowed lower limit is very small, and only the upper limit $2a$ is obtained.
Fitted results were previously obtained by three possible solutions in Ref. Roca:2003uk . Apart from the measured decays of $K_{1}(1270)$ and $K_{1}(1400)$ mesons given in Eqs. (21-22), the average values of the three fitted central values are Roca:2003uk
$\mathcal{B}(a_{1}(1260)\to\rho\pi)=62.75\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(a_{1}(1260)\to K^{*}K)=7.53\%,$ (25)
$\mathcal{B}(b_{1}(1235)\to\omega\pi)=73.47\%,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}(b_{1}(1235)\to K^{*}K)=5.92\%.$ (26)
Our prediction of $\mathcal{B}(b_{1}(1235)\to K^{*}K)$ is smaller than that in Ref. Roca:2003uk ; nevertheless, the other three branching ratios are consistent with each other.
The sensitivity of $\mathcal{B}(A\to VP)$ to the constrained parameters is shown in Figs. 4-6. As shown in Figs. 4-6, the upper limits of $K_{1}(1270)\to K^{*}\pi,\rho K$ and $K_{1}(1400)\to\rho K$ and the lower limits of $K_{1}(1270)\to K^{*}\pi,\omega K$ and $K_{1}(1400)\to K^{*}\pi,\rho K,\omega K$ give the effective constraints on the relevant non-perturbative parameters. $\mathcal{B}(K_{1}(1270)\to K^{*}\pi)$ and $\mathcal{B}(K_{1}(1270)\to K^{*}\eta)$ are sensitive to $F$ and $\theta_{K1}$. $\mathcal{B}(K_{1}(1400)\to\rho K)$, $\mathcal{B}(K_{1}(1400)\to\omega K)$ and $\mathcal{B}(K_{1}(1400)\to K^{*}\eta)$ are sensitive to $D$ and $\theta_{K1}$. $\mathcal{B}(K_{1}(1400)\to\rho K)$ and $\mathcal{B}(K_{1}(1400)\to\omega K)$ have similar change trends, and the inflection point values appear when $D\approx-1.0$ and $\theta_{K_{1}}\approx 56^{\circ}$. As displayed in Fig. 6, many $\mathcal{B}(a_{1}/b_{1}/f_{1}/h_{1}\to VP)$ are very sensitive to $F$, $D$, $\theta_{3_{P1}}$ and $\theta_{1_{P1}}$. Any future measurement of these decays can obviously place constraints on the relevant non-perturbative parameters.
Figure 4: The constrained effects of the non-perturbative parameters in the $K_{1}(1270)\to VP$ decays (in units of $10^{-2}$).
Figure 5: The constrained effects of the non-perturbative parameters in the $K_{1}(1400)\to VP$ decays (in units of $10^{-2}$).
Figure 6: The constrained effects of the non-perturbative parameters in the $a_{1}/b_{1}/f_{1}/h_{1}\to VP$ decays ( in units of $10^{-2}$).
II.3 Results of the $D\to A(A\to VP)\ell^{+}\nu$ decays
For the four-body semileptonic decays with resonances, in the case where the decay widths of the resonances are very narrow, the resonant decay branching ratios respect a simple factorization relation
$\mathcal{B}(D\to R\ell^{+}\nu_{\ell},R\to VP)=\mathcal{B}(D\to R\ell^{+}\nu_{\ell})\times\mathcal{B}(R\to VP),$ (27)
where $R$ denotes the resonance. This relation is also a good approximation for wider resonances.
At present, only experimental upper limits of $\mathcal{B}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\omega\pi^{0})$ and $\mathcal{B}(D^{0}\to b_{1}(1235)^{-}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{-}\to\omega\pi^{-})$ have been reported PDG2022 ; BESIII:2020jan
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\omega\pi^{0})\leq 1.75\times 10^{-4},$ (28)
$\mathcal{B}(D^{0}\to b_{1}(1235)^{-}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{-}\to\omega\pi^{-})\leq 1.12\times 10^{-4}.$ (29)
Branching ratios of $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays can be obtained by using $\mathcal{B}(D\rightarrow A\ell^{+}\nu_{\ell})$ given in Sec. II.1 and $\mathcal{B}(A\rightarrow VP)$ given in Sec. II.2 as well as considering the further constraints in Eqs. (28-29), and our predictions are listed in the second and fourth columns of Tabs. 7-9.
[…]
Decay modes $\mathcal{B}_{[A]}(\ell=e)$ $\mathcal{B}_{[T]}(\ell=e)$ $\mathcal{B}_{[A]}(\ell=\mu)$ $\mathcal{B}_{[T]}(\ell=\mu)$
$D^{0}\to K^{*-}\pi^{0}\ell^{+}\nu_{\ell}$ $7.29\pm 3.48_{[K_{1}(1270)^{-}]}$ $0.11\pm 0.05_{[K^{*}_{2}(1430)^{-}]}$ $6.29\pm 3.00_{[K_{1}(1270)^{-}]}$ $0.08\pm 0.03_{[K^{*}_{2}(1430)^{-}]}$
$0.60\pm 0.60_{[K_{1}(1400)^{-}]}$ $0.47\pm 0.47_{[K_{1}(1400)^{-}]}$
$D^{0}\to\overline{K}^{*0}\pi^{-}\ell^{+}\nu_{\ell}$ $14.30\pm 6.83_{[K_{1}(1270)^{-}]}$ $0.22\pm 0.09_{[K^{*}_{2}(1430)^{-}]}$ $12.33\pm 5.89_{[K_{1}(1270)^{-}]}$ $0.15\pm 0.05_{[K^{*}_{2}(1430)^{-}]}$
$1.19\pm 1.19_{[K_{1}(1400)^{-}]}$ $0.93\pm 0.93_{[K_{1}(1400)^{-}]}$
$D^{0}\to\rho^{0}K^{-}\ell^{+}\nu_{\ell}$ $15.77\pm 1.97_{[K_{1}(1270)^{-}]}$ $0.04\pm 0.02_{[K^{*}_{2}(1430)^{-}]}$ $13.61\pm 1.70_{[K_{1}(1270)^{-}]}$ $0.03\pm 0.01_{[K^{*}_{2}(1430)^{-}]}$
$0.03\pm 0.03_{[K_{1}(1400)^{-}]}$ $0.02\pm 0.02_{[K_{1}(1400)^{-}]}$
$D^{0}\to\rho^{-}\overline{K}^{0}\ell^{+}\nu_{\ell}$ $30.64\pm 3.88_{[K_{1}(1270)^{-}]}$ $0.08\pm 0.03_{[K^{*}_{2}(1430)^{-}]}$ $26.44\pm 3.36_{[K_{1}(1270)^{-}]}$ $0.05\pm 0.02_{[K^{*}_{2}(1430)^{-}]}$
$0.06\pm 0.06_{[K_{1}(1400)^{-}]}$ $0.05\pm 0.05_{[K_{1}(1400)^{-}]}$
$D^{0}\to\omega K^{-}\ell^{+}\nu_{\ell}$ $8.93\pm 1.12_{[K_{1}(1270)^{-}]}$ $0.04\pm 0.02_{[K^{*}_{2}(1430)^{-}]}$ $7.70\pm 0.96_{[K_{1}(1270)^{-}]}$ $0.03\pm 0.02_{[K^{*}_{2}(1430)^{-}]}$
$0.03\pm 0.03_{[K_{1}(1400)^{-}]}$ $0.02\pm 0.02_{[K_{1}(1400)^{-}]}$
$D^{0}\to K^{*-}\eta\ell^{+}\nu_{\ell}$ $0.48\pm 0.20_{[K_{1}(1270)^{-}]}$ $0.008\pm 0.003_{[K^{*}_{2}(1430)^{-}]}$ $0.41\pm 0.17_{[K_{1}(1270)^{-}]}$ $0.005\pm 0.002_{[K^{*}_{2}(1430)^{-}]}$
$0.04\pm 0.04_{[K_{1}(1400)^{-}]}$ $0.03\pm 0.03_{[K_{1}(1400)^{-}]}$
$D^{0}\to\phi K^{-}\ell^{+}\nu_{\ell}$ $0.07\pm 0.07_{[K_{1}(1400)^{-}]}$ $0.0005\pm 0.0002_{[K^{*}_{2}(1430)^{-}]}$ $0.06\pm 0.06_{[K_{1}(1400)^{-}]}$ $0.0004\pm 0.0002_{[K^{*}_{2}(1430)^{-}]}$
$D^{0}\to\rho^{-}\pi^{0}\ell^{+}\nu_{\ell}$ $1.87\pm 1.09_{[a_{1}(1260)^{-}]}$ $0.12\pm 0.05_{[a_{2}(1320)^{-}]}$ $1.64\pm 0.97_{[a_{1}(1260)^{-}]}$ $0.09\pm 0.03_{[a_{2}(1320)^{-}]}$
$D^{0}\to\rho^{0}\pi^{-}\ell^{+}\nu_{\ell}$ $1.87\pm 1.08_{[a_{1}(1260)^{-}]}$ $0.12\pm 0.05_{[a_{2}(1320)^{-}]}$ $1.64\pm 0.96_{[a_{1}(1260)^{-}]}$ $0.09\pm 0.03_{[a_{2}(1320)^{-}]}$
$D^{0}\to K^{*0}K^{-}\ell^{+}\nu_{\ell}$ $0.15\pm 0.07_{[a_{1}(1260)^{-}]}$ $0.0003\pm 0.0001_{[a_{2}(1320)^{-}]}$ $0.13\pm 0.06{[a_{1}(1260)^{-}]}$ $0.0003\pm 0.0001_{[a_{2}(1320)^{-}]}$
$0.02\pm 0.02_{[b_{1}(1235)^{-}]}$ $0.02\pm 0.01_{[b_{1}(1235)^{-}]}$
$D^{0}\to K^{*-}K^{0}\ell^{+}\nu_{\ell}$ $0.15\pm 0.07_{[a_{1}(1260)^{-}]}$ $0.0004\pm 0.0002_{[a_{2}(1320)^{-}]}$ $0.13\pm 0.06_{[a_{1}(1260)^{-}]}$ $0.0003\pm 0.0001_{[a_{2}(1320)^{-}]}$
$0.02\pm 0.02_{[b_{1}(1235)^{-}]}$ $0.02\pm 0.02_{[b_{1}(1235)^{-}]}$
$D^{0}\to\omega\pi^{-}\ell^{+}\nu_{\ell}$ $1.89\pm 1.26_{[b_{1}(1235)^{-}]}$ $1.65\pm 1.11_{[b_{1}(1235)^{-}]}$
$D^{0}\to\rho^{-}\eta\ell^{+}\nu_{\ell}$ $0.16\pm 0.10_{[b_{1}(1235)^{-}]}$ $0.14\pm 0.09_{[b_{1}(1235)^{-}]}$
[…]
Decay modes $\mathcal{B}_{[A]}(\ell=e)$ $\mathcal{B}_{[T]}(\ell=e)$ $\mathcal{B}_{[A]}(\ell=\mu)$ $\mathcal{B}_{[T]}(\ell=\mu)$
$D^{+}\to K^{*-}\pi^{+}\ell^{+}\nu_{\ell}$ $37.70\pm 17.93_{[\overline{K}_{1}(1270)^{0}]}$ $0.59\pm 0.23_{[\overline{K}^{*}_{2}(1430)^{0}]}$ $32.59\pm 15.51_{[\overline{K}_{1}(1270)^{0}]}$ $0.39\pm 0.14_{[\overline{K}^{*}_{2}(1430)^{0}]}$
$3.17\pm 3.17_{[\overline{K}_{1}(1400)^{0}]}$ $2.49\pm 2.49_{[\overline{K}_{1}(1400)^{0}]}$
$D^{+}\to\overline{K}^{*0}\pi^{0}\ell^{+}\nu_{\ell}$ $18.73\pm 8.91_{[\overline{K}_{1}(1270)^{0}]}$ $0.29\pm 0.11_{[\overline{K}^{*}_{2}(1430)^{0}]}$ $16.19\pm 7.71_{[\overline{K}_{1}(1270)^{0}]}$ $0.19\pm 0.07_{[\overline{K}^{*}_{2}(1430)^{0}]}$
$1.58\pm 1.58_{[\overline{K}_{1}(1400)^{0}]}$ $1.24\pm 1.24_{[\overline{K}_{1}(1400)^{0}]}$
$D^{+}\to\rho^{+}K^{-}\ell^{+}\nu_{\ell}$ $82.47\pm 10.58_{[\overline{K}_{1}(1270)^{0}]}$ $0.21\pm 0.09_{[\overline{K}^{*}_{2}(1430)^{0}]}$ $71.30\pm 9.16_{[\overline{K}_{1}(1270)^{0}]}$ $0.14\pm 0.06_{[\overline{K}^{*}_{2}(1430)^{0}]}$
$0.16\pm 0.16_{[\overline{K}_{1}(1400)^{0}]}$ $0.13\pm 0.13_{[\overline{K}_{1}(1400)^{0}]}$
$D^{+}\to\rho^{0}\overline{K}^{0}\ell^{+}\nu_{\ell}$ $39.54\pm 5.02_{[\overline{K}_{1}(1270)^{0}]}$ $0.10\pm 0.04_{[\overline{K}^{*}_{2}(1430)^{0}]}$ $34.20\pm 4.36_{[\overline{K}_{1}(1270)^{0}]}$ $0.07\pm 0.03_{[\overline{K}^{*}_{2}(1430)^{0}]}$
$0.08\pm 0.08_{[\overline{K}_{1}(1400)^{0}]}$ $0.06\pm 0.06_{[\overline{K}_{1}(1400)^{0}]}$
$D^{+}\to\omega\overline{K}^{0}\ell^{+}\nu_{\ell}$ $21.65\pm 2.78_{[\overline{K}_{1}(1270)^{0}]}$ $0.11\pm 0.06_{[\overline{K}^{*}_{2}(1430)^{0}]}$ $18.72\pm 2.41_{[\overline{K}_{1}(1270)^{0}]}$ $0.07\pm 0.04_{[\overline{K}^{*}_{2}(1430)^{0}]}$
$0.08\pm 0.08_{[\overline{K}_{1}(1400)^{0}]}$ $0.06\pm 0.06_{[\overline{K}_{1}(1400)^{0}]}$
$D^{+}\to\overline{K}^{*0}\eta\ell^{+}\nu_{\ell}$ $1.03\pm 0.44_{[\overline{K}_{1}(1270)^{0}]}$ $0.02\pm 0.01_{[\overline{K}^{*}_{2}(1430)^{0}]}$ $0.89\pm 0.38_{[\overline{K}_{1}(1270)^{0}]}$ $0.01\pm 0.005_{[\overline{K}^{*}_{2}(1430)^{0}]}$
$0.09\pm 0.09_{[\overline{K}_{1}(1400)^{0}]}$ $0.07\pm 0.07_{[\overline{K}_{1}(1400)^{0}]}$
$D^{+}\to\phi\overline{K}^{0}\ell^{+}\nu_{\ell}$ $0.18\pm 0.18_{[\overline{K}_{1}(1400)^{0}]}$ $0.001\pm 0.0006_{[\overline{K}^{*}_{2}(1430)^{0}]}$ $0.14\pm 0.14_{[\overline{K}_{1}(1400)^{0}]}$ $0.0009\pm 0.0004_{[\overline{K}^{*}_{2}(1430)^{0}]}$
$D^{+}\to\rho^{+}\pi^{-}\ell^{+}\nu_{\ell}$ $2.41\pm 1.39_{[a_{1}(1260)^{0}]}$ $0.16\pm 0.06_{[a_{2}(1320)^{0}]}$ $2.12\pm 1.24_{[a_{1}(1260)^{0}]}$ $0.13\pm 0.05_{[a_{2}(1320)^{0}]}$
$1.40\pm 1.05_{[h_{1}(1170)^{0}]}$ $1.25\pm 0.94_{[[h_{1}(1170)^{0}]}$
$0.03\pm 0.03_{[h_{1}(1415)^{0}]}$ $0.02\pm 0.02_{[[h_{1}(1415)^{0}]}$
$D^{+}\to\rho^{-}\pi^{+}\ell^{+}\nu_{\ell}$ $2.41\pm 1.39_{[a_{1}(1260)^{0}]}$ $0.16\pm 0.06_{[a_{2}(1320)^{0}]}$ $2.12\pm 1.24_{[a_{1}(1260)^{0}]}$ $0.13\pm 0.05_{[a_{2}(1320)^{0}]}$
$1.40\pm 1.05_{[h_{1}(1170)^{0}]}$ $1.25\pm 0.94_{[[h_{1}(1170)^{0}]}$
$0.03\pm 0.03_{[h_{1}(1415)^{0}]}$ $0.02\pm 0.02_{[[h_{1}(1415)^{0}]}$
$D^{+}\to\omega\pi^{0}\ell^{+}\nu_{\ell}$ $2.47\pm 1.66_{[b_{1}(1235)^{0}]}$ $2.17\pm 1.45_{[b_{1}(1235)^{0}]}$
$D^{+}\to\rho^{0}\eta\ell^{+}\nu_{\ell}$ $0.20\pm 0.13_{[b_{1}(1235)^{0}]}$ $0.17\pm 0.12_{[b_{1}(1235)^{0}]}$
$D^{+}\to\rho^{0}\pi^{0}\ell^{+}\nu_{\ell}$ $1.40\pm 1.05_{[h_{1}(1170)^{0}]}$ $1.26\pm 0.95_{[h_{1}(1170)^{0}]}$
$0.03\pm 0.03_{[h_{1}(1415)^{0}]}$ $0.02\pm 0.02_{[h_{1}(1415)^{0}]}$
$D^{+}\to\omega\eta\ell^{+}\nu_{\ell}$ $0.12\pm 0.10_{[h_{1}(1170)^{0}]}$ $0.11\pm 0.09_{[h_{1}(1170)^{0}]}$
$0.007\pm 0.007_{[h_{1}(1415)^{0}]}$ $0.006\pm 0.006_{[h_{1}(1415)^{0}]}$
$D^{+}\to\phi\eta\ell^{+}\nu_{\ell}$ $0.004\pm 0.004_{[h_{1}(1170)^{0}]}$ $0.004\pm 0.004_{[h_{1}(1170)^{0}]}$
$0.001\pm 0.001_{[h_{1}(1415)^{0}]}$ $0.0008\pm 0.0008_{[h_{1}(1415)^{0}]}$
$D^{+}\to K^{*+}K^{-}\ell^{+}\nu_{\ell}$ $0.10\pm 0.05_{[a_{1}(1260)^{0}]}$ $0.0003\pm 0.0001_{[a_{2}(1320)^{0}]}$ $0.09\pm 0.04_{[a_{1}(1260)^{0}]}$ $0.0002\pm 0.00008_{[a_{2}(1320)^{0}]}$
$0.02\pm 0.01_{[b_{1}(1235)^{0}]}$ $0.0006\pm 0.0003_{[f_{2}(1270)]}$ $0.01\pm 0.01_{[b_{1}(1235)^{0}]}$ $0.0005\pm 0.0002_{[f_{2}(1270)]}$
$0.01\pm 0.01_{[f_{1}(1285)]}$ $0.00003\pm 0.00002_{[f_{2}^{\prime}(1525)]}$ $0.009\pm 0.009_{[f_{1}(1285)]}$ $0.00002\pm 0.00001_{[f_{2}^{\prime}(1525)]}$
$0.15\pm 0.15_{[f_{1}(1420)]}$ $0.11\pm 0.11_{[f_{1}(1420)]}$
$0.05\pm 0.05_{[h_{1}(1170)]}$ $0.05\pm 0.04_{[h_{1}(1170)]}$
$0.007\pm 0.007_{[h_{1}(1415)]}$ $0.006\pm 0.006_{[h_{1}(1415)]}$
$D^{+}\to K^{*-}K^{+}\ell^{+}\nu_{\ell}$ $0.10\pm 0.05_{[a_{1}(1260)^{0}]}$ $0.0003\pm 0.0001_{[a_{2}(1320)^{0}]}$ $0.09\pm 0.04_{[a_{1}(1260)^{0}]}$ $0.0002\pm 0.00008_{[a_{2}(1320)^{0}]}$
$0.02\pm 0.01_{[b_{1}(1235)^{0}]}$ $0.0006\pm 0.0003_{[f_{2}(1270)]}$ $0.01\pm 0.01_{[b_{1}(1235)^{0}]}$ $0.0005\pm 0.0002_{[f_{2}(1270)]}$
$0.01\pm 0.01_{[f_{1}(1285)]}$ $0.00003\pm 0.00002_{[f_{2}^{\prime}(1525)]}$ $0.009\pm 0.009_{[f_{1}(1285)]}$ $0.00002\pm 0.00001_{[f_{2}^{\prime}(1525)]}$
$0.15\pm 0.15_{[f_{1}(1420)]}$ $0.11\pm 0.11_{[f_{1}(1420)]}$
$0.05\pm 0.05_{[h_{1}(1170)]}$ $0.05\pm 0.04_{[h_{1}(1170)]}$
$0.007\pm 0.007_{[h_{1}(1415)]}$ $0.006\pm 0.006_{[h_{1}(1415)]}$
$D^{+}\to K^{*0}\overline{K}^{0}\ell^{+}\nu_{\ell}$ $0.09\pm 0.04_{[a_{1}(1260)^{0}]}$ $0.0002\pm 0.00009_{[a_{2}(1320)^{0}]}$ $0.08\pm 0.04_{[a_{1}(1260)^{0}]}$ $0.0002\pm 0.00006_{[a_{2}(1320)^{0}]}$
$0.01\pm 0.01_{[b_{1}(1235)^{0}]}$ $0.0005\pm 0.0002_{[f_{2}(1270)]}$ $0.01\pm 0.009_{[b_{1}(1235)^{0}]}$ $0.0004\pm 0.0002_{[f_{2}(1270)]}$
$0.002\pm 0.002_{[f_{1}(1285)]}$ $0.00003\pm 0.00002_{[f_{2}^{\prime}(1525)]}$ $0.002\pm 0.002_{[f_{1}(1285)]}$ $0.00002\pm 0.00001_{[f_{2}^{\prime}(1525)]}$
$0.13\pm 0.13_{[f_{1}(1420)]}$ $0.10\pm 0.10_{[f_{1}(1420)]}$
$0.05\pm 0.05_{[h_{1}(1170)]}$ $0.04\pm 0.04_{[h_{1}(1170)]}$
$0.006\pm 0.006_{[h_{1}(1415)]}$ $0.005\pm 0.005_{[h_{1}(1415)]}$
$D^{+}\to\overline{K}^{*0}K^{0}\ell^{+}\nu_{\ell}$ $0.09\pm 0.04_{[a_{1}(1260)^{0}]}$ $0.0002\pm 0.00009_{[a_{2}(1320)^{0}]}$ $0.08\pm 0.04_{[a_{1}(1260)^{0}]}$ $0.0002\pm 0.00006_{[a_{2}(1320)^{0}]}$
$0.01\pm 0.01_{[b_{1}(1235)^{0}]}$ $0.0005\pm 0.0002_{[f_{2}(1270)]}$ $0.01\pm 0.009_{[b_{1}(1235)^{0}]}$ $0.0004\pm 0.0002_{[f_{2}(1270)]}$
$0.002\pm 0.002_{[f_{1}(1285)]}$ $0.00003\pm 0.00002_{[f_{2}^{\prime}(1525)]}$ $0.002\pm 0.002_{[f_{1}(1285)]}$ $0.00002\pm 0.00001_{[f_{2}^{\prime}(1525)]}$
$0.13\pm 0.13_{[f_{1}(1420)]}$ $0.10\pm 0.10_{[f_{1}(1420)]}$
$0.05\pm 0.05_{[h_{1}(1170)]}$ $0.04\pm 0.04_{[h_{1}(1170)]}$
$0.006\pm 0.006_{[h_{1}(1415)]}$ $0.005\pm 0.005_{[h_{1}(1415)]}$
[…]
Decay modes $\mathcal{B}_{[A]}(\ell=e)$ $\mathcal{B}_{[T]}(\ell=e)$ $\mathcal{B}_{[A]}(\ell=\mu)$ $\mathcal{B}_{[T]}(\ell=\mu)$
$D_{s}^{+}\to K^{*-}\pi^{+}\ell^{+}\nu_{\ell}$ $1.82\pm 0.87_{[K_{1}(1270)^{0}]}$ $0.06\pm 0.02_{[K^{*}_{2}(1430)^{0}]}$ $1.64\pm 0.78_{[K_{1}(1270)^{0}]}$ $0.04\pm 0.02_{[K^{*}_{2}(1430)^{0}]}$
$0.19\pm 0.19_{[K_{1}(1400)^{0}]}$ $0.16\pm 0.16_{[K_{1}(1400)^{0}]}$
$D_{s}^{+}\to\overline{K}^{*0}\pi^{0}\ell^{+}\nu_{\ell}$ $0.91\pm 0.43_{[K_{1}(1270)^{0}]}$ $0.03\pm 0.01_{[K^{*}_{2}(1430)^{0}]}$ $0.81\pm 0.39_{[K_{1}(1270)^{0}]}$ $0.02\pm 0.008_{[K^{*}_{2}(1430)^{0}]}$
$0.09\pm 0.09_{[K_{1}(1400)^{0}]}$ $0.08\pm 0.08_{[K_{1}(1400)^{0}]}$
$D_{s}^{+}\to\rho^{+}K^{-}\ell^{+}\nu_{\ell}$ $4.00\pm 0.54_{[K_{1}(1270)^{0}]}$ $0.02\pm 0.009_{[K^{*}_{2}(1430)^{0}]}$ $3.59\pm 0.48_{[K_{1}(1270)^{0}]}$ $0.02\pm 0.007_{[K^{*}_{2}(1430)^{0}]}$
$0.01\pm 0.01_{[K_{1}(1400)^{0}]}$ $0.008\pm 0.008_{[K_{1}(1400)^{0}]}$
$D_{s}^{+}\to\rho^{0}\overline{K}^{0}\ell^{+}\nu_{\ell}$ $1.92\pm 0.25_{[K_{1}(1270)^{0}]}$ $0.01\pm 0.004_{[K^{*}_{2}(1430)^{0}]}$ $1.72\pm 0.23_{[K_{1}(1270)^{0}]}$ $0.008\pm 0.003_{[K^{*}_{2}(1430)^{0}]}$
$0.005\pm 0.005_{[K_{1}(1400)^{0}]}$ $0.004\pm 0.004_{[K_{1}(1400)^{0}]}$
$D_{s}^{+}\to\omega\overline{K}^{0}\ell^{+}\nu_{\ell}$ $1.04\pm 0.14_{[K_{1}(1270)^{0}]}$ $0.01\pm 0.006_{[K^{*}_{2}(1430)^{0}]}$ $0.94\pm 0.12_{[K_{1}(1270)^{0}]}$ $0.008\pm 0.004_{[K^{*}_{2}(1430)^{0}]}$
$0.005\pm 0.005_{[K_{1}(1400)^{0}]}$ $0.004\pm 0.004_{[K_{1}(1400)^{0}]}$
$D_{s}^{+}\to\overline{K}^{*0}\eta\ell^{+}\nu_{\ell}$ $0.05\pm 0.02_{[K_{1}(1270)^{0}]}$ $0.002\pm 0.0009_{[K^{*}_{2}(1430)^{0}]}$ $0.05\pm 0.02_{[K_{1}(1270)^{0}]}$ $0.001\pm 0.0006_{[K^{*}_{2}(1430)^{0}]}$
$0.005\pm 0.005_{[K_{1}(1400)^{0}]}$ $0.004\pm 0.004_{[K_{1}(1400)^{0}]}$
$D_{s}^{+}\to\phi\overline{K}^{0}\ell^{+}\nu_{\ell}$ $0.01\pm 0.01_{[K_{1}(1400)^{0}]}$ $0.0001\pm 0.00006_{[K^{*}_{2}(1430)^{0}]}$ $0.009\pm 0.009_{[K_{1}(1400)^{0}]}$ $0.0001\pm 0.00005_{[K^{*}_{2}(1430)^{0}]}$
$D_{s}^{+}\to\rho^{+}\pi^{-}\ell^{+}\nu_{\ell}$ $4.21\pm 4.21_{[h_{1}(1170)]}$ $3.86\pm 3.86_{[h_{1}(1170)]}$
$3.47\pm 3.47_{[h_{1}(1415)]}$ $2.92\pm 2.92_{[[h_{1}(1415)]}$
$D_{s}^{+}\to\rho^{-}\pi^{+}\ell^{+}\nu_{\ell}$ $4.21\pm 4.21_{[h_{1}(1170)]}$ $3.86\pm 3.86_{[h_{1}(1170)]}$
$3.47\pm 3.47_{[h_{1}(1415)]}$ $2.92\pm 2.92_{[[h_{1}(1415)]}$
$D_{s}^{+}\to\rho^{0}\pi^{0}\ell^{+}\nu_{\ell}$ $4.23\pm 4.23_{[h_{1}(1170)]}$ $3.89\pm 3.88_{[h_{1}(1170)]}$
$3.48\pm 3.48_{[h_{1}(1415)]}$ $2.93\pm 2.93_{[[h_{1}(1415)]}$
$D_{s}^{+}\to\omega\eta\ell^{+}\nu_{\ell}$ $0.41\pm 0.41_{[h_{1}(1170)]}$ $0.37\pm 0.37_{[h_{1}(1170)]}$
$0.89\pm 0.89_{[h_{1}(1415)]}$ $0.75\pm 0.75_{[[h_{1}(1415)]}$
$D_{s}^{+}\to K^{*+}K^{-}\ell^{+}\nu_{\ell}$ $0.51\pm 0.46_{[f_{1}(1285)]}$ $0.0007\pm 0.0004_{[f_{2}(1270)]}$ $0.46\pm 0.41_{[f_{1}(1285)]}$ $0.0006\pm 0.0003_{[f_{2}(1270)]}$
$5.24\pm 5.24_{[f_{1}(1420)]}$ $0.11\pm 0.04_{[f_{2}^{\prime}(1525)]}$ $4.36\pm 4.36_{[f_{1}(1420)]}$ $0.07\pm 0.03_{[f_{2}^{\prime}(1525)]}$
$0.37\pm 0.37_{[h_{1}(1170)]}$ $0.34\pm 0.34_{[h_{1}(1170)]}$
$4.04\pm 3.70_{[h_{1}(1415)]}$ $3.39\pm 3.11_{[h_{1}(1415)]}$
$D_{s}^{+}\to K^{*-}K^{+}\ell^{+}\nu_{\ell}$ $0.51\pm 0.46_{[f_{1}(1285)]}$ $0.0007\pm 0.0004_{[f_{2}(1270)]}$ $0.46\pm 0.41_{[f_{1}(1285)]}$ $0.0006\pm 0.0003_{[f_{2}(1270)]}$
$5.24\pm 5.24_{[f_{1}(1420)]}$ $0.11\pm 0.04_{[f_{2}^{\prime}(1525)]}$ $4.36\pm 4.36_{[f_{1}(1420)]}$ $0.07\pm 0.03_{[f_{2}^{\prime}(1525)]}$
$0.37\pm 0.37_{[h_{1}(1170)]}$ $0.34\pm 0.34_{[h_{1}(1170)]}$
$4.04\pm 3.70_{[h_{1}(1415)]}$ $3.39\pm 3.11_{[h_{1}(1415)]}$
$D_{s}^{+}\to K^{*0}\overline{K}^{0}\ell^{+}\nu_{\ell}$ $0.11\pm 0.10_{[f_{1}(1285)]}$ $0.0007\pm 0.0003_{[f_{2}(1270)]}$ $0.10\pm 0.09_{[f_{1}(1285)]}$ $0.0005\pm 0.0003_{[f_{2}(1270)]}$
$4.71\pm 4.71_{[f_{1}(1420)]}$ $0.10\pm 0.04_{[f_{2}^{\prime}(1525)]}$ $3.92\pm 3.92_{[f_{1}(1420)]}$ $0.07\pm 0.02_{[f_{2}^{\prime}(1525)]}$
$0.35\pm 0.35_{[h_{1}(1170)]}$ $0.32\pm 0.32_{[h_{1}(1170)]}$
$3.43\pm 3.16_{[h_{1}(1415)]}$ $2.88\pm 2.65_{[h_{1}(1415)]}$
[…]
$4.71\pm 4.71_{[f_{1}(1420)]}$ $0.10\pm 0.04_{[f_{2}^{\prime}(1525)]}$ $3.92\pm 3.92_{[f_{1}(1420)]}$ $0.07\pm 0.02_{[f_{2}^{\prime}(1525)]}$
$0.35\pm 0.35_{[h_{1}(1170)]}$ $0.32\pm 0.32_{[h_{1}(1170)]}$
$3.43\pm 3.16_{[h_{1}(1415)]}$ $2.88\pm 2.65_{[h_{1}(1415)]}$
Comparing the experimental upper limits in Eqs. (28-29) and our predictions in Tab. 8, we can see our predictions of $\mathcal{B}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\omega\pi^{0})$ and $\mathcal{B}(D^{0}\to b_{1}(1235)^{-}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{-}\to\omega\pi^{-})$ are much smaller than the experimental upper limits. After we completed the analysis of the axial-vector resonant contributions, we saw the experimental report regarding $\mathcal{B}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\omega\pi^{0})$ and $\mathcal{B}(D^{0}\to b_{1}(1235)^{-}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{-}\to\omega\pi^{-})$ from BESIII BESIII:2024pwp
$\mathcal{B}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\omega\pi^{0})$ $=$ $(1.16\pm 0.44\pm 0.16)\times 10^{-4},$ (30)
$\mathcal{B}(D^{0}\to b_{1}(1235)^{-}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{-}\to\omega\pi^{-})$ $=$ $(0.72\pm 0.18^{+0.06}_{-0.08})\times 10^{-4}.$ (31)
The above data are slightly larger than our predictions in Tab. 8, and they are consistent within $2\sigma$ errors. As mentioned earlier, for some branching ratio predictions, due to poor experimental data at present and our conservative choice of the non-perturbative parameters, their errors are as large as their central values, which means that we can only give the upper limits. Many predictions of $\mathcal{B}(D\to A\ell^{+}\nu_{\ell},A\to VP)$ are on the order of $\mathcal{O}(10^{-5}-10^{-4})$, and they are expected to be measured in the near future.
III Semileptonic $D\rightarrow T(T\to PV)\ell^{+}\nu_{\ell}$ decays
III.1 Semileptonic $D\rightarrow T\ell^{+}\nu_{\ell}$ decays
III.1.1 Theoretical framework for the $D\to T\ell^{+}\nu_{\ell}$ decays
With the same effective Hamiltonian given in Eq. (1), the decay amplitudes of the $D\to T\ell^{+}\nu_{\ell}$ decays are similar to those of the $D\to A\ell^{+}\nu_{\ell}$ decays given in Eqs. (2-4) by replacing $A$ with $T$.
The form factors of the $D\to T$ transitions are similar to those of the $B\to T$ transitions Chen:2021ywv
$T(p,\varepsilon^{*})\left|\bar{u}_{k}\gamma_{\mu}(1-\gamma_{5})b\right|D(p_{D})\right>$ $=$ $\frac{2iV^{T}(q^{2})}{m_{D}+m_{T}}\epsilon_{\mu\nu\alpha\beta}e^{*\nu}p^{\alpha}_{D}p^{\beta}$ (32)
$+2m_{T}\frac{e^{*}\cdot q}{q^{2}}q_{\mu}A^{T}_{0}(q^{2})+(m_{D}+m_{T})\Big{(}e^{*}_{\mu}-\frac{e^{*}\cdot q}{q^{2}}q_{\mu}\Big{)}A^{T}_{1}(q^{2})$
$-\frac{e^{*}\cdot q}{m_{D}+m_{T}}\Big{(}(p_{D}+p)_{\mu}-\frac{m_{D}^{2}-m^{2}_{T}}{q^{2}}q_{\mu}\Big{)}A^{T}_{2}(q^{2}),$
where $q^{2}\equiv(p_{D}-p)^{2}$, $V^{T}(q^{2})$, and $A^{T}_{0,1,2}(q^{2})$ are the form factors of the $D\to T$ transition, $m_{T}$ is the mass of the tensor mesons, and $e^{*\nu}\equiv\frac{\varepsilon^{*\mu\nu}\cdot p_{D\mu}}{m_{D}}$. The hadronic amplitudes of $D\to T\ell^{+}\nu_{\ell}$ decays can then be written as
$H^{T}_{\pm}$ $=$ $\frac{2|\vec{p}_{T}|}{\sqrt{6}m_{T}}\left[(m_{D_{q}}+m_{T})A^{T}_{1}(q^{2})\mp\frac{2m_{D_{q}}|\vec{p}_{T}|}{(m_{D_{q}}+m_{T})}V^{T}(q^{2})\right],$ (33)
$H^{T}_{0}$ $=$ $\frac{|\vec{p}_{T}|}{\sqrt{2}m_{T}}\frac{1}{2m_{T}\sqrt{q^{2}}}\left[(m_{D_{q}}^{2}-m_{T}^{2}-q^{2})(m_{D_{q}}+m_{T})A^{T}_{1}(q^{2})-\frac{4m_{D_{q}}^{2}|\vec{p}_{T}|^{2}}{m_{D_{q}}+m_{T}}A^{T}_{2}(q^{2})\right],$ (34)
$H^{T}_{t}$ $=$ $\frac{|\vec{p}_{T}|}{\sqrt{2}m_{T}}\frac{2m_{D_{q}}|\vec{p}_{T}|}{\sqrt{q^{2}}}A^{T}_{0}(q^{2}),$ (35)
where $|\vec{p}_{T}|\equiv\sqrt{\lambda(m_{D_{q}}^{2},m_{T}^{2},q^{2})}/2m_{D_{q}}$ with $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2ab-2ac-2bc$.
Similar to the $D\to A\ell^{+}\nu_{\ell}$ decays, the hadronic amplitude relations or form factor relations of the $D\to T\ell^{+}\nu_{\ell}$ decays can be obtained from Eq. (12) with $M=T$ by using the SU(3) flavor symmetry approach, and are listed in Tab. 11. The SU(3) amplitudes of the $D\to T\ell^{+}\nu_{\ell}$ decays contain non-perturbation coefficients $T_{1,2,3,4}$; if neglecting the SU(3) flavor breaking $c^{T}_{1}$ and $c^{T}_{2}$ terms, there is only one non-perturbation parameter $T_{1}=T_{2}=T_{3}=T_{4}=c^{T}_{0}$.
[…]
One can see that the predictions of $\mathcal{B}(D^{0}\to K^{*}_{2}(1430)^{-}\ell^{+}\nu_{\ell})$, $\mathcal{B}(D^{+}\to\overline{K}^{*}_{2}(1430)^{0}\ell^{+}\nu_{\ell})$, $\mathcal{B}(D_{s}^{+}\to f_{2}(1270)\ell^{+}\nu_{\ell})$, $\mathcal{B}(D_{s}^{+}\to f_{2}^{\prime}(1525)\ell^{+}\nu_{\ell})$, and $\mathcal{B}(D^{+}\to f_{2}(1270)e^{+}\nu_{e})$ reached the order of $10^{-5}$. Other branching ratios are obviously suppressed by the CKM matrix element $V^{*}_{cd}$.
Table 10: The hadronic amplitudes for $D\to T\ell^{+}\nu$ decays. $T_{1}\equiv c^{T}_{0}+c^{T}_{1}-2c^{T}_{2}$, $T_{2}\equiv c^{T}_{0}-2c^{T}_{1}-2c^{T}_{2}$, $T_{3}\equiv c^{T}_{0}+c^{T}_{1}+c^{T}_{2}$, $T_{4}\equiv c^{T}_{0}-2c^{T}_{1}+c^{T}_{2}$. $T_{1}=T_{2}=T_{3}=T_{4}=c^{T}_{0}$ if neglecting the SU(3) flavor breaking $c^{T}_{1}$ and $c^{T}_{2}$ terms.
Hadronic helicity amplitudes SU(3) amplitudes Hadronic helicity amplitudes SU(3) amplitudes
$H(D^{0}\to K^{*}_{2}(1430)^{-}\ell^{+}\nu_{\ell})$ $T_{1}\leavevmode\nobreak\ V^{*}_{cs}$ $H(D^{0}\to a_{2}(1320)^{-}\ell^{+}\nu_{\ell})$ $T_{3}\leavevmode\nobreak\ V^{*}_{cd}$
$H(D^{+}\to\overline{K}^{*}_{2}(1430)^{0}\ell^{+}\nu_{\ell})$ $T_{1}\leavevmode\nobreak\ V^{*}_{cs}$ $H(D^{+}\to a_{2}(1320)^{0}\ell^{+}\nu_{\ell})$ $-\frac{1}{\sqrt{2}}T_{3}\leavevmode\nobreak\ V^{*}_{cd}$
$H(D_{s}^{+}\to f_{2}(1270)\ell^{+}\nu_{\ell})$ $T_{2}\leavevmode\nobreak\ V^{*}_{cs}\leavevmode\nobreak\ sin\theta_{f_{2}}$ $H(D^{+}\to f_{2}(1270)\ell^{+}\nu_{\ell})$ $\frac{1}{\sqrt{2}}T_{3}\leavevmode\nobreak\ V^{*}_{cd}\leavevmode\nobreak\ cos\theta_{f_{2}}$
$H(D_{s}^{+}\to f_{2}^{\prime}(1525)\ell^{+}\nu_{\ell})$ $-T_{2}\leavevmode\nobreak\ V^{*}_{cs}\leavevmode\nobreak\ cos\theta_{f_{2}}$ $H(D^{+}\to f_{2}^{\prime}(1525)\ell^{+}\nu_{\ell})$ $\frac{1}{\sqrt{2}}T_{3}\leavevmode\nobreak\ V^{*}_{cd}\leavevmode\nobreak\ sin\theta_{f_{2}}$
$H(D_{s}^{+}\to K^{*}_{2}(1430)^{0}\ell^{+}\nu_{\ell})$ $T_{4}\leavevmode\nobreak\ V^{*}_{cd}$
[…]
      Branching ratios       Predictions $(\ell=e)$       Predictions $(\ell=\mu)$
      $\mathcal{B}(D^{0}\to K^{*}_{2}(1430)^{-}\ell^{+}\nu_{\ell})$       $1.36\pm 0.49$       $0.91\pm 0.31$
      $\mathcal{B}(D^{+}\to\overline{K}^{*}_{2}(1430)^{0}\ell^{+}\nu_{\ell})$       $3.40\pm 1.21$       $2.25\pm 0.76$
      $\mathcal{B}(D_{s}^{+}\to f_{2}(1270)\ell^{+}\nu_{\ell})$       $1.04\pm 0.54$       $0.87\pm 0.44$
      $\mathcal{B}(D_{s}^{+}\to f_{2}^{\prime}(1525)\ell^{+}\nu_{\ell})$       $1.83\pm 0.69$       $1.25\pm 0.45$
      $\mathcal{B}(D^{0}\to a_{2}(1320)^{-}\ell^{+}\nu_{\ell})$       $0.35\pm 0.13$       $0.27\pm 0.09$
      $\mathcal{B}(D^{+}\to a_{2}(1320)^{0}\ell^{+}\nu_{\ell})$       $0.47\pm 0.17$       $0.36\pm 0.12$
      $\mathcal{B}(D^{+}\to f_{2}(1270)\ell^{+}\nu_{\ell})$       $0.78\pm 0.29$       $0.62\pm 0.21$
      $\mathcal{B}(D^{+}\to f_{2}^{\prime}(1525)\ell^{+}\nu_{\ell})$       $(5.35\pm 2.83)\times 10^{-4}$       $(2.90\pm 1.48)\times 10^{-4}$
      $\mathcal{B}(D_{s}^{+}\to K^{*}_{2}(1430)^{0}\ell^{+}\nu_{\ell})$       $0.34\pm 0.13$       $0.26\pm 0.09$
III.2 Nonleptonic $T\to VP$ decays
The branching ratios of the $T\to VP$ decays can be written as Li:2015xka
$\mathcal{B}(T\to VP)=\frac{\tau_{T}|\lambda_{TVP}|^{2}(q^{\prime}_{T})^{5}}{20\pi},$ (36)
where $q^{\prime}_{T}\equiv\frac{1}{2m_{T}}\lambda^{1/2}(m_{T}^{2},m_{P}^{2},m_{V}^{2})$, and $\lambda_{TVP}$ is the coefficient of the $TVP$ vertex, and $\tau_{T}$ is the lifetime of the tensor meson. Since most of the $T\to VP$ modes decay by small phases, the mass distribution of tensor resonances and vector mesons is considered in all $T\to VP$ decays, which are similar to the $A\to VP$ decays in Eq. (19) by replacing $A$ with $T$.
Under the SU(3) flavor symmetry, the $\lambda_{TVP}$ of different $T\to VP$ decays can be related. The $TVP$ vertex coefficients can be obtained by replacing $A$ with $T$ in Eq. (20), and the results are listed in Tab. 12. After ignoring the SU(3) flavor breaking terms, we see that all the vertex coefficients of the $T\to VP$ decays are related to the non-perturbative coefficients $c^{\prime T}_{0}$. Apart from the mixing angle $\theta_{P}$, there remains only $\theta_{f_{2}}$ in the $f_{2}(1270)/f_{2}^{\prime}(1525)\to VP$ decay modes.
Table 12: The vertex coefficients of $T\to VP$ decays under the SU(3) flavor symmetry. $a^{T}_{1}=c^{\prime T}_{0}-c^{\prime T}_{1}+c^{\prime T}_{3}$, $a^{T}_{2}=c^{\prime T}_{0}-c^{\prime T}_{1}-2c^{\prime T}_{3}$, $a^{T}_{3}=c^{\prime T}_{0}+2c^{\prime T}_{1}+c^{\prime T}_{3}$, $a^{T}_{4}=c^{\prime T}_{0}+2c^{\prime T}_{1}-2c^{\prime T}_{3}$, $a^{T}_{5}=c^{\prime T}_{0}-4c^{\prime T}_{1}+c^{\prime T}_{3}$. $a^{T}_{1}=a^{T}_{2}=a^{T}_{3}=a^{T}_{4}=a^{T}_{5}=c^{\prime T}_{0}$ if neglecting the SU(3) flavor breaking $c^{\prime T}_{1}$ and $c^{\prime T}_{3}$ terms.
Helicity amplitudes SU(3) flavor amplitudes Helicity amplitudes SU(3) flavor amplitudes
$\overline{K}_{2}^{*}(1430)^{0}\to\rho^{+}K^{-}$ $-a^{T}_{1}$ $\overline{K}_{2}^{*}(1430)^{0}\to\overline{K}^{*0}\eta$ $\frac{1}{\sqrt{6}}\cos\theta_{P}(a^{T}_{1}+2a^{T}_{2})-\frac{1}{\sqrt{3}}\sin\theta_{P}(a^{T}_{1}-a^{T}_{2})$
$\overline{K}_{2}^{*}(1430)^{0}\to\rho^{0}\overline{K}^{0}$ $\frac{1}{\sqrt{2}}a^{T}_{1}$ $K_{2}^{*}(1430)^{-}\to K^{*-}\eta$ $\frac{1}{\sqrt{6}}\cos\theta_{P}(a^{T}_{1}+2a^{T}_{2})-\frac{1}{\sqrt{3}}\sin\theta_{P}(a^{T}_{1}-a^{T}_{2})$
$K_{2}^{*}(1430)^{-}\to\rho^{0}K^{-}$ $-\frac{1}{\sqrt{2}}a^{T}_{1}$ $\overline{K}_{2}^{*}(1430)^{0}\to\overline{K}^{*0}\eta^{\prime}$ $\frac{1}{\sqrt{6}}\sin\theta_{P}(a^{T}_{1}+2a^{T}_{2})+\frac{1}{\sqrt{3}}\cos\theta_{P}(a^{T}_{1}-a^{T}_{2})$
$K_{2}^{*}(1430)^{-}\to\rho^{-}\overline{K}^{0}$ $-a^{T}_{1}$ $K_{2}^{*}(1430)^{-}\to K^{*-}\eta^{\prime}$ $\frac{1}{\sqrt{6}}\sin\theta_{P}(a^{T}_{1}+2a^{T}_{2})+\frac{1}{\sqrt{3}}\cos\theta_{P}(a^{T}_{1}-a^{T}_{2})$
$\overline{K}_{2}^{*}(1430)^{0}\to K^{*-}\pi^{+}$ $a^{T}_{1}$ $\overline{K}_{2}^{*}(1430)^{0}\to\omega\overline{K}^{0}$ $-\frac{1}{\sqrt{2}}a^{T}_{1}$
$\overline{K}_{2}^{*}(1430)^{0}\to\overline{K}^{*0}\pi^{0}$ $-\frac{1}{\sqrt{2}}a^{T}_{1}$ $K_{2}^{*}(1430)^{-}\to\omega K^{-}$ $-\frac{1}{\sqrt{2}}a^{T}_{1}$
$K_{2}^{*}(1430)^{-}\to K^{*-}\pi^{0}$ $\frac{1}{\sqrt{2}}a^{T}_{1}$ $\overline{K}_{2}^{*}(1430)^{0}\to\phi\overline{K}^{0}$ $a^{T}_{2}$
$K_{2}^{*}(1430)^{-}\to\overline{K}^{*0}\pi^{-}$ $a^{T}_{1}$ $K_{2}^{*}(1430)^{-}\to\phi K^{-}$ $a^{T}_{2}$
$a_{2}(1320)^{0}\to\rho^{+}\pi^{-}$ $\sqrt{2}a^{T}_{3}$ $a_{2}(1320)^{0}\to K^{*0}\overline{K}^{0}$ $-\frac{1}{\sqrt{2}}a^{T}_{4}$
$a_{2}(1320)^{0}\to\rho^{-}\pi^{+}$ $-\sqrt{2}a^{T}_{3}$ $a_{2}(1320)^{0}\to\overline{K}^{*0}K^{0}$ $\frac{1}{\sqrt{2}}a^{T}_{4}$
$a_{2}(1320)^{-}\to\rho^{-}\pi^{0}$ $\sqrt{2}a^{T}_{3}$ $a_{2}(1320)^{0}\to K^{*+}K^{-}$ $\frac{1}{\sqrt{2}}a^{T}_{4}$
$a_{2}(1320)^{-}\to\rho^{0}\pi^{-}$ $-\sqrt{2}a^{T}_{3}$ $a_{2}(1320)^{0}\to K^{*-}K^{+}$ $-\frac{1}{\sqrt{2}}a^{T}_{4}$
$a_{2}(1320)^{-}\to K^{*0}K^{-}$ $a^{T}_{4}$
$a_{2}(1320)^{-}\to K^{*-}K^{0}$ $-a^{T}_{4}$
$f_{2}(1270)\to K^{*+}K^{-}$ $a^{T}_{4}\frac{\cos\theta_{f_{2}}}{\sqrt{2}}-a^{T}_{5}\sin\theta_{f_{2}}$ $f_{2}^{\prime}(1525)\to K^{*+}K^{-}$ $a^{T}_{4}\frac{\sin\theta_{f_{2}}}{\sqrt{2}}+a^{T}_{5}\cos\theta_{f_{2}}$
$f_{2}(1270)\to K^{*-}K^{+}$ $-a^{T}_{4}\frac{\cos\theta_{f_{2}}}{\sqrt{2}}+a^{T}_{5}\sin\theta_{f_{2}}$ $f_{2}^{\prime}(1525)\to K^{*-}K^{+}$ $-a^{T}_{4}\frac{\sin\theta_{f_{2}}}{\sqrt{2}}-a^{T}_{5}\cos\theta_{f_{2}}$
$f_{2}(1270)\to K^{*0}\overline{K}^{0}$ $a^{T}_{4}\frac{\cos\theta_{f_{2}}}{\sqrt{2}}-a^{T}_{5}\sin\theta_{f_{2}}$ $f_{2}^{\prime}(1525)\to K^{*0}\overline{K}^{0}$ $a^{T}_{4}\frac{\sin\theta_{f_{2}}}{\sqrt{2}}+a^{T}_{5}\cos\theta_{f_{2}}$
$f_{2}(1270)\to\overline{K}^{*0}K^{0}$ $-a^{T}_{4}\frac{\cos\theta_{f_{2}}}{\sqrt{2}}+a^{T}_{5}\sin\theta_{f_{2}}$ $f_{2}^{\prime}(1525)\to\overline{K}^{*0}K^{0}$ $-a^{T}_{4}\frac{\sin\theta_{f_{2}}}{\sqrt{2}}-a^{T}_{5}\cos\theta_{f_{2}}$
Now only some $T\to VP$ decays have been measured PDG2022
$\mathcal{B}(K_{2}^{*}(1430)\to K^{*}\pi)$ $=$ $(24.7\pm 1.5)\%,$ (37)
$\mathcal{B}(K_{2}^{*}(1430)\to\rho K)$ $=$ $(8.7\pm 0.8)\%,$ (38)
$\mathcal{B}(K_{2}^{*}(1430)\to\omega K)$ $=$ $(2.9\pm 0.8)\%,$ (39)
$\mathcal{B}(a_{2}(1320)\to\rho\pi)$ $=$ $(70.1\pm 2.7)\%.$ (40)
The above experimental data could be used to determine the $TVP$ vertex coefficients $\lambda_{TVP}$, and we obtain the following: $|\lambda_{TVP}|=8.22\pm 0.33$ from $K_{2}^{*}(1430)\to K^{*}\pi$, $|\lambda_{TVP}|=7.14\pm 0.39$ from $K_{2}^{*}(1430)\to\rho K$, $|\lambda_{TVP}|=8.81\pm 1.32$ from $K_{2}^{*}(1430)\to\omega K$, and $|\lambda_{TVP}|=8.44\pm 0.23$ from $a_{2}(1320)\to\rho\pi$. One can see that although $|\lambda_{TVP}|$ from $K_{2}^{*}(1430)\to\rho K$ is slightly small, the other three values are coincident with each other within $1\sigma$ error.
In our numerical results, we use each constrained $\lambda_{TVP}$ to predict their specific processes. For examples, $|\lambda_{TVP}|=8.22\pm 0.33$ from $K_{2}^{*}(1430)\to K^{*}\pi$ is used to predict the branching ratios of the $\overline{K}_{2}^{*}(1430)^{0}\to K^{*-}\pi^{+}$, $\overline{K}_{2}^{*}(1430)^{0}\to\overline{K}^{*0}\pi^{0}$, $K_{2}^{*}(1430)^{-}\to K^{*-}\pi^{0}$, $K_{2}^{*}(1430)^{-}\to\overline{K}^{*0}\pi^{-}$ decays, and $|\lambda_{TVP}|=8.44\pm 0.23$ from $a_{2}(1320)\to\rho\pi$ is used to predict the branching ratios of the $a_{2}(1320)^{0}\to\rho^{+}\pi^{-}$, $a_{2}(1320)^{0}\to\rho^{-}\pi^{+}$, $a_{2}(1320)^{-}\to\rho^{-}\pi^{0}$, $a_{2}(1320)^{-}\to\rho^{0}\pi^{-}$ decays. For the as-yet-unmeasured processes, for examples, $K_{2}^{*}(1430)\to K^{*}\eta$, $a_{2}(1320)\to K^{*}K$ and $f_{2}(1270)\to K^{*}K$, we will use $|\lambda_{TVP}|=8.22\pm 0.33$ from $K_{2}^{*}(1430)\to K^{*}\pi$ to predict their branching ratios.
Our predictions are listed in Tab. 13. One can see that $\mathcal{B}(K^{*}_{2}(1430)\to\rho K)$, $\mathcal{B}(K^{*}_{2}(1430)\to K^{*}\pi)$, $\mathcal{B}(K^{*}_{2}(1430)\to\omega K)$, $\mathcal{B}(a_{2}(1320)\to\rho\pi)$, and $\mathcal{B}(f^{\prime}_{2}(1525)\to K^{*}K)$ are on the order of $10^{-2}-10^{-1}$; other processes are strongly suppressed by their small phase spaces.
Table 13: Branching ratios of the $T\to VP$ decays within $1\sigma$ error (in units of $10^{-2}$). Edenotes the experimental data.
Our predictions Our predictions
$\mathcal{B}(\overline{K}_{2}^{*}(1430)^{0}\to\rho^{+}K^{-})$ $6.07\pm 0.63$ $\mathcal{B}(\overline{K}_{2}^{*}(1430)^{0}\to\overline{K}^{*0}\eta)$ $0.56\pm 0.08$
$\mathcal{B}(\overline{K}_{2}^{*}(1430)^{0}\to\rho^{0}\overline{K}^{0})$ $2.93\pm 0.31$ $\mathcal{B}(K_{2}^{*}(1430)^{-}\to K^{*-}\eta)$ $0.56\pm 0.08$
$\mathcal{B}(K_{2}^{*}(1430)^{-}\to\rho^{0}K^{-})$ $2.86\pm 0.30$ $\mathcal{B}(\overline{K}_{2}^{*}(1430)\to K^{*}\eta)$ $0.56\pm 0.08$
$\mathcal{B}(K_{2}^{*}(1430)^{-}\to\rho^{-}\overline{K}^{0})$ $5.57\pm 0.58$ $\mathcal{B}(\overline{K}_{2}^{*}(1430)^{0}\to\omega\overline{K}^{0})$ $2.94\pm 0.85$
$\mathcal{B}(K_{2}^{*}(1430)\to\rho K)$ $8.70\pm 0.80^{E}$ $\mathcal{B}(K_{2}^{*}(1430)^{-}\to\omega K^{-})$ $2.88\pm 0.84$
$\mathcal{B}(\overline{K}_{2}^{*}(1430)^{0}\to K^{*-}\pi^{+})$ $17.01\pm 1.18$ $\mathcal{B}(K_{2}^{*}(1430)\to\omega K)$ $2.90\pm 0.80^{E}$
$\mathcal{B}(\overline{K}_{2}^{*}(1430)^{0}\to\overline{K}^{*0}\pi^{0})$ $8.34\pm 0.58$ $\mathcal{B}(\overline{K}_{2}^{*}(1430)^{0}\to\phi\overline{K}^{0})$ $0.040\pm 0.008$
$\mathcal{B}(K_{2}^{*}(1430)^{-}\to K^{*-}\pi^{0})$ $8.25\pm 0.58$ $\mathcal{B}(K_{2}^{*}(1430)^{-}\to\phi K^{-})$ $0.038\pm 0.008$
$\mathcal{B}(K_{2}^{*}(1430)^{-}\to\overline{K}^{*0}\pi^{-})$ $15.82\pm 1.10$ $\mathcal{B}(K_{2}^{*}(1430)\to\phi K)$ $0.039\pm 0.007$
$\mathcal{B}(K_{2}^{*}(1430)\to K^{*}\pi)$ $24.70\pm 1.50^{E}$
$\mathcal{B}(a_{2}(1320)^{0}\to\rho^{+}\pi^{-})$ $34.99\pm 1.37$ $\mathcal{B}(a_{2}(1320)^{0}\to K^{*0}\overline{K}^{0})$ $0.043\pm 0.005$
$\mathcal{B}(a_{2}(1320)^{0}\to\rho^{-}\pi^{+})$ $34.99\pm 1.37$ $\mathcal{B}(a_{2}(1320)^{0}\to\overline{K}^{*0}K^{0})$ $0.043\pm 0.005$
$\mathcal{B}(a_{2}(1320)^{-}\to\rho^{-}\pi^{0})$ $35.33\pm 1.38$ $\mathcal{B}(a_{2}(1320)^{0}\to K^{*+}K^{-})$ $0.055\pm 0.006$
$\mathcal{B}(a_{2}(1320)^{-}\to\rho^{0}\pi^{-})$ $34.99\pm 1.40$ $\mathcal{B}(a_{2}(1320)^{0}\to K^{*-}K^{+})$ $0.055\pm 0.006$
$\mathcal{B}(a_{2}(1320)\to\rho\pi)$ $70.10\pm 2.70^{E}$ $\mathcal{B}(a_{2}(1320)^{-}\to K^{*0}K^{-})$ $0.094\pm 0.010$
$\mathcal{B}(a_{2}(1320)^{-}\to K^{*-}K^{0})$ $0.10\pm 0.01$
$\mathcal{B}(a_{2}(1320)\to K^{*}K)$ $0.19\pm 0.02$
$\mathcal{B}(f_{2}(1270)\to K^{*+}K^{-})$ $0.072\pm 0.013$ $\mathcal{B}(f_{2}^{\prime}(1525)\to K^{*+}K^{-})$ $5.89\pm 0.75$
$\mathcal{B}(f_{2}(1270)\to K^{*-}K^{+})$ $0.072\pm 0.013$ $\mathcal{B}(f_{2}^{\prime}(1525)\to K^{*-}K^{+})$ $5.89\pm 0.75$
$\mathcal{B}(f_{2}(1270)\to K^{*0}\overline{K}^{0})$ $0.064\pm 0.011$ $\mathcal{B}(f_{2}^{\prime}(1525)\to K^{*0}\overline{K}^{0})$ $5.25\pm 0.66$
$\mathcal{B}(f_{2}(1270)\to\overline{K}^{*0}K^{0})$ $0.064\pm 0.011$ $\mathcal{B}(f_{2}^{\prime}(1525)\to\overline{K}^{*0}K^{0})$ $5.25\pm 0.66$
$\mathcal{B}(f_{2}(1270)\to K^{*}K)$ $0.273\pm 0.048$ $\mathcal{B}(f_{2}^{\prime}(1525)\to K^{*}K)$ $22.28\pm 2.80$
III.3 Results of the the $D\to T(T\to VP)\ell^{+}\nu_{\ell}$ decays
Using the predictions of $\mathcal{B}(D\rightarrow T\ell^{+}\nu_{\ell})$ given in Tab. 11 and $\mathcal{B}(T\rightarrow VP)$ given in Tab. 13, the branching ratio predictions of the $D\to VP\ell^{+}\nu_{\ell}$ decays with tensor resonances can be obtained, and they are listed in the third and fifth columns of Tabs. 7-9.
As given in Tabs. 7-9, some processes have both tensor and axial-vector resonant contributions, while other processes only receive axial-vector resonant contributions. Comparing the tensor resonant contributions with the axial-vector resonant contributions in the decays with the same initial states and the same final states, one can see that the axial-vector resonant contributions are dominant.
IV Summary
Four-body semileptonic decays $D\to VP\ell^{+}\nu_{\ell}$ with axial-vector resonance states and tensor resonance states were studied using the SU(3) flavor symmetry/breaking approach. The decay amplitudes (vertex coefficients) of the $D\to M\ell^{+}\nu_{\ell}$ ($M\to VP$) decays were related, and the allowed spaces of relevant non-perturbative parameters of $D\to A\ell^{+}\nu_{\ell}$, $A\to VP$, and $T\to VP$ decays were obtained by using the present experimental data. The form factors of $D\to T\ell^{+}\nu_{\ell}$ decays were taken from the results in the light-front quark model since no experimental data are available at present. Finally, in terms of the predicted branching ratios of the $D\to M\ell^{+}\nu_{\ell}$ and $M\to VP$ decays, the branching ratios of the $D\to M(M\to VP)\ell^{+}\nu_{\ell}$ decays were obtained using the narrow width approximation. In addition, the sensitivity of the branching ratios of the $D\to A\ell^{+}\nu_{\ell}$ and $A\to VP$ decays to the constrained non-perturbative parameter spaces was explored. Our main results can be summarized as follows.
1) Decays $D\to A(A\to VP)\ell^{+}\nu_{\ell}$: $\mathcal{B}(D\to K_{1}(1270)\ell^{+}\nu_{\ell})$ and $\mathcal{B}(D_{s}\to f_{1}/h_{1}\ell^{+}\nu_{\ell})$ are on the order of $\mathcal{O}(10^{-4}-10^{-3})$, which may be measured experimentally in the near future. $\mathcal{B}(D\to K_{1}(1400)\ell^{+}\nu_{\ell})$ are strongly suppressed by the mixing angle $\theta_{K_{1}}$; $\mathcal{B}(D\to a_{1}/b_{1}/f_{1}/h_{1}\ell^{+}\nu_{\ell})$ and $\mathcal{B}(D_{s}\to K_{1}(1270)/K_{1}(1400)\ell^{+}\nu_{\ell})$ are strongly suppressed by the CKM matrix element $V_{cd}$, and they are on the order of $\mathcal{O}(10^{-6}-10^{-5})$. We also found that some decay branching ratios are very sensitive to the non-perturbative parameters $A$ and $B$, or the mixing angles $\theta_{K_{1}}$, $\theta_{3P_{1}}$, or $\theta_{1P_{1}}$. The experimental data for the $A\to VP$ decays give relatively strong constraints on the non-perturbative parameters and the mixing angles. Some branching ratios were obtained with the large errors due to our conservative choice of parameters and the large experimental errors. We found that $\mathcal{B}(K_{1}(1270)\to K^{*}\pi,K^{*}\eta)$ are sensitive to $F$ and $\theta_{K1}$. $\mathcal{B}(K_{1}(1400)\to\rho K,\omega K,K^{*}\eta)$ are sensitive to $D$ and $\theta_{K1}$, $\mathcal{B}(K_{1}(1400)\to\rho K,\omega K)$ have the inflection point values when $D\approx-1.0$ and $\theta_{K_{1}}\approx 56^{\circ}$, and many $\mathcal{B}(a_{1}/b_{1}/f_{1}/h_{1}\to VP)$ are very sensitive to $F$, $D$, $\theta_{3_{P1}}$ or $\theta_{1_{P1}}$. Any future measurement of these decays can obviously constrain the relevant parameters. Our predictions of $\mathcal{B}(D^{+}\to b_{1}(1235)^{0}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{0}\to\omega\pi^{0})$ and $\mathcal{B}(D^{0}\to b_{1}(1235)^{-}e^{+}\nu_{e},\leavevmode\nobreak\ b_{1}(1235)^{-}\to\omega\pi^{-})$ are much smaller than their present experimental upper limits. We found that many $\mathcal{B}(D\to A\ell^{+}\nu_{\ell},A\to VP)$ are on the order of $\mathcal{O}(10^{-5}-10^{-4})$, and they are expected to be measured in the near future.
2) Decays $D\to T(T\to VP)\ell^{+}\nu_{\ell}$: We found that the predictions of $\mathcal{B}(D^{0}\to K^{*}_{2}(1430)^{-}\ell^{+}\nu_{\ell})$, $\mathcal{B}(D^{+}\to\overline{K}^{*}_{2}(1430)^{0}\ell^{+}\nu_{\ell})$, $\mathcal{B}(D_{s}^{+}\to f_{2}(1270)\ell^{+}\nu_{\ell})$, $\mathcal{B}(D_{s}^{+}\to f_{2}^{\prime}(1525)\ell^{+}\nu_{\ell})$, and $\mathcal{B}(D^{+}\to f_{2}(1270)e^{+}\nu_{e})$ could reach on the order of $10^{-5}$; nevertheless, other branching ratios of the $D\to T\ell^{+}\nu_{\ell}$ are obviously suppressed by the CKM matrix element $V^{*}_{cd}$. For the $T\to VP$ decays, $\mathcal{B}(K^{*}_{2}(1430)\to\rho K)$, $\mathcal{B}(K^{*}_{2}(1430)\to K^{*}\pi)$, $\mathcal{B}(K^{*}_{2}(1430)\to\omega K)$, $\mathcal{B}(a_{2}(1320)\to\rho\pi)$ and $\mathcal{B}(f^{\prime}_{2}(1525)\to K^{*}K)$ are on the order of $10^{-2}-10^{-1}$; however, $\mathcal{B}(K^{*}_{2}(1430)\to K^{*}\eta,\phi K)$, $\mathcal{B}(a_{2}(1320)\to K^{*}K)$, and $\mathcal{B}(f_{2}(1270)\to K^{*}K)$ are on the order of $10^{-4}-10^{-3}$. As for the branching ratios of the $D\to T(T\to VP)\ell^{+}\nu_{\ell}$ decays, some processes received both the tensor and axial-vector resonant contributions. We found that the axial-vector resonant contributions were dominant, and the tensor resonant contributions were obviously suppressed.
Although SU(3) flavor symmetry is approximate, it can still provide very useful information about these decays. According to our rough predictions, some decays of $D\to M\ell^{+}\nu_{\ell}$, $M\to VP$, and $D\to M(M\to VP)\ell^{+}\nu_{\ell}$ might be measured experimentally in the near future, and they are useful for testing and understanding the non-perturbative effects in the semileptonic decays. Nevertheless, some decay branching ratios are very small and could not be measured recently.
ACKNOWLEDGEMENTS
We thank Hai-Bo Li, Mao-Zhi Yang and Xian-Wei Kang for useful discussions. The work was supported by the National Natural Science Foundation of China (No. 12175088 and No. 12365014).
Appendix
We present the meson multiplets of the SU(3) flavor group in this section. Charmed pseudoscalar mesons are composed of a charm quark and a light anti-quark
$D^{i}=\Big{(}D^{0}(c\bar{u}),\leavevmode\nobreak\ D^{+}(c\bar{d}),\leavevmode\nobreak\ D^{+}_{s}(c\bar{s})\Big{)}.$ (41)
Light pseudoscalar mesons and vector mesons are represented as He:2018joe
$P$ $=$ $\left(\begin{array}[]{ccc}\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta_{8}}{\sqrt{6}}+\frac{\eta_{1}}{\sqrt{3}}&\pi^{+}&K^{+}\\ \pi^{-}&-\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta_{8}}{\sqrt{6}}+\frac{\eta_{1}}{\sqrt{3}}&K^{0}\\ K^{-}&\overline{K}^{0}&-\frac{2\eta_{8}}{\sqrt{6}}+\frac{\eta_{1}}{\sqrt{3}}\end{array}\right)\,,$ (45)
$V$ $=$ $\left(\begin{array}[]{ccc}\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}&\rho^{+}&K^{*+}\\ \rho^{-}&-\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}&K^{*0}\\ K^{*-}&\overline{K}^{*0}&\phi\end{array}\right)\,.$ (49)
The pseudoscalar mesons $\eta$ and $\eta^{\prime}$ are the mixtures of $\eta_{1}=\frac{u\bar{u}+d\bar{d}+s\bar{s}}{\sqrt{3}}$ and $\eta_{8}=\frac{u\bar{u}+d\bar{d}-2s\bar{s}}{\sqrt{6}}$ with a mixing angle denoted as $\theta_{P}$ Bramon:1992kr
$\left(\begin{array}[]{c}\eta\\ \eta^{\prime}\end{array}\right)\,=\left(\begin{array}[]{cc}\mbox{cos}\theta_{P}&-\mbox{sin}\theta_{P}\\ \mbox{sin}\theta_{P}&\mbox{cos}\theta_{P}\end{array}\right)\,\left(\begin{array}[]{c}\eta_{8}\\ \eta_{1}\end{array}\right)\,.$ (56)
In our numerical analysis, we will utilize the mixing angle range of $\theta_{P}=[-20^{\circ},-10^{\circ}]$ from the Particle Data Group (PDG) PDG2022 .
The multiplets of the light axial-vector mesons are presented in the matrix of $A$ and $B$, where $A$ represents the axial-vector mesons with $J^{PC}=1^{++}$ and $B$ corresponds to the mesons with $J^{PC}=1^{+-}$ Roca:2003uk
$A$ $=$ $\left(\begin{array}[]{ccc}\frac{a^{0}_{1}}{\sqrt{2}}+\frac{f_{1}}{\sqrt{3}}+\frac{f_{8}}{\sqrt{6}}&a^{+}_{1}&K^{+}_{1A}\\ a^{-}_{1}&-\frac{a^{0}_{1}}{\sqrt{2}}+\frac{f_{1}}{\sqrt{3}}+\frac{f_{8}}{\sqrt{6}}&K^{0}_{1A}\\ K^{-}_{1A}&\overline{K}^{0}_{1A}&\frac{f_{1}}{\sqrt{3}}-\frac{2f_{8}}{\sqrt{6}}\end{array}\right)\,,$ (60)
$B$ $=$ $\left(\begin{array}[]{ccc}\frac{b^{0}_{1}}{\sqrt{2}}+\frac{h_{1}}{\sqrt{3}}+\frac{h_{8}}{\sqrt{6}}&b^{+}_{1}&K^{+}_{1B}\\ b^{-}_{1}&-\frac{b^{0}_{1}}{\sqrt{2}}+\frac{h_{1}}{\sqrt{3}}+\frac{h_{8}}{\sqrt{6}}&K^{0}_{1B}\\ K^{-}_{1B}&\overline{K}^{0}_{1B}&\frac{h_{1}}{\sqrt{3}}-\frac{2h_{8}}{\sqrt{6}}\end{array}\right)\,.$ (64)
Mesons $K_{1}(1270)$ and $K_{1}(1400)$ are the mixtures of $K_{1A}$ and $K_{1B}$ by the mixing angle $\theta_{K_{1}}$
$\left(\begin{array}[]{c}K_{1}(1270)\\ K_{1}(1400)\end{array}\right)\,=\left(\begin{array}[]{cc}sin\theta_{K_{1}}&cos\theta_{K_{1}}\\ cos\theta_{K_{1}}&-sin\theta_{K_{1}}\end{array}\right)\,\left(\begin{array}[]{c}K_{1A}\\ K_{1B}\end{array}\right).$ (71)
Mesons $f_{1}(1285)$, $f_{1}(1420)$ ($h_{1}(1170)$, $h_{1}(1415)$) can be expressed by the mixing angle $\theta_{3P_{1}}$ ($\theta_{1P_{1}}$)
$\left(\begin{array}[]{c}f_{1}(1285)\\ f_{1}(1420)\end{array}\right)\,=\left(\begin{array}[]{cc}cos\theta_{3P_{1}}&sin\theta_{3P_{1}}\\ -sin\theta_{3P_{1}}&cos\theta_{3P_{1}}\end{array}\right)\,\left(\begin{array}[]{c}f_{1}\\ f_{8}\end{array}\right),$ (78)
$\left(\begin{array}[]{c}h_{1}(1170)\\ h_{1}(1415)\end{array}\right)\,=\left(\begin{array}[]{cc}cos\theta_{1P_{1}}&sin\theta_{1P_{1}}\\ -sin\theta_{1P_{1}}&cos\theta_{1P_{1}}\end{array}\right)\,\left(\begin{array}[]{c}h_{1}\\ h_{8}\end{array}\right).$ (85)
The value of mixing angle $\theta_{K_{1}}$ has been investigated in various studies, for examples, $33^{\circ}$ and $57^{\circ}$ in Ref. Suzuki:1993yc , $37^{\circ}$ or $58^{\circ}$ Cheng:2003bn , [$35^{\circ},55^{\circ}$] Burakovsky:1997dd , $60^{\circ}$ Tayduganov:2011ui , $(33.6\pm 4.3)^{\circ}$ Divotgey:2013jba , etc. These ranges will not be used in later numerical analysis. We will obtain the ranges of $\theta_{K_{1}}$, $\theta_{1P_{1}}$ and $\theta_{3P_{1}}$ from the $D\to A(A\to VP)\ell^{+}\nu_{\ell}$ decays.
The lowest multiplet of p-wave tensor mesons with $J^{PC}=2^{++}$ take the form as Ecker:2007us ; Chen:2023ybr
$T$ $=$ $\left(\begin{array}[]{ccc}\frac{a_{2}^{0}}{\sqrt{2}}+\frac{f_{2}^{q}}{\sqrt{2}}&a^{+}_{2}&K_{2}^{*+}\\ a^{-}_{2}&-\frac{a^{0}_{2}}{\sqrt{2}}+\frac{f^{q}_{2}}{\sqrt{2}}&K^{*0}_{2}\\ K^{*-}_{2}&\overline{K}^{*0}_{2}&f^{s}_{2}\end{array}\right)\,.$ (89)
$f_{2}(1270)$ and $f_{2}^{\prime}(1525)$ are mixed by $f_{2}^{q}$ and $f_{2}^{s}$ with the mixing angle $\theta_{f_{2}}$
$\left(\begin{array}[]{c}f_{2}(1270)\\ f_{2}^{\prime}(1525)\end{array}\right)\,=\left(\begin{array}[]{cc}cos\theta_{f_{2}}&sin\theta_{f_{2}}\\ sin\theta_{f_{2}}&-cos\theta_{f_{2}}\end{array}\right)\,\left(\begin{array}[]{c}f^{q}_{2}\\ f^{s}_{2}\end{array}\right),$ (96)
where $\leavevmode\nobreak\ \theta_{f_{2}}\in[8^{\circ},10^{\circ}]$ Cheng:2010yd will be used in our calculation.